### v1

# Orthogonal roots, Macdonald representations, and quasiparabolic sets

Abstract.
Let $W$ be a simply laced Weyl group of finite type and rank $n$. If $W$ has type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even, then the root system of $W$ has subsystems of type $nA_{1}$. This gives rise to an irreducible Macdonald representation of $W$ spanned by $n$-roots, which are products of $n$ orthogonal roots in the symmetric algebra of the reflection representation. We prove that in these cases, the set of all maximal sets of orthogonal positive roots has the structure of a quasiparabolic set in the sense of Rains–Vazirani. The quasiparabolic structure can be described in terms of certain quadruples of orthogonal positive roots which we call crossings, nestings, and alignments. This leads to nonnesting and noncrossing bases for the Macdonald representation, as well as some highly structured partially ordered sets. We use the $8$-roots in type $E_{8}$ to give a concise description of a graph that is known to be non-isomorphic but quantum isomorphic to the orthogonality graph of the $E_{8}$ root system.
Key words and phrases:
root system, Macdonald representation, quasiparabolic set, canonical basis, quantum isomorphism
1991 Mathematics Subject Classification:
Primary: 20F55; Secondary: 05E18, 05E10, 06A07.
1. Introduction
If $W$ is a finite simply laced Weyl group, then it is possible to find a basis of the reflection representation $V$ of $W$ that consists of orthogonal positive roots when $W$ has type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even. The goal of this paper is to demonstrate that the set $X=X(W)$ of such bases has a rich combinatorial structure, both by identifying $X$ with a subset of a Macdonald representation of $W$, and by regarding $X$ as a quasiparabolic set. A quasiparabolic set for a Weyl group $W$ is a $W$-set equipped with an integer-valued height function satisfying two axioms that specify how the action of a reflection changes the height. The theory of quasiparabolic sets was developed by Rains and Vazirani, and many examples of quasiparabolic sets may be found in their paper [37] and in Lu’s thesis [32].
Let $W$ be a finite Weyl group with root system $\Phi$ and Dynkin diagram $\Gamma$, and let $V$ be the reflection representation of $W$ defined over ${\mathbb{Q}}$. Let $\Psi$ be a subsystem of $\Phi$, meaning that $\emptyset\neq\Psi\subset\Phi$ and that $\Psi$ is also a root system, and let $\Phi^{+}$ and $\Psi^{+}$ be the set of positive roots in $\Phi$ and $\Psi$, respectively. The (positive definite) inner product on $V$ allows us to identify $V$ with its dual, $V^{*}$. The symmetric algebra $\text{\rm Sym}(V^{*})$ then becomes a $W$-module via the contragredient action $(w.\phi)(x)=\phi(w^{-1}.x)$. Following [33], we define the Macdonald representation $j^{\Gamma}_{\Psi}({\rm{sgn}})=j^{\Phi}_{\Psi}({\rm{sgn}})$ of $W$ given by $\Psi$ to be the cyclic ${\mathbb{Q}}W$-submodule of $\text{\rm Sym}(V^{*})$ generated by $\pi_{\Psi}$, where
$\pi_{\Psi}=\prod_{\alpha\in\Psi^{+}}\alpha.$
A short argument shows that the Macdonald representation is an absolutely irreducible $W$-module. When $W$ has type $E_{7},E_{8}$ or $D_{n}$ for $n$ even, $\Phi$ contains subsystems of type $nA_{1}$. The set $\Psi^{+}$ for any such subsystem $\Psi$ consists of $n$ orthogonal positive roots, and we call an element of the form $w.\pi_{\Psi}$ (for $w\in W$) an $n$-root. Thus, an $n$-root has the form $\alpha=\prod_{i=1}^{n}\beta_{i}$ where the elements $\beta_{i}$ are orthogonal roots. Conversely, since $W$ acts transitively on the set of maximal sets of orthogonal roots (Lemma 3.1.2), every product of $n$ orthogonal roots is an $n$-root.
Given an $n$-root $\alpha=\prod_{i=1}^{n}\beta_{i}$, the factors $\beta_{i}$ are unique up to reordering and multiplication by nonzero scalars, because they are the irreducible factors of $\alpha$ in the unique factorization domain ${\mathbb{Q}}[\alpha_{1},\alpha_{2},\ldots,\alpha_{n}]$, where the $\alpha_{i}$ correspond to the simple roots of $\Phi$. We say $\alpha$ is positive if all the factors $\beta_{i}$ may be taken to be positive, or equivalently if evenly many of the components are negative. If $\alpha=\prod_{i=1}^{n}\beta_{i}$ is a positive root with all the $\beta_{i}$ positive, we call the roots $\beta_{i}$ the components of $\alpha$. An $n$-root $\alpha$ is negative if $-\alpha$ is positive. It is immediate from the definitions that if $\alpha$ is an $n$-root then $-\alpha$ is also an $n$-root, and that precisely one of $\alpha$ and $-\alpha$ is positive, similarly to how roots appear in positive-negative pairs in ordinary root systems. If $\alpha$ is either a root or an $n$-root, we define the absolute value of $\alpha$, denoted $\lvert\alpha\rvert$, to be the positive element in the pair $\{\alpha,-\alpha\}$. (We may view both ordinary roots and $n$-roots as special cases of $k$-roots, by which we mean products in $\text{\rm Sym}(V^{*})$ of $k$ orthogonal roots of $W$ for any fixed integer $1\leq k\leq n$. The notion of $k$-roots plays an important role in our previous papers [24] and [23], and we will occasionally speak of $4$-roots, even when $n\neq 4$, in this paper.)
The set $\Phi^{+}_{n}$ of all positive $n$-roots admits a natural $W$-action given by $w(\alpha)=\lvert w.\alpha\rvert$. Similarly, the set $X$ of sets of $n$ orthogonal roots admits a natural $W$-action given by $w(\{\beta_{1},\cdots,\beta_{n}\})=\{\lvert w(\beta_{1})\rvert,\cdots,\lvert w(\beta_{n})\rvert\}$. The map sending each set $\{\beta_{1},\cdots,\beta_{n}\}\in X$ to the product $\prod_{i=1}^{n}\beta_{i}\in\Phi^{+}_{n}$ respects these two $W$-actions, and we use it to identify $X$ with $\Phi^{+}_{n}$.
We show that $X$ has the structure of a quasiparabolic set under a suitable height function $\lambda$ (Theorem 4.1.5). As we explain in sections 3 and 4, to understand this structure it is useful to consider quadruples $Q=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ of four orthogonal roots with the property that $(\beta_{1}+\beta_{2}+\beta_{3}+\beta_{4})/2$ is also a root. We call such quadruples coplanar quadruples, and show that they fall into three distinct types, called crossings, nestings, and alignments. The height function $\lambda$ is given by $\lambda(\gamma)=C(\gamma)+2N(\gamma)$, where $C(\gamma)$ and $N(\gamma)$ are the numbers of crossings and nestings in $\gamma$, respectively, for each $\gamma\in X$. The terms “crossing”, “nesting”, and “alignment” are motivated by the theory of perfect matchings (Remark 3.2.3).
As a quasiparabolic set, the set $X$ is equiped with a partial order $\leq_{Q}$, which is the weakest partial order such that $x\leq_{Q}rx$ whenever $r$ is a reflection such that $\lambda(x)\leq\lambda(rx)$. We use the theory of quasiparabolic sets to prove that $X$ has a unique maximally aligned $n$-root, $\theta_{A}$, and a unique maximally nesting $n$-root, $\theta_{N}$; these two elements are the unique minimal and maximal element of $X$ with respect to $\leq_{Q}$, respectively (Proposition 4.3.1). The $n$-roots that avoid alignments, or the alignment-free $n$-roots, form a quasiparabolic set $X_{I}\subset X$ of a certain maximal parabolic subgroup $W_{I}$ of $W$. The corresponding partial order on $X_{I}$ allows us to show that $X$ also has a maximally crossing element, $\theta_{C}$, and that it is the unique minimal element of $X_{I}$ (Proposition 5.1.2 (iii)). The set $X_{I}$ has a natural bipartite structure, with the $n$-roots in $X_{I}$ with even levels and those with odd levels partitioning $X_{I}$ into two equal-sized components that are interchanged by every reflection in $W_{I}$ (Remark 5.1.3).
The alignment-free $n$-roots in $X$ are one of three families that avoid a particular type of coplanar quadruple, the other two being those that avoid crossings and avoid nestings. Section 5 studies these three families together. We use a version of Bergman’s diamond lemma to show that the noncrossing elements of $X$ form a basis for the Macdonald representation, as do the nonnesting positive $n$-roots (Theorem 5.2.2). In addition, the noncrossing basis may be viewed as a canonical basis and behaves somewhat like a simple system in a root system (Theorem 5.3.2). The nonnesting basis is naturally parametrized by a particular interval $[1,w_{N}]$ in the weak Bruhat order of $W$ and has the structure of a distributive lattice (Theorem 5.4.5). The element $w_{N}$, which we call the nonnesting element of $W$, is the unique shortest element taking the maximally crossing $n$-root $\theta_{C}$ to the maximally aligned element $\theta_{A}$. We note that the noncrossing basis is essentially the same as a basis that appears in the work of Fan [20, Section 6] and others, although the realization of the basis as polynomials in roots seems to be new.
We say that two positive $n$-roots are sum equivalent or $\sigma$-equivalent if their sets of components have the same sum. We show that the $\sigma$-equivalence classes are in canonical bijection with the nonnesting and the noncrossing elements of $X$ in the following way: each $\sigma$-equivalence class is an interval $[\beta_{1},\beta_{2}]$ in the quasiparabolic order $\leq_{Q}$, where the minimal element $\beta_{1}$ and maximal element $\beta_{2}$ are the unique nonnesting and noncrossing $n$-roots in the class. The alignment-free elements in $X$ form a single $\sigma$-equivalence class that is maximal with respect to a natural order (Proposition 5.5.1). Any set of $\sigma$-equivalence class representatives forms a basis for the Macdonald representation, and the change of basis matrix between any two such bases, such as the one between the nonnesting and noncrossing basis, is always unitriangular with integer entries (Theorem 5.5.2).
As we explain in Section 7.3, the feature-avoiding $n$-roots and $\sigma$-equivalence classes in the set $X$ can all be characterized abstractly using the quasiparabolic structure of $X$, without using the combinatorics of sets of roots. This raises the possibility of extending the notions of alignment-free, noncrossing, and nonnesting elements to more general quasiparabolic sets.
The theory developed in this paper has various natural connections to many previous works. In type $D_{n}$ for an even integer $n=2k$, the $n$-roots correspond naturally to perfect matchings of the set $[n]=\{1,2,\dots,n\}$, and the crossings, nestings, and alignments in $n$-roots recover the correponding notions in the theory of matchings. Besides matchings, the quasiparabolic set $X$ can be identified with the set of fixed-point free involutions in $S_{n}$, which is one of the original motivating examples of a quasiparabolic set [37, Section 4]. The level function $\lambda=C+2N$ appears as a useful statistic on matchings in [45, 19, 14], and has a natural interpretation in the context of combinatorial game theory [29]; see Section 6.1.
The Macdonald representation $j^{\Phi}_{nA_{1}}({\rm{sgn}})$ in type $D_{n}$ for $n$ even recovers a Specht module in a very natural way: the action of the Weyl group $W$ factors through an obvious sign-forgetting map (Equation (2.3.1)) to induce an $S_{n}$-module structure on the Macdonald representation for the symmetric group $S_{n}$, and the resulting module is isomorphic to a realization of the Specht module corresponding to the two-row partition $(k,k)$ due to Rhoades [41] (Proposition 6.1.2). The noncrossing bases and nonnesting bases have been studied extensively as the web basis and the Specht basis, respectively, of the Specht module [42, 28, 27, 25]; see Section 6.1.
In type $E_{7}$, the Macdonald representation contains 135 positive 7-roots and has degree 15 [17, Proposition 4.12]. This representation has a long history, going back the work of Coble in 1916 [15, (65)] on the Göpel variety. There are also applications of $7$-roots to quantum information theory and supergravity [11, Section IV G], [18]. In this case, the elements of the quasiparabolic set $X_{I}$ are in canonical bijection with the 30 distinct labellings of the Fano plane, and the maximal and minimal elements are given by $\{136,145,127,235,246,347,567\}$ and $\{123,145,246,257,347,356,167\}$, respectively (Proposition 6.2.2).
In type $E_{8}$, there are 2025 positive $8$-roots. The bases of orthogonal roots have applications to physics, where they can be used to prove the Kochen–Specker theorem in quantum mechanics [50] (Section 7.2). The Macdonald representation in this case has degree $50$, but seems not to have been studied much before. The quasiparabolic set $X_{I}$ in this case is a bipartite structure with 240 elements. As we explain in Section 6.3, either partite component can be used to define a graph that has an interesting relationship with two strongly regular graphs studied recently by Schmidt [44] (Remark 6.3.5). Those two graphs each have 120 vertices, and they have the remarkable property of being quantum isomorphic (in the sense of [3]) but not isomorphic.
The paper is organized as follows. We recall the basics of root systems in Section 2. Section 3 introduces the key notions of crossings, nestings, and alignments in an $n$-root, and we connect them to the theory of quasiparabolic sets in Section 4. Section 5 studies the alignment-free, noncrossing, and nonnesting $n$-roots. Section 6 discusses the details of $n$-roots in the types $D_{n}$ with $n$ even, $E_{7}$ and $E_{8}$. Section 7 concludes the paper, and includes discussions of the Poincaré polynomial of the set $X$ and of orbits of $n$-roots under the action of Coxeter elements.
2. Review of root systems
In Section 2, we recall the basic properties of simply laced root systems of finite type. We will mostly follow the notation of the first two chapters of [26], except in the case of type $E_{7}$, where we follow [21, Section 4].
2.1. Weyl groups, root systems, and reflection representations
The root systems in this paper will be irreducible root systems of finite type, whose Dynkin diagrams are shown in Figure 1. The vertices of the Dynkin diagram $\Gamma$ index the simple roots $\Pi=\{\alpha_{i}:i\in\Gamma\}$. The root lattice ${\mathbb{Z}}\Pi$ is the free ${\mathbb{Z}}$-module on $\Pi$. We define a ${\mathbb{Z}}$-bilinear form $B$ on ${\mathbb{Z}}\Pi\times{\mathbb{Z}}\Pi$ by
$B(\alpha_{i},\alpha_{j})=\begin{cases}2&\text{\ if \ }i=j;\\ -1&\text{\ if \ $i$\ and\ $j$\ are\ adjacent\ in\ }\Gamma;\\ 0&\text{\ otherwise.}\\ \end{cases}$
If $\alpha_{i}\in\Pi$, then we define the simple reflection $s_{i}=s_{\alpha_{i}}$ to be the ${\mathbb{Z}}$-linear operator ${\mathbb{Z}}\Pi\rightarrow{\mathbb{Z}}\Pi$ given by
$s_{i}(\beta)=\beta-B(\alpha_{i},\beta)\alpha_{i}.$
The Weyl group $W=W(\Gamma)$ is the finite group generated by the simple reflections.
[…]
(a) $A_{n}$
[…]
(b) $D_{n}$
111122223333444455556666
(c) $E_{6}$
1111222233334444555566667777
(d) $E_{7}$
11112222333344445555666677778888
(e) $E_{8}$
Figure 1. Dynkin diagrams of irreducible simply laced Weyl groups
The root system of $W$ is the set $\Phi=\{w(\alpha_{i}):\alpha_{i}\in\Pi,w\in W\}$. Each element of $\Phi$ is called a root. The group $W$ acts transitively on $\Phi$, and the form $B$ is $W$-invariant in the sense that $B(\alpha,\beta)=B(w.\alpha,w.\beta)$ for all $w\in W$ and all $\alpha,\beta\in\Phi$. We say two roots are $\alpha,\beta\in\Phi$ are orthogonal if $B(\alpha,\beta)=0$.
Each root $\alpha\in\Phi$ gives rise to a reflection in $W$, which is the self-inverse ${\mathbb{Z}}$-linear operator $s_{\alpha}:{\mathbb{Z}}\Pi\rightarrow{\mathbb{Z}}\Pi$ generalizing simple reflections and given by the formula
(2.1.1) $s_{\alpha}(\beta)=\beta-B(\alpha,\beta)\alpha.$
The reflections in $W$ form a single conjugacy class. The ${\mathbb{Q}}$-vector space $V:={\mathbb{Q}}\otimes_{\mathbb{Z}}{\mathbb{Z}}\Pi$ affords the reflection representation of $W$, where each reflection $s_{\alpha}$ acts by Equation (2.1.1).
A subset $\Psi$ of $\Phi$ is called a subsystem if $\Psi$ is itself a root system. For each root $\alpha\in\Phi$, the set $\Phi_{\alpha}:=\{\beta\in\Phi:B(\alpha,\beta)=0\}$ is automatically a subsystem.
2.2. Positive and simple systems
A subset $\Delta$ of a root system $\Phi$ is called a simple system if $\Delta$ is a vector space basis for $V$ and every root is a linear combination of $\Delta$ with coefficients of like sign. Given such a system $\Delta$, we say a root $\alpha=\sum_{i\in\Gamma}c_{i}\alpha_{i}$ is positive (with respect to $\Delta$) if $c_{i}\geq 0$ for all $i$, and we call $\alpha$ negative if $c_{i}\leq 0$ for all $i$. The sets of positive and negative roots are denoted by $\Phi_{\Delta}^{+}$ and $\Phi_{\Delta}^{-}$, and they are setwise negations of each other. For each root $\alpha$, the integer $\text{\rm ht}(\alpha)=\sum_{i\in\Gamma}c_{i}$ is called the height of $\alpha$.
Each simple system $\Delta$ also gives rise to a partial order $\leq_{\Delta}$ on $\Phi$, which is defined by the condition that $\alpha\leq\beta$ if and only if $\beta-\alpha$ is a nonnegative linear combination of $\Delta$. With respect to this partial order, $\Phi$ has a unique maximal element, $\theta_{\Delta}$, called the highest root.
The set of simple roots $\Pi$ is an example of a simple system, and the corresponding set of positive roots is an example of a positive system. Recall that each positive system $P$ contains a unique simple system, which is the set $\Delta_{P}$ of all elements in $P$ that cannot be expressed as positive linear combinations of other elements of $P$. The maps $\Delta\mapsto\Phi_{\Delta}^{+}$ and $P\mapsto\Delta_{P}$ are mutually inverse bijections between the sets of simple systems and positive systems in $\Phi$ [26, Theorem 1.3]. The simple systems of $\Phi$ are conjugate to each other under the action of $W$, as are the positive systems.
From now on, we choose $\Pi$ as the default simple system of $\Phi$ and choose $\Phi^{+}_{\Pi}$ as the default positive system. For each notion defined above relative to a general simple system $\Delta$, an omission of the subscript in the correponding notation will indicate that $\Pi$ is chosen as $\Delta$. For example, the set of positive roots with respect to $\Pi$ will be denoted by $\Phi^{+}$.
For any subsystem $\Psi$ of $\Phi$, the set $\Psi^{+}:=\Psi\cap\Phi^{+}$ is automatically a positive system of $\Psi$. We call $\Psi^{+}$ the induced positive system of $\Psi$ with respect to $\Phi^{+}$. We call the corresponding simple system of $\Psi$ the induced simple system of $\Psi$.
If $\Psi$ is a subsystem of the form $\Phi_{\alpha}$, i.e., if $\Psi$ is the subsystem of roots orthogonal to a root $\alpha$, then we denote the induced simple system by $\Pi_{\alpha}$. The elements of $\Pi_{\alpha}$ are thus the positive roots orthogonal to $\alpha$ that cannot be expressed as a nonnegative linear combination of other positive roots orthogonal to $\alpha$. Note that the elements of $\Pi_{\alpha}$ may not all lie in $\Pi$, but every simple root $\alpha_{i}\in\Psi$ that is orthogonal to $\alpha$ will lie in $\Pi$. We denote the Weyl group corresponding to $\Pi_{\alpha}$ by $W_{\alpha}$, so that $\Phi_{\alpha}$ is the root system of $W_{\alpha}$. It is known that $W_{\alpha}$ is the stabilizer of $\alpha$ under the action of $W$, and $W_{\alpha}$ is a direct product of irreducible simply laced Weyl groups [7, 1]. We will recall the known Dynkin type of the groups $W_{\alpha}$ in the next subsection.
Example 2.2.1.
Let $W$ be the Weyl group of type $D_{5}$ and let $\alpha_{2}$ be the simple root of $W$ under the labelling used in Figure 1. The induced simple system of $\Phi_{\alpha_{2}}$ is given by the set
$\Pi_{\alpha_{2}}=\{\beta_{1}=\alpha_{2}+2\alpha_{3}+\alpha_{4}+\alpha_{5}\}\ \cup\ \{\beta_{2}=\alpha_{1}+\alpha_{2}+\alpha_{3},\beta_{3}=\alpha_{4},\beta_{4}=\alpha_{5}\}.$
The corresponding Weyl group is the direct product $W(A_{1})\times W(A_{3})$ of the Weyl groups of types $A_{1}$ and $A_{3}$, generated respectively by the sets $\{s_{\beta_{1}}\}$ and $\{s_{\beta_{2}},s_{\beta_{3}},s_{\beta_{4}}\}$.
2.3. Explicit constructions
We now recall well-known explicit realizations of the root systems of types $A,D$ and $E$ in coordinate systems. Let $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}$ be an orthonormal basis for ${\mathbb{R}}^{n}$. The vectors $\{\varepsilon_{i}-\varepsilon_{j}:1\leq i\neq j\leq n\}$ form a root system of type $A_{n-1}$. The simple roots $\Pi=\{\alpha_{1},\alpha_{2},\ldots,\alpha_{n-1}\}$ are given by $\alpha_{i}=\varepsilon_{i}-\varepsilon_{i+1}$. A root $\varepsilon_{i}-\varepsilon_{j}$ is positive if $i<j$ and negative if $i>j$. The highest root (with respect to $\Pi$) is $\varepsilon_{1}-\varepsilon_{n}$. The bilinear form $B$ is the usual dot product, and two roots are orthogonal if and only if they have disjoint support. The Weyl group is isomorphic to $S_{n}$ and acts by permuting the orthonormal basis $\varepsilon_{1},..,\varepsilon_{n}$. The stabilizer $W_{\alpha}$ of each root $\alpha$ is a Weyl group of type $A_{n-3}$, which is trivial if $n\leq 3$.
The vectors {±$\{\pm\varepsilon_{i}\pm\varepsilon_{j}:1\leq i<j\leq n\}$ form a root system of type $D_{n}$. The simple roots $\{\alpha_{1},\alpha_{2},\ldots,\alpha_{n}\}$ are given by $\alpha_{i}=\varepsilon_{i}-\varepsilon_{i+1}$ for $i<n$, and $\alpha_{n}=\varepsilon_{i}+\varepsilon_{i+1}$. If $i<j$, then the roots $\varepsilon_{i}\pm\varepsilon_{j}$ are positive, and the roots $-\varepsilon_{i}\pm\varepsilon_{j}$ are negative. The highest root is $\varepsilon_{1}+\varepsilon_{2}$. The bilinear form $B$ is the usual dot product, and two roots $\alpha$ and $\beta$ are orthogonal if and only if either (a) $\alpha$ and $\beta$ have disjoint support or (b) $\alpha$ and $\beta$ have the same support and $\alpha\neq\pm\beta$. The Weyl group has order $2^{n-1}n!$, and acts by signed permutations of the orthonormal basis, with the restriction that each element effects an even number of sign changes [26, Section 2.10]. The stabilizer of a root is a Weyl group of type $A_{1}+D_{n-2}$, meaning $W(A_{1})\times W(D_{n-2})$, where we interpret $D_{3}$ as $A_{3}$ and $D_{2}$ as $A_{1}+A_{1}$. There is a well-known homomorphism $\phi$ from $W(D_{2n})$ to the symmetric group $S_{2n}$ resulting from forgetting the signs in a signed permutation; it is given by the following formula:
(2.3.1) $\phi(s_{i})=\begin{cases}(i,i+1)&\text{if\ }i<2n\\ (2n-1,2n)&\text{otherwise.}\end{cases}$
Let $\varepsilon_{0},\varepsilon_{1},\ldots,\varepsilon_{7}$ be an orthonormal basis of ${\mathbb{R}}^{8}$. The root system of type $E_{7}$ may be regarded as a subset of ${\mathbb{R}}^{8}$ as follows. There are 56 roots of the form ±$\pm 2(\varepsilon_{i}-\varepsilon_{j})$ where $0\leq i\neq j\leq 7$, and there are 70 roots of the form $\sum_{i=0}^{7}\pm\varepsilon_{i}$, where the signs are chosen so that there are four $+$ and four --. The simple roots are $\alpha_{1},\alpha_{2},\ldots,\alpha_{7}$, where $\alpha_{i}=2(\varepsilon_{i}-\varepsilon_{i+1})$ for $1\leq i\leq 6$, and
$\alpha_{7}=-\varepsilon_{0}-\varepsilon_{1}-\varepsilon_{2}-\varepsilon_{3}+\varepsilon_{4}+\varepsilon_{5}+\varepsilon_{6}+\varepsilon_{7}.$
A root of the form $2(\varepsilon_{i}-\varepsilon_{j})$ is positive if $0<i<j$ or $j=0$, and negative otherwise. A root of the form $\sum_{i=0}^{7}\pm\varepsilon_{i}$ is positive if and only if $\varepsilon_{0}$ occurs with negative coefficient. The highest root is $2(\varepsilon_{1}-\varepsilon_{0})$. The bilinear form $B$ is $1/4$ of the usual dot product. The stabilizer of a root is a Weyl group of type $D_{6}$. We call the coordinates $\varepsilon_{i}$ Fano coordinates because they are particularly compatible with the combinatorics of the Fano plane; this will be important in Section 6.2.
Let $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{8}$ be an orthonormal basis of ${\mathbb{R}}^{8}$. The root system of type $E_{8}$ may be regarded as a subset of ${\mathbb{R}}^{8}$ as follows. There are 112 roots of the form ±$\pm 2(\varepsilon_{i}\pm\varepsilon_{j})$ where $1\leq i\neq j\leq 8$, and there are 128 roots of the form $\sum_{i=1}^{8}\pm\varepsilon_{i}$, where the signs are chosen so that the total number of -- is even. The simple roots are $\alpha_{1},\alpha_{2},\ldots,\alpha_{8}$, where
$\alpha_{1}=\varepsilon_{1}-\varepsilon_{2}-\varepsilon_{3}-\varepsilon_{4}-\varepsilon_{5}-\varepsilon_{6}-\varepsilon_{7}+\varepsilon_{8},$
$\alpha_{2}=\varepsilon_{1}+\varepsilon_{2}$, and $\alpha_{i}=\varepsilon_{i-1}-\varepsilon_{i-2}$ for all $3\leq i\leq 8$. If $k$ is the largest integer such that $\varepsilon_{k}$ appears in $\alpha$ with nonzero coefficient $c$, then $\alpha$ is positive if and only if $c>0$. The highest root is $2(\varepsilon_{7}+\varepsilon_{8})$. The bilinear form $B$ is $1/4$ of the usual dot product. The stabilizer of a root is a Weyl group of type $E_{7}$. We call the coordinates $\varepsilon_{i}$ the standard coordinates for $E_{8}$.
3. Combinatorics of coplanar quadruples
A matching of $[2n]:=\{1,2,\ldots,2n\}$ is a collection of pairwise disjoint size-2 subsets, or 2-blocks, of $[2n]$. The matching is perfect if the union of the 2-blocks is the whole of $[2n]$. If $1\leq a<b<c<d\leq 2n$, then a crossing is a subset of the matching of the form $\{\{a,c\},\ \{b,d\}\}$, a nesting is a subset of the form $\{\{a,d\},\ \{b,c\}\}$, and an alignment is a subset of the form $\{\{a,b\},\ \{c,d\}\}$. For convenience, we will often denote each 2-block $\{a,b\}$ in a matching simply by $ab$ from now on.
In this section, we generalize crossings, nestings, and alignments to the context of orthogonal sets of roots (Definition 3.2.1, Remark 3.2.3). The resulting generalizations, collectively called coplanar quadruples, will be very useful for understanding the action of $W$ on $n$-roots. We develop some key properties of coplanar quadruples in Theorem 3.2.2, which is the main result of Section 3. We also show that crossings, nestings, and alignments can be distinguished from each other based on the heights of the roots that they contain (Proposition 3.2.5), and we give a precise description of the ways in which two coplanar quadruples can overlap (Proposition 3.3.6).
3.1. Coplanar quadruples
We gather a few facts about $n$-roots and define coplanar quadruples in this subsection. The following two results on maximal orthogonal sets of roots are well known, but we include proofs for ease of reference.
Lemma 3.1.1.
Let $W$ be a Weyl group of type $D_{n}$ for $n$ even. Suppose that $n=2k\geq 4$.
  • (i)
    Every maximal orthogonal set of roots is of the form
[…]
    ±$\pm(\varepsilon_{\sigma(n-1)}+\varepsilon_{\sigma(n)}),\ \pm(\varepsilon_{\sigma(n-1)}-\varepsilon_{\sigma(n)})\}$
    for some permutation $\sigma\in S_{n}$, where the signs are chosen independently.
  • (ii)
    Every maximal orthogonal set of positive roots is of the form
    $\{\varepsilon_{i_{1}}-\varepsilon_{j_{1}},\ \varepsilon_{i_{1}}+\varepsilon_{j_{1}},\ \varepsilon_{i_{2}}-\varepsilon_{j_{2}},\ \varepsilon_{i_{2}}+\varepsilon_{j_{2}},\ \ldots,\varepsilon_{i_{k}}-\varepsilon_{j_{k}},\ \varepsilon_{i_{k}}+\varepsilon_{j_{k}}\},$
    and these sets are in bijection with perfect matchings $\{\{i_{1},j_{1}\},\ \{i_{2},j_{2}\},\ \ldots,\{i_{k},j_{k}\}\}$ of the set $[n]$ that satisfy $i_{r}<j_{r}$ for all $1\leq r\leq k$.
Proof.
Let $R$ be a maximal orthogonal set of roots. By symmetry, we may reduce to the case where all the roots in $R$ are positive. If $R$ contains the root $\varepsilon_{i}\pm\varepsilon_{j}$, then $R$ must also contain the root $\varepsilon_{i}\mp\varepsilon_{j}$, because otherwise $R\cup\{\varepsilon_{i}\mp\varepsilon_{j}\}$ would be a set of orthogonal roots that was larger than $R$. It follows that $R$ consists of $n/2$ pairs of roots such that each pair has the same support, and roots from distinct pairs have disjoint supports. This completes the proof of (i).
Part (ii) follows from (i) and the fact that if $1\leq i<j\leq n$, then the roots $\varepsilon_{i}\pm\varepsilon_{j}$ are positive. ∎
It follows from Lemma 3.1.1 (ii) that a maximal orthogonal set of positive roots in type $D_{n}$ (for $n$ even) contains the root $\varepsilon_{i}-\varepsilon_{j}$ if and only if it contains the root $\varepsilon_{i}+\varepsilon_{j}$. We will call such a pair of roots $\{\varepsilon_{i}\pm\varepsilon_{j}\}$ a collinear pair of roots.
Lemma 3.1.2.
If $W$ is a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ with $n\geq 4$ even, then $W$ acts transitively on the set $\mathcal{M}(W)$ of unordered maximal sets of orthogonal roots of $W$.
Proof.
Recall from Section 2 that the group $W(D_{n})$ can be regarded as the group of signed permutations of $n$ objects in which there is an even number of sign changes. Such a group acts transitively on the set described in Lemma 3.1.1 (i).
Now suppose $W$ has type $E_{7}$, and let $\alpha$ be a root of $W$. Then by Section 2, the stabilizer $W_{\alpha}$ is a Weyl group of type $D_{6}$ whose root system is the set $\Phi_{\alpha}$ of roots that are orthogonal to $\alpha$. Since $W_{\alpha}$ acts transitively on $\Phi_{\alpha}$, it follows that there is a well-defined bijection $[R]\mapsto[R\cup\{\alpha\}]$ from the set of $W_{\alpha}$-orbits on $\mathcal{M}(W_{\alpha})$ to the set of $W$-orbits on $\mathcal{M}(W)$, where the orbit $[R]$ of every 6-tuple $R\in\mathcal{M}(W_{\alpha})$ is sent to the orbit $[R\cup\{\alpha\}]$ of the 7-tuple $R\cup\{\alpha\}$. It then follows that $W$ acts transitively on $\mathcal{M}(W)$, as desired.
Finally, if $W$ has type $E_{8}$, then for each root $\alpha$ of $W$ the stablizer $W_{\alpha}$ is of type $E_{7}$. A similar argument to the one above shows that $W$ acts transitively on $\mathcal{M}(W)$ because $W_{\alpha}$ acts transitively on $\mathcal{M}(W_{\alpha})$. ∎
We are ready to define coplanar quadruples. The following proposition offers multiple equivalent characterizations of them.
Proposition 3.1.3.
Let $Q=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ be a set of four mutually orthogonal roots for a simply laced Weyl group $W$ with root system $\Phi$, and let $\gamma=(\beta_{1}+\beta_{2}+\beta_{3}+\beta_{4})/2$. The following are equivalent:
  • (i)
    $\gamma$ is a root (i.e., the elements of $Q$ sum to twice a root);
  • (ii)
    $Q$ is contained in a subsystem $\Psi$ of type $D_{4}$;
  • (iii)
    there is a unique subsystem $\Psi$ of type $D_{4}$ such that $(Q\cup\{\gamma\})\subset\Psi\subseteq\Phi$, and we have
    $\Psi=\left\{\pm\beta_{1},\ \pm\beta_{2},\ \pm\beta_{3},\ \pm\beta_{4},\ (\pm\beta_{1}\pm\beta_{2}\pm\beta_{3}\pm\beta_{4})/2\right\},$
    where all the signs are chosen independently.
Proof.
We first prove that (i) implies (iii). Assume that $\gamma$ is a root. Any root subsystem containing $Q\cup\{\gamma\}$ also contains $s_{\beta_{i}}(\beta_{i})=-\beta_{i}$ for each $i$, as well as all roots of the form
$s_{\beta_{1}}^{\epsilon_{1}}s_{\beta_{2}}^{\epsilon_{2}}s_{\beta_{3}}^{\epsilon_{3}}s_{\beta_{4}}^{\epsilon_{4}}(\gamma),$
where we have $\epsilon_{i}\in\{0,1\}$ for all $i$. The 16 roots listed above can also be expressed as
(±$(\pm\beta_{1}\pm\beta_{2}\pm\beta_{3}\pm\beta_{4})/2.$
We have constructed all 24 roots in the set $\Psi$ listed in the statement, and this is the cardinality of a root system of type $D_{4}$. To prove (iii) it now suffices to show that $\Psi$ is a root system of type $D_{4}$. Because the elements of $Q$ are orthogonal vectors of the same length, we may choose Euclidean coordinates $\beta_{1}=\varepsilon_{1}-\varepsilon_{2}$, $\beta_{2}=\varepsilon_{1}+\varepsilon_{2}$, $\beta_{3}=\varepsilon_{3}-\varepsilon_{4}$, and $\beta_{4}=\varepsilon_{3}+\varepsilon_{4}$. With respect to these coordinates, we have
$\Psi=\{\pm\varepsilon_{i}\pm\varepsilon_{j}:1\leq i<j\leq 4\},$
which indeed forms a root system of type $D_{4}$, as desired.
It is immediate that (iii) implies (ii).
In the usual notation for the simple roots of type $D_{4}$, the orthogonal roots $\alpha_{1}$, $\alpha_{3}$, $\alpha_{4}$, and $\alpha_{1}+2\alpha_{2}+\alpha_{3}+\alpha_{4}$ sum to $2\alpha$, where $\alpha$ is the root $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$. Lemma 3.1.2 applied to a root system of type $D_{4}$ then implies that the sum of every orthogonal quadruple of roots in a root system of type $D_{4}$ is equal to $2\alpha^{\prime}$ for some root $\alpha^{\prime}$. It follows that (ii) implies (i), which completes the proof. ∎
Definition 3.1.4.
A set $Q$ of four mutually orthogonal roots for a simply laced Weyl group is called a coplanar quadruple if it satisfies the equivalent conditions of Proposition 3.1.3. In this case, we call the set $\Psi$ from Proposition 3.1.3 the $D_{4}$-subsystem associated to $Q$.
Coplanar quadruples can be described explicitly in coordinates in type $D$:
Lemma 3.1.5.
Let $W$ be a Weyl group of type $D_{n}$ for $n$ even and $n\geq 4$. Then four positive roots of $W$ form a coplanar quadruple if and only if they consist of two collinear pairs of roots, i.e., if and only if they are of the form $\varepsilon_{i}+\varepsilon_{j},\varepsilon_{i}-\varepsilon_{j},\varepsilon_{k}+\varepsilon_{l},\varepsilon_{k}-\varepsilon_{l}$ for four distinct indices $i,j,k,l$ where $i<j$ and $k<l$.
Remark 3.1.6.
In the setting of Lemma 3.1.5, we may naturally identify the coplanar quadruple $\{\varepsilon_{i}\pm\varepsilon_{j},\varepsilon_{k}\pm\varepsilon_{l}\}$ with the matching $\{ij,kl\}$ of the set $\{i,j,k,l\}$.
Remark 3.1.7.
Recall that reflections in $W(D_{n})$ act on the reflection representation as signed permutations, with $s_{\alpha}(\varepsilon_{i})=\varepsilon_{j}$ if $\alpha=\varepsilon_{i}-\varepsilon_{j}$ and $s_{\alpha}(\varepsilon_{i})=-\varepsilon_{j}$ if $\alpha=\varepsilon_{i}+\varepsilon_{j}$. It follows that $W(D_{n})$ acts on the set $\{\varepsilon_{1}^{2},\cdots,\varepsilon_{n}^{2}\}$ as ordinary premutations, with $s_{\alpha}(\varepsilon_{i}^{2})=\varepsilon_{j}^{2}$ for all distinct $i,j$. When $n$ is even, it then follows from Lemmas 3.1.1 and 3.1.5 that the action of $W(D_{n})$ on $n$-roots factors through the map $\phi:W(D_{n})\rightarrow S_{n}$ from Equation (2.3.1) to induce an action of $S_{n}$ on $n$-roots. In particular, each reflection $r\in W(D_{n})$ acts in the same way as $\phi(r)$ on every $n$-root of $W(D_{n})$.
Proof of Lemma 3.1.5.
The “if” implication holds since the four roots in the given form sum to twice the positive root $\varepsilon_{i}+\varepsilon_{k}$. To prove the “only if” implication, let $Q=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ be a coplanar quadruple. Recall from Section 2 that two roots in type $D_{n}$ are orthogonal if and only if they have the same or disjoint support. It follows that if no two roots in $Q$ have the same support, then the supports of $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ contain a total of eight distinct coordinates $\varepsilon_{i}$, in which case the sum $\gamma=\beta_{1}+\beta_{2}+\beta_{3}+\beta_{4}$ cannot be twice a root. We may therefore assume, without loss of generality, that $\{\beta_{1},\beta_{2}\}=\{\varepsilon_{i}\pm\varepsilon_{j}\}$ for some $i<j$. This implies $\beta_{1}+\beta_{2}=2\varepsilon_{i}$. The condition that $Q$ is an orthogonal set summing to twice a root then forces us to have $\{\beta_{3},\beta_{4}\}=\{\varepsilon_{k}\pm\varepsilon_{l}\}$ for some elements $k,l$ distinct from $i$ and $j$ with $k<l$. ∎
The next proposition shows that the action of $W$ on $n$-roots is local to coplanar quadruples in the following sense: whenever a reflection in $W$ does not fix a maximal orthogonal set $R$ of roots, it must change exactly four elements of $R$ that form a coplanar quadruple, and it changes these four elements to another coplanar quadruple with the same associated $D_{4}$-subsystem.
Proposition 3.1.8.
Let $W$ be a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even, let $\alpha$ be a root, and let $R$ be a maximal set of orthogonal positive roots. Suppose that neither $\alpha$ nor $-\alpha$ is an element of $R$.
  • (i)
    The root $\alpha$ is orthogonal to all but precisely four elements $Q=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ of $R$. The elements of $Q$ form a coplanar quadruple, and we have $2\alpha=\pm\beta_{1}\pm\beta_{2}\pm\beta_{3}\pm\beta_{4}$ for suitable choices of signs.
  • (ii)
    Let $\Psi$ be the $D_{4}$ subsystem associated to $Q$. Then we have $\alpha\in\Psi$, and the set $s_{\alpha}(Q)=\{s_{\alpha}(\beta_{i}):1\leq i\leq 4\}$ is also a coplanar quadruple whose associated $D_{4}$-subsystem is $\Psi$.
Proof.
To prove (i), it suffices by Lemma 3.1.2 to do so for a fixed $R$. Suppose first that $W$ is of type $D_{n}$, and choose
$R=\{\varepsilon_{1}+\varepsilon_{2},\ \varepsilon_{1}-\varepsilon_{2},\ \varepsilon_{3}+\varepsilon_{4},\ \varepsilon_{3}-\varepsilon_{4},\ \ldots,\ \varepsilon_{n-1}+\varepsilon_{n},\ \varepsilon_{n-1}-\varepsilon_{n}\}.$
The root $\alpha$ must be of the form ±$\pm\varepsilon_{i}\pm\varepsilon_{j}$, where $i$ and $j$ come from different parts of the partition $\{\{1,2\},\{3,4\},\ldots,\{n-1,n\}\}$. It follows that the support of $\alpha$ has one element in common with the support of each of precisely four elements of $R$ making up two collinear pairs, and that $\alpha$ is orthogonal to all the other elements of $R$. Furthermore, the roots $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ that are not orthogonal to $\alpha$ are of the form ±$\pm\varepsilon_{h}\pm\varepsilon_{i}$ and ±$\pm\varepsilon_{j}\pm\varepsilon_{k}$, where |$|\{h,i,j,k\}|=4$. It follows that $2\alpha$ can be written in the form ±$\pm\beta_{1}\pm\beta_{2}\pm\beta_{3}\pm\beta_{4}$ for suitable choices of signs.
Next, suppose that $W$ has type $E_{8}$, and choose
[…]
Now suppose that $W$ has type $E_{7}$. By Section 2.3, we may identify the root system of $W$ with the set of roots orthogonal to the highest root $\theta$ in the root system of type $E_{8}$, so that $R\cup\{\theta\}$ is a maximal set of roots in type $E_{8}$. By the above paragraph, $\alpha$ is orthogonal to four of the roots in $R\cup\{\theta\}$, but one of these roots is $\theta$ itself. It follows that $\alpha$ is orthogonal to three elements of $R$, and that $2\alpha$ can be expressed as a signed sum of the other four elements of $R$, as required. This completes the proof of (i).
It follows from (i) that we have $\alpha\in\Psi$, because condition (i) implies condition (iii) in Proposition 3.1.3. The element $\gamma=(\beta_{1}+\beta_{2}+\beta_{3}+\beta_{4})/2$ is a root because $Q$ is a coplanar quadruple, and the set $s_{\alpha}(Q)$ is a coplanar quadruple because its elements sum to twice $s_{\alpha}(\gamma)$, which is a root because $\gamma$ is. We also have $s_{\alpha}(Q)\subseteq\Psi$ because both $\alpha$ and $Q$ are in $\Psi$. This implies that $\Psi$ must be the $D_{4}$-subsystem associated to $s_{\alpha}(Q)$, proving (ii). ∎
3.2. Crossings, nestings, and alignments
We examine coplanar quadruples more closely in this subsection and classify them into three types, namely, crossings, nestings, and alignments. As we will explain in Remark 3.2.3, our terminology comes from the theory of matchings, but the following definition makes sense for all the simply laced root systems considered in this paper.
Definition 3.2.1.
Let $Q=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ be a coplanar quadruple of positive orthogonal roots, let $\Psi$ be the $D_{4}$-subsystem associated to $Q$, let $\leq$ be the partial order on $\Psi$ relative to the induced simple roots of $\Psi$, and let $\gamma$ be the root $(\beta_{1}+\beta_{2}+\beta_{3}+\beta_{4})/2$. We say that $Q$ is
  • (i)
    a crossing if $\beta_{i}\leq\gamma$ for all $i$ and $Q$ contains the unique $\leq$-maximal element of $Q\cup(-s_{\gamma}(Q))$;
  • (ii)
    a nesting if $\beta_{i}\leq\gamma$ for all $i$ and $Q$ contains the unique $\leq$-minimal element of $Q\cup(-s_{\gamma}(Q))$;
  • (iii)
    an alignment otherwise.
We also call each crossing, nesting, and alignment a feature of type A, type C, and type N, respectively.
Theorem 3.2.2.
Let $\Phi$ be a root system for a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even. Let $Q$ be a coplanar quadruple of positive roots of $\Phi$, let $\Psi$ be the associated $D_{4}$-subsystem, and let $\Psi^{+}=\Psi\cap\Phi^{+}$ be the induced positive system of $\Psi$.
  • (i)
    The set $\Psi^{+}$ contains precisely three distinct quadruples of mutually orthogonal roots. These quadruples are pairwise disjoint and partition $\Psi^{+}$.
  • (ii)
    The three quadruples of orthogonal roots in $\Psi^{+}$ are all coplanar. Among them there is exactly one crossing, $\Psi^{+}_{C}$, exactly one nesting, $\Psi^{+}_{N}$, and exactly one alignment, $\Psi^{+}_{A}$. In particular, the quadruple $Q$ cannot be both a crossing and a nesting, and the three conditions in Definition 3.2.1 are mutually exclusive.
  • (iii)
    Each quadruple in $\{\Psi^{+}_{C},\Psi^{+}_{N},\Psi^{+}_{A}\}$ uniquely determines both of the other two.
  • (iv)
    If $R$ is a set of mutually orthogonal roots that is disjoint from $\Psi$, then either each of the three sets $\{R\cup\Psi^{+}_{C},\ R\cup\Psi^{+}_{N},\ R\cup\Psi^{+}_{A}\}$ consists of mutually orthogonal roots, or none of them does.
  • (v)
    The crossing $\Phi^{+}_{C}$ contains no root from the induced simple system of $\Psi$.
  • (vi)
    For each $x\in\{C,N,A\}$, let $\gamma_{x}$ be the product of the roots in $\Psi^{+}_{x}$. Then we have $\sigma(\gamma_{A})<\sigma(\gamma_{N})=\sigma(\gamma_{C})$ and $\gamma_{C}=\gamma_{N}+\gamma_{A}.$ Moreover, if $\alpha$ is any component in one of the three $4$-roots $\gamma_{C},\gamma_{N}$ and $\gamma_{A}$, then the reflection $s_{\alpha}$ sends the other two $4$-roots to each other; for example, if $\alpha\in\Psi^{+}_{C}$ then $s_{\alpha}$ sends $\gamma_{N}$ and $\gamma_{A}$ to each other.
Proof.
By Section 2.3, the roots orthogonal to a given root in type $D_{4}$ form a subsystem of type $3A_{1}$. Therefore each positive root lies in a unique quadruple of mutually orthogonal positive roots, which proves (i).
Let $\{\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}\}$ be the induced simple roots of $\Psi$, with $\alpha_{2}$ corresponding to the branch node in the Dynkin diagram. Then the three quadruples from (i) are given by
$\Psi^{+}_{1}$ =$=\{\alpha_{1}+\alpha_{2},\ \alpha_{2}+\alpha_{3},\ \alpha_{2}+\alpha_{4},\ \alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\},$
$\Psi^{+}_{2}$ =$=\{\alpha_{2},\ \alpha_{1}+\alpha_{2}+\alpha_{3},\ \alpha_{1}+\alpha_{2}+\alpha_{4},\ \alpha_{2}+\alpha_{3}+\alpha_{4}\},\quad\text{\ and\ }$
$\Psi^{+}_{3}$ =$=\{\alpha_{1},\ \alpha_{3},\ \alpha_{4},\ \alpha_{1}+2\alpha_{2}+\alpha_{3}+\alpha_{4}\}.$
The roots in $\Psi^{+}_{3}$ add up to twice the root $\alpha=\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$. The root $\alpha$ is strictly lower in the $\leq$ order than one of the roots in $\Psi^{+}_{3}$, therefore $\Psi^{+}_{3}$ is an alignment by Definition 3.2.1. The roots in $\Psi^{+}_{1}$ and $\Psi^{+}_{2}$ both add up to $2\theta$, where $\theta=\alpha_{1}+2\alpha_{2}+\alpha_{3}+\alpha_{4}$. The root $\theta$ is strictly higher in the $\leq$ order than each element of $\Psi^{+}_{1}$ and $\Psi^{+}_{2}$. Note that we have $s_{\theta}(\Psi^{+}_{1})=-\Psi^{+}_{2}$ and $s_{\theta}(\Psi^{+}_{2})=-\Psi^{+}_{1}$. Furthermore, the roots $\alpha_{2}$ and $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$ are the unique $\leq$-minimal and unique $\leq$-maximal elements of $\Psi^{+}_{1}\ \dot{\cup}\ \Psi^{+}_{2}$, respectively. This implies that $\Psi^{+}_{1}$ is a crossing and that $\Psi^{+}_{2}$ is a nesting, and it also follows that none of $\Psi^{+}_{1},\Psi^{+}_{2}$ and $\Psi^{+}_{3}$ is both a crossing and a nesting. The quadruple $Q$ must be one of $\Psi^{+}_{1},\Psi^{+}_{2}$ and $\Psi^{+}_{3}$, and (ii) follows.
Part (iii) follows from (ii), since each of $\Psi^{+}_{C},\Psi^{+}_{N}$ and $\Psi^{+}_{A}$ uniquely determines $\Psi^{+}$ as its associated $D_{4}$-subsystem by Proposition 3.1.3 (iii).
Part (iv) holds as each of the quadruples in $\{\Psi^{+}_{C},\Psi^{+}_{N},\Psi^{+}_{A}\}$ is a basis for the span of $\Psi^{+}$, so that any root that is orthogonal to every element of one quadruple is also orthogonal to every element of each of the other quadruples.
Finally, the claims in (v) and (vi) can all be verified by inspection or direct computation based on the description of $\Psi^{+}_{C}=\Psi^{+}_{1},\Psi^{+}_{N}=\Psi^{+}_{2}$ and $\Psi^{+}_{A}=\Psi^{+}_{3}$. For the equation $\gamma_{C}=\gamma_{N}+\gamma_{A}$ and the assertion about $s_{\alpha}$ in (vi), one can alternatively prove them using the usual realizations of the root system and Weyl group of type $D_{4}$, where the simple roots are $\alpha_{1}=\varepsilon_{1}-\varepsilon_{2}$, $\alpha_{2}=\varepsilon_{2}-\varepsilon_{3}$, $\alpha_{3}=\varepsilon_{3}-\varepsilon_{4}$ and $\alpha_{4}=\varepsilon_{3}+\varepsilon_{4}$ and the group $W$ acts as signed permutations. Under this realization, we have
(3.2.1) $\gamma_{C}=(\varepsilon_{1}^{2}-\varepsilon_{3}^{2})(\varepsilon_{2}^{2}-\varepsilon_{4}^{2}),\;\gamma_{N}=(\varepsilon_{1}^{2}-\varepsilon_{4}^{2})(\varepsilon_{2}^{2}-\varepsilon_{3}^{2}),\;\text{and}\;\gamma_{A}=(\varepsilon_{1}^{2}-\varepsilon_{2}^{2})(\varepsilon_{3}^{2}-\varepsilon_{4}^{2}),$
and the equation $\gamma_{C}=\gamma_{N}+\gamma_{A}$ follows as the terms expressed in coordinates satisfy the Ptolemy relation
$(A-C)(B-D)=(A-D)(B-C)+(A-B)(C-D).\qed$
Remark 3.2.3.
In the setting of Theorem 3.2.2, the coordinate forms of the $4$-roots $\gamma_{C},\gamma_{N},\gamma_{A}$ given in Equation (3.2.1) correspond via the bijection of Lemma 3.1.1 (ii) to the perfect matchings of the set $[4]$ given by the crossing $m_{C}=\{\{1,3\},\{2,4\}\}$, the nesting $m_{N}=\{\{1,4\},\{2,3\}\},$ and the alignment $m_{A}=\{\{1,2\},\{3,4\}\}$, respectively. Definition 3.2.1 generalizes the notion of crossings, nestings, and alignments in the sense that a coplanar quadruple in the sense of the definition is a crossing, nesting, or alignment if and only if the corresponding perfect matching of the set $[4]$ is a crossing, nesting, or alignment, respectively, in the context of matchings.
Remark 3.2.4.
Let $R$ be a maximal orthogonal set of positive roots and let $\alpha$ be a positive root not in $R$. Proposition 3.1.8 shows that the reflection $s_{\alpha}$ changes precisely four elements in $R$ which form a coplanar quadruple $Q$, and that $Q^{\prime}=s_{\alpha}(Q)$ is another coplanar quadruple with the same associated $D_{4}$-subsystem as $Q$. Theorem 3.2.2 (vi) reveals more about $\alpha,Q$ and $Q^{\prime}$: it shows that $Q$ and $Q^{\prime}$ are two distinct features from the set $\{\Psi^{+}_{C},\Psi^{+}_{N},\Psi^{+}_{A}\}$ inside the $D_{4}$-subsystem $\Psi$ associated to $Q$, while $\alpha$ is in the remaining feature. We will say that $s_{\alpha}$ moves $Q$ in this case; we will also say that $s_{\alpha}$ (or $\alpha$) moves an $X$ to a $Y$ and call $s_{\alpha}$ an XY move, where $X$ and $Y$ are the distinct types of $Q$ and $Q^{\prime}$, respectively. Note that knowledge of $Q$ and $Y$ is enough to determine $Q^{\prime}$ by Theorem 3.2.2 (iii), even if $\alpha$ is not known. Note also that Theorem 3.2.2 (vi) guarantees that $XY$ moves exist for any distinct elements $X$, $Y$ in $\{C,N,A\}$, so any two coplanar quadruples sharing the same $D_{4}$-subsystem can be connected, up to sign, by a reflection.
The next proposition shows how to distinguish crossings, nestings, and alignments from each other using only the heights of their components.
Proposition 3.2.5.
Let $W$ be a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even. Let $Q=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ be a coplanar quadruple of orthogonal positive roots, ordered so that we have $h_{1}\leq h_{2}\leq h_{3}\leq h_{4}$, where $h_{i}$ is the height $\text{\rm ht}(\beta_{i})$ of $\beta_{i}$ (with respect to the simple roots of $\Phi$) for each $i$.
  • (i)
    We have $h_{1}+h_{2}+h_{3}\neq h_{4}$ and $h_{2}+h_{3}\neq h_{1}+h_{4}$.
  • (ii)
    If $h_{1}+h_{2}+h_{3}<h_{4}$, then $Q$ is an alignment.
  • (iii)
    If $h_{1}+h_{2}+h_{3}>h_{4}$ and $h_{2}+h_{3}>h_{1}+h_{4}$, then $Q$ is a nesting. In this case, we also have $h_{1}<h_{2}$.
  • (iv)
    If $h_{1}+h_{2}+h_{3}>h_{4}$ and $h_{2}+h_{3}<h_{1}+h_{4}$, then $Q$ is a crossing. In this case, we also have $h_{1}>1$ and $h_{3}<h_{4}$.
Proof.
Let $\Psi$ be the type $D_{4}$ subsystem associated to $Q$, and let $\{\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}\}$ be the simple system of $\Psi$ induced by $\Phi^{+}$, with $\alpha_{2}$ corresponding to the branch node in the Dynkin diagram. Then as in the proof of Theorem 3.2.2, the set $\Psi^{+}$ decomposes into the union of the crossing $\Psi^{+}_{C}$, nesting $\Psi^{+}_{N}$, and alignment $\Psi^{+}_{A}$ given below:
$\Psi^{+}_{C}$ =$=\{\alpha_{1}+\alpha_{2},\ \alpha_{2}+\alpha_{3},\ \alpha_{2}+\alpha_{4},\ \alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\},$
$\Psi^{+}_{N}$ =$=\{\alpha_{2},\ \alpha_{1}+\alpha_{2}+\alpha_{3},\ \alpha_{1}+\alpha_{2}+\alpha_{4},\ \alpha_{2}+\alpha_{3}+\alpha_{4}\},\quad\text{\ and\ }$
$\Psi^{+}_{A}$ =$=\{\alpha_{1},\ \alpha_{3},\ \alpha_{4},\ \alpha_{1}+2\alpha_{2}+\alpha_{3}+\alpha_{4}\}.$
If $Q=\Psi^{+}_{A}$, then we have $\beta_{4}-\beta_{1}-\beta_{2}-\beta_{3}=2\alpha_{2}$, and $\beta_{4}+\beta_{1}-\beta_{2}-\beta_{3}=2\alpha_{2}+2\alpha_{i}$ for some $i\in\{1,3,4\}$.
If $Q=\Psi^{+}_{N}$, then we have $\beta_{1}<\beta_{j}$ for all $j\in\{2,3,4\}$. We also have $\beta_{4}-\beta_{1}-\beta_{2}-\beta_{3}=-2\alpha_{i}-2\alpha_{2}$ for some $i\in\{1,3,4\}$, and $\beta_{4}+\beta_{1}-\beta_{2}-\beta_{3}=-2\alpha_{i}$ for some $i\in\{1,3,4\}$.
If $Q=\Psi^{+}_{C}$, then we have $\beta_{4}>\beta_{j}$ for all $j\in\{1,2,3\}$. We also have $\beta_{4}-\beta_{1}-\beta_{2}-\beta_{3}=-2\alpha_{2}$ and $\beta_{4}+\beta_{1}-\beta_{2}-\beta_{3}=2\alpha_{i}$ for some $i\in\{1,3,4\}$. Furthermore, none of the $\beta_{i}$ is a simple root of $\Phi$, because none of the $\beta_{i}$ is even a simple root of $\Psi$.
All the claims in the proposition follow from the above observations. Note that the conditions in (ii), (iii), and (iv) are exclusive and exhaustive because of part (i). ∎
Remark 3.2.6.
It can be shown that if $\{h_{1},h_{2},\ldots,h_{n}\}$ is the multiset of heights of the components of an $n$-root $\beta$ of a Weyl group $W$ as in Proposition 3.2.5, then the number $\sum_{i=1}^{n}h_{i}^{2}$ depends only on $W$ and is independent of the choice of $\beta$.
3.3. Intersections of coplanar quadruples
Let $R$ be a maximal orthogonal set of roots of $W$. In this subsection, we first focus on type $E_{8}$ and show that coplanar quadruples in $R$ gives rise to a Steiner quadruple system in this type. We will then use this result to count coplanar quadruples in $R$ and deduce how coplanar quadruples in $R$ can overlap with each other, in the general case.
Definition 3.3.1.
A Steiner system $S(t,k,N)$ is a collection $B$ of $k$-element subsets of the set $[N]=\{1,2,\dots,N\}$ with the property that every $t$-element subset is contained in a unique element of $B$. The elements of $B$ are called blocks, and we write each block $\{a,b,c,\dots\}$ where $a<b<c<\dots$ as $abc\dots$. We call $B$ a Steiner triple system if $k=3$ and a Steiner quadruple system if $k=4$.
Remark 3.3.2.
It is well known that, up to isomorphism (by permutations), there is a unique Steiner triple system $S(2,3,7)$ and a unique Steiner quadruple system $S(3,4,8)$ [4]. The following 14 quadruples form an example of a Steiner quadruple system, and removing the element 8 from all the quadruples on the left results in a Steiner triple system.
1 2 4 8  3 5 6 7
2 3 5 8  1 4 6 7
3 4 6 8  1 2 5 7
4 5 7 8  1 2 3 6
1 5 6 8  2 3 4 7
2 6 7 8  1 3 4 5
1 3 7 8  2 4 5 6
Lemma 3.3.3.
Let $\beta_{1},\beta_{2},\beta_{3}$ be three mutually orthogonal positive roots of type $E_{8}$.
  • (i)
    There exists a unique positive root $\beta_{4}$ such that $\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ is a coplanar quadruple.
  • (ii)
    If $R$ is a maximal orthogonal set of positive roots of type $E_{8}$, and $\{\beta_{1},\beta_{2},\beta_{3}\}\subset R$, then we have $\beta_{4}\in R$.
Proof.
By [10, Lemma 11 (iii)] and its proof, if $W=W(E_{8})$, then $W$ acts transitively on ordered triples of orthogonal roots, and the set of roots orthogonal to three given mutually orthogonal roots is a root system of type $A_{1}+D_{4}$. Since the action of orthogonal triples of roots is transitive, it suffices to prove (i) for a fixed choice of $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$. If we choose $\beta_{1}=\alpha_{3}=2(\varepsilon_{2}-\varepsilon_{1})$, $\beta_{2}=\alpha_{5}=2(\varepsilon_{4}-\varepsilon_{3})$, and $\beta_{3}=\alpha_{7}=2(\varepsilon_{6}-\varepsilon_{5})$, then it follows from the explicit description of the root system in Section 2.3 that the only positive root that forms a coplanar quadruple with $\beta_{1},\beta_{2}$ and $\beta_{3}$ is the root $\beta_{4}=2(\varepsilon_{8}-\varepsilon_{7})$. This proves (i).
The uniqueness property of (i) proves that the set of coplanar quadruples corresponds to the $A_{1}$ summand of the $A_{1}+D_{4}$ subsystem. This implies that if $\beta$ is any positive root that is orthogonal to all of $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$, then either $\beta_{4}=\beta$, or $\beta_{4}$ is orthogonal to $\beta$. The maximality of $R$ in the statement of (ii) implies that $\beta_{4}$ cannot be orthogonal to all elements of $R$. It follows that $\beta_{4}\in R$, which proves (ii). ∎
Lemma 3.3.4.
Let $W$ be a Weyl group of type $E_{8}$, and let $R$ be a maximal orthogonal set of roots.
  • (i)
    The coplanar quadruples of $R$ endow $R$ with the structure of a Steiner quadruple system $S(3,4,8)$.
  • (ii)
    Any two coplanar quadruples of $R$ intersect in either 00, $2$, or $4$ elements.
Proof.
Part (i) is immediate from Lemma 3.3.3. To prove part (ii), we need to show that any two distinct quadruples from the Steiner quadruple system are either disjoint or overlap in precisely two elements. This can be proved by an exhaustive check, or by arguing as follows.
The quadruples in the left column of the table consist of the element 8 together with three points forming a line in the Fano plane (see Section 6.2). Any two such quadruples intersect in two elements: the element 8 and the unique point on the intersection of the two lines of the Fano plane. The general case follows by combining this observation with the fact that each quadruple in the right column is the complement of the corresponding quadruple in the left column. ∎
Corollary 3.3.5.
Let $W$ be a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even and $n=2k\geq 4$. Let $R$ be a maximal orthogonal set of positive roots. Then the number $M$ of coplanar quadruples contained in $R$ does not depend on $R$. We have $M=\binom{k}{2}$ if $W$ has type $D_{2k}$, $M=7$ if $W$ has type $E_{7}$, and $M=14$ if $W$ has type $E_{8}$.
Proof.
If $W$ has type $D_{2k}$, then $R$ determines a perfect matching of the set $\{1,2,\ldots,2k\}$ with $k$ blocks by Lemma 3.1.1, and the coplanar quadruples in $R$ correspond bijectively to pairs of these blocks by Remark 3.1.6. It follows that $M$ does not depend on $R$ and equals $\binom{k}{2}$.
In type $E_{8}$, the result follows from Lemma 3.3.4 (i).
If $W$ has type $E_{7}$, then as in the proof of Proposition 3.1.8, we may again identify the root system of $W$ with the set of roots orthogonal to the the highest root $\theta$ in the root system of type $E_{8}$. The set $R\cup\{\theta\}$ is then a maximal orthogonal set of roots in type $E_{8}$. The coplanar quadruples of $R$ are in bijection with the quadruples of $S(3,4,8)$ that exclude a fixed element, and there are $7$ such quadruples, so we are done. ∎
Proposition 3.3.6.
Let $W$ be a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even, and let $R$ be a maximal orthogonal subset of positive roots. If $Q_{1}$ and $Q_{2}$ are coplanar quadruples of roots contained in $R$, then we have the following.
  • (i)
    The intersection $Q_{1}\cap Q_{2}$ has size 0, 2, or 4.
  • (ii)
    If $|Q_{1}\cap Q_{2}|=2$, then there is a root subsystem $\Psi\subseteq\Phi$ of type $D_{6}$ that contains both $Q_{1}$ and $Q_{2}$. In this case, each of the sets $Q_{1}$, $Q_{2}$, $Q_{1}\cap Q_{2}$, and $Q_{1}\cup Q_{2}$ consists of collinear pairs of roots with respect to $\Psi$, and the symmetric difference $Q_{1}\ \Delta\ Q_{2}$ is also a coplanar quadruple.
Proof.
If $W$ has type $D_{n}$, then the assertions follow from Lemma 3.1.5.
Suppose that $W$ has type $E_{8}$. In this case, part (i) follows from Lemma 3.3.4 (ii). If $|Q_{1}\cap Q_{2}|=2$, then $|Q_{1}\cup Q_{2}|=6$, and there are precisely two elements $\alpha,\beta\in R$ that are orthogonal to every root in $Q_{1}\cup Q_{2}$. Let $\Psi$ be the set of roots in $\Phi$ that are orthogonal to $\alpha$ and $\beta$. Then $\Psi$ forms a root system of type $D_{6}$ by Section 2.3, proving the first assertion of (ii), and the sets $Q_{1},Q_{2},Q_{1}\cap Q_{2}$ and $Q_{1}\cup Q_{2}$ all lie in $\Psi$. The other assertions of (ii) now follow by applying the type $D_{n}$ case of the result to $\Psi$.
Finally, suppose that $W$ has type $E_{7}$. We identify the root system with the set of roots orthogonal to the the highest root $\theta$ in the root system of type $E_{8}$ as usual. Then $R\cup\{\theta\}$ is a maximal orthogonal set of roots in type $E_{8}$. The assertions in this case follow by applying the argument of the previous paragraph with $\alpha=\theta$. ∎
4. Quasiparabolic structure
Let $X=X(W)$ be the set of all maximal orthogonal sets of positive roots of $W$, and recall from the introduction that $W$ acts on $X$ naturally via the action $w(\{\beta_{1},\cdots,\beta_{n}\})=\{\lvert w(\beta_{1})\rvert,\cdots,\lvert w(\beta_{n})\rvert\}$. In Section 4, we recall the notion of a quasiparabolic set as defined by Rains and Vazirani [37], and we use the concepts of crossings and nestings to endow the $W$-set $X$ with a quasiparabolic structure.
4.1. Quasiparabolic sets
Quasiparabolic sets were introduced by Rains and Vazirani for a general Coxeter system as follows.
Definition 4.1.1.
[37, Section 2, Section 5] Let $W$ be a Weyl group with generating set $S$ and set of reflections $T$. A scaled $W$-set is a pair $(X,\lambda)$, where $X$ is a $W$-set and $\lambda:X\rightarrow{\mathbb{Z}}$ is a function satisfying $|\lambda(sx)-\lambda(x)|\leq 1$ for all $s\in S$. An element $x\in X$ is $W$-minimal if $\lambda(sx)\geq\lambda(x)$ and is $W$-maximal if $\lambda(sx)\leq\lambda(x)$ for all $s\in S$.
A quasiparabolic set for $W$ is a scaled $W$-set $X$ satisfying the following two properties:
  • (QP1)
    for any $r\in T$ and $x\in X$, if $\lambda(rx)=\lambda(x)$, then $rx=x$;
  • (QP2)
    for any $r\in T$, $x\in X$, and $s\in S$, if $\lambda(rx)>\lambda(x)$ and $\lambda(srx)<\lambda(sx)$, then $rx=sx$.
For a quasiparabolic set $X$, we define $\leq_{Q}$ to be the weakest partial order such that $x\leq_{Q}rx$ whenever $x\in X$, $r\in T$, and $\lambda(x)\leq\lambda(rx)$.
Rains and Vazirani call $\lambda(x)$ the height of $x$, and $\leq_{Q}$ the Bruhat order, but we will refer to them as the level of $x$ and the quasiparabolic order because of the potential for confusion in the context of this paper. It follows from [37, Proposition 5.16] that $\lambda$ is a rank function with respect to the partial order $\leq_{Q}$, so that every covering relation $x<_{Q}y$ satisfies $\lambda(y)=\lambda(x)+1$.
We will show that the set $X$ forms a quasiparabolic set for $W$ in type $E_{7},E_{8}$ or $D_{n}$ with $n$ even under a suitable level function defined in terms of coplanar quadruples. We define the level function and some other useful statistics below.
Definition 4.1.2.
Let $W$ be a Weyl group of type $E_{7},E_{8}$ or $D_{n}$ for $n$ even. Let $R$ be a set of mutually orthogonal roots of $W$, and let $\beta$ be a positive $n$-root of $W$.
  • (i)
    We define the crossing number $C(R)$, the nesting number $N(R)$, and the alignment number $A(R)$ of $R$ to be the numbers of crossings, nestings, and alignments contained in $R^{+}$, respectively.
  • (ii)
    We define the type of $R$ to be the monomial $A^{A(R)}C^{C(R)}N^{N(R)}$, and define the level $\lambda(R)$ of $R$ to be $C(R)+2N(R)$.
  • (iii)
    If $R$ is a maximal orthogonal set of roots, then we say that $R$ is noncrossing, nonnesting, and alignment-free if $C(R)=0$, $N(R)=0$ and $A(R)=0$, respectively; we also call $R$ maximally crossing, maximally nesting, or maximally aligned if we have $N(R)=A(R)=0$, $C(R)=A(R)=0$, or $C(R)=N(R)=0$, respectively.
We also apply all the above definitions to $x$ by applying them to the set of components of $\beta$ in its normal form. (Note that the definitions of type in (ii) and in Definition 3.2.1 are consistent.)
Remark 4.1.3.
By Corollary 3.3.5, when $R$ is a maximal orthogonal set of positive roots, the sum of the numbers $C(R),N(R)$ and $A(R)$ is a constant depending only on $W$ and not on $R$, therefore each of these numbers achieves the maximal possible value when the other two equal zero. This justifies the terms “maximally crossing”, “maximally nesting”, and “maximally aligned” in Definition 4.1.2.(iii).
Example 4.1.4.
Suppose that $W$ has type $D_{6}$, and that $R=\{\varepsilon_{1}\pm\varepsilon_{2},\varepsilon_{3}\pm\varepsilon_{6},\varepsilon_{4}\pm\varepsilon_{5}\}$ is the maximal orthogonal set of positive roots corresponding to the matching $\{12,36,45\}$ (via the natural bijection of Lemma 3.1.1 (ii)). In this case, $R$ contains two alignments, corresponding to the pairs $\{12,45\}$ and $\{12,36\}$, and one nesting, corresponding to the pair $\{36,45\}$. The type of $R$ is therefore $A^{2}C^{0}N^{1}$ (or $AAN$).
We now state the main theorem of this section. Its proof will occupy the next subsection.
Theorem 4.1.5.
Let $W$ be a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even, and let $X$ be the set of maximal orthogonal sets of positive roots of $W$, regarded as a $W$-set under the action
$w(\{\beta_{1},\cdots,\beta_{n}\})=\{\lvert w(\beta_{1})\rvert,\cdots,\lvert w(\beta_{n})\rvert\}.$
Then the set $(X,\lambda)$ is a quasiparabolic set for $W$, where $\lambda:X\rightarrow{\mathbb{Z}}$ is the level function $\lambda(x)=C(x)+2N(x)$.
4.2. Proof of Theorem 4.1.5
We will prove Theorem 4.1.5 by showing that the set $X$ is a scaled $W$-set satisfying the axioms (QP1) and (QP2) of Definition 4.1.1. To this end, we first study how the action of a reflection $s_{\alpha}$ corresponding to a root $\alpha$ can affect the level of a maximal orthogonal set $R$ of roots. Recall from Remark 3.2.4 that $s_{\alpha}$ must replace a coplanar quadruple in $R$ with a feature of a different type whenever $R$ does not contain ±$\pm\alpha$. We will therefore examine how such feature replacements affect the level function $\lambda$.
Example 4.2.1.
Let $W=W(D_{6})$ and $R=\{\varepsilon_{1}\pm\varepsilon_{2},\varepsilon_{3}\pm\varepsilon_{6},\varepsilon_{4}\pm\varepsilon_{5}\}$ be as in Example 4.1.4. Let $\alpha=\varepsilon_{2}-\varepsilon_{4}$, so that $s_{\alpha}$ acts as the transposition $(2,4)$. In this case, the set $s_{\alpha}(R)$ corresponds to the matching $\{14,25,36\}$ and has type $CCC$. The reflection $s_{\alpha}$ changes the alignment $Q=\{\varepsilon_{1}\pm\varepsilon_{2},\ \varepsilon_{4}\pm\varepsilon_{5}\}$ to the crossing $Q^{\prime}=s_{\alpha}(Q)=\{\varepsilon_{1}\pm\varepsilon_{4},\ \varepsilon_{2}\pm\varepsilon_{5}\}$, so $\alpha$ moves an $A$ to a $C$. Note that while $s_{\alpha}$ changes the quadruple $Q$ from an $A$ to a $C$ locally, globally $s_{\alpha}$ does not change the type of $R$ from $AAN$ to $ACN$ but to $CCC$. This is because after the application of $s_{\alpha}$, each collinear pair of roots in $Q$ becomes a new collinear pair that forms a new type of coplanar quadruple with the collinear pair of roots $\varepsilon_{3}\pm\varepsilon_{6}$ outside $Q$.
Part (ii) of the next proposition, however, will imply that if $\alpha$ is minimal among roots moving an $A$ to a $C$, then the global change in the type of $R$ will mirror this local change, so that if $R$ has type $A^{p}C^{q}N^{r}$ then $s_{\alpha}(R)$ has type $A^{p-1}C^{q+1}N^{r}$. In our example, the root $\alpha^{\prime}=\varepsilon_{2}-\varepsilon_{3}$ satisfies the minimality condition since it is simple. The reflection $s_{\alpha^{\prime}}$ changes the coplanar quadruple $\{\varepsilon_{1}\pm\varepsilon_{2},\varepsilon_{3}\pm\varepsilon_{6}\}$ of type $A$ to the coplanar quadruple $\{\varepsilon_{1}\pm\varepsilon_{3},\varepsilon_{2}\pm\varepsilon_{6}\}$ of type $C$, and changes the set $R$ of type $AAN$ to a maximal orthogonal set of type $ACN$.
Proposition 4.2.2.
Let $W$ be a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ with $n$ even, and let $R$ be a maximal set of orthogonal positive roots of type $A^{p}C^{q}N^{r}$. Let $\lambda=C+2N$ be the level function from Definition 4.1.2.
  • (i)
    If $\alpha_{i}$ is a simple root, then either $\alpha_{i}\in R$, or $\lambda(s_{\alpha_{i}}(R))\neq\lambda(R)$. If $\lambda(s_{\alpha_{i}}(R))>\lambda(R)$, then we have $\lambda(s_{\alpha_{i}}(R))=\lambda(R)+1$, and either (1) $s_{\alpha_{i}}$ moves an $A$ to a $C$ and $s_{\alpha_{i}}(R)$ has type $A^{p-1}C^{q+1}N^{r}$, or (2) $s_{\alpha_{i}}$ moves a $C$ to an $N$ and $s_{\alpha_{i}}(R)$ has type $A^{p}C^{q-1}N^{r+1}$.
  • (ii)
    If $Q$ is an alignment in $R$, $Q^{\prime}$ is the corresponding crossing quadruple, and $R^{\prime}=(R\setminus Q)\cup Q^{\prime}$, then $d=\lambda(R^{\prime})-\lambda(R)$ is a positive odd number. If there is a positive root $\alpha$ such that $s_{\alpha}(Q)=Q^{\prime}$ but no positive root $\alpha^{\prime}<\alpha$ moves an $A$ to a $C$ in $R$, then $R^{\prime}$ has type $A^{p-1}C^{q+1}N^{r}$.
  • (iii)
    If $Q$ is a crossing in $R$, $Q^{\prime}$ is the corresponding nesting quadruple, and $R^{\prime}=(R\setminus Q)\cup Q^{\prime}$, then $A(R)=A((R\backslash Q)\cup Q^{\prime})$ and $d=\lambda(R^{\prime})-\lambda(R)$ is a positive odd number. If there is a positive root $\alpha$ such that $s_{\alpha}(Q)=Q^{\prime}$ but no positive root $\alpha^{\prime}<\alpha$ moves a $C$ to an $N$ in $R$, then $R^{\prime}$ has type $A^{p}C^{q-1}N^{r+1}$.
  • (iv)
    If $Q$ is an alignment in $R$, $Q^{\prime}$ is the corresponding nesting quadruple, and $R^{\prime}=(R\setminus Q)\cup Q^{\prime}$, then $d=\lambda((R\backslash Q)\cup Q^{\prime})-\lambda(R)$ is a strictly positive even number.
Lemma 4.2.3.
Proposition 4.2.2 holds if $W$ has type $D_{6}$.
Proof of Lemma 4.2.3.
Throughout the proof, we will identify both $R$ and the coplanar quadruples within $R$ with their corresponding matchings (as in Lemma 3.1.1 and Remark 3.1.6). Recall that we will often write a 2-block $\{a,b\}$ in a matching as $ab$.
Suppose that $\alpha_{i}$ is a simple root, so that the reflection $s_{\alpha_{i}}$ acts as the transposition $(i,i+1)$. We will assume that $\alpha_{i}\not\in R$, so that we have $R=\{ai,(i+1)b,ef\}$ and the coplanar quadruple moved by $s_{\alpha_{i}}$ is $Q=\{ai,(i+1)b\}$. Let $Q^{\prime}$ be any other coplanar quadruple in $R$. Then $Q^{\prime}$ is of the form $\{xy,ef\}$ with $x\in\{a,b\}$ and $y\in\{i,i+1\}$, and we have $s_{\alpha_{i}}(Q^{\prime})=\{\{x,s_{\alpha_{i}}(y)\},ef\}$. Since the numbers $i$ and $i+1$ are only distance 1 apart, the four elements $x,s_{\alpha_{i}}(y),e,f$ appearing in $Q^{\prime}$ have the same relative order as the numbers $x,y,e,f$, therefore $Q^{\prime}$ and $s_{\alpha_{i}}(Q^{\prime})$ have the same type. It follows that $Q$ is the only quadruple in $R$ that is changed to a quadruple of another type by $s_{\alpha_{i}}$. Note that $Q$ will be an alignment if $a<i<i+1<b$; $Q$ will be a crossing if $a<b<i$ or $i+1<a<b$ or $b<i<i+1<a$; and $Q$ will be a nesting if $b<a<i$ or $i+1<b<a$. We have $\lambda(s_{\alpha_{i}}(R))=\lambda(R)+1$ in the first three of these six cases and $\lambda(s_{\alpha_{i}}(R))=\lambda(R)-1$ in the last three cases. The first of the six cases corresponds to the situation in (1), and the second and third cases correspond to the situation in (2). Part (i) follows.
Suppose that $Q$ and $Q^{\prime}$ are as in the statement of (ii), with $Q$ being the alignment $\{a_{1}a_{2},b_{1}b_{2}\}$ for some $a_{1}<a_{2}<b_{1}<b_{2}$. If $b_{1}=a_{2}+1$, then the simple root $s_{a_{2}}$ moves an $A$ to a $C$ by moving $Q$, and we have $R^{\prime}=s_{\alpha_{2}}(R)$, therefore $R^{\prime}$ has type $A^{p-1}C^{q+1}N^{r}$ by (i). If $b_{1}>a_{2}+1$, then $Q$ must be one of the following five quadruples
[…]
Direct computation shows that $d$ equals $5,1,3,1$, and $3$ in these cases, respectively. It follows that $d$ is a positive odd number.
To prove the second assertion in (ii), we prove its contrapositive. If $R^{\prime}$ does not have type $A^{p-1}C^{q+1}N^{r}$, then $Q$ must be one of five quadruples listed in the last paragraph. We claim that in each case, for every positive root $\alpha$ such that $s_{\alpha}(Q)=Q^{\prime}$, there exists a positive root $\alpha^{\prime}<\alpha$ that moves another alignment in $R$ other than $Q$ to a crossing. Specifially, we may always take $\alpha^{\prime}$ to be $\varepsilon_{2}-\varepsilon_{3}$ in the first three cases and $\varepsilon_{4}-\varepsilon_{5}$ in the last two cases. For example, the only possiblities for $\alpha$ if $Q=\{12,56\}$ are $\varepsilon_{2}\pm\varepsilon_{5}$ and $\varepsilon_{1}\pm\varepsilon_{6}$, and for all these possibilities the root $\alpha^{\prime}=\varepsilon_{2}-\varepsilon_{3}$ is smaller than $\alpha$ and moves an alignment in $R$ other than $Q$ to a crossing. This completes the proof of the desired contrapositive.
A similar argument proves (iii). This time, we have $Q=\{a_{1}a_{2},b_{1}b_{2}\}$ for some $a_{1}<b_{1}<a_{2}<b_{2}$. If we have $b_{1}=a_{1}+1$ or $b_{2}=a_{2}+1$, then the simple root $\alpha_{a_{1}}$ or $\alpha_{a_{2}}$ moves $Q$ to an nesting $Q^{\prime}$, so $R^{\prime}$ has type $A^{p}C^{q-1}N^{r+1}$ by (i). The only remaining possibility for $Q$ is $\{14,36\}$. If $\alpha$ is a positive root such that $s_{\alpha}(Q)=Q^{\prime}$, then $\alpha\in\{\varepsilon_{1}\pm\varepsilon_{3},\varepsilon_{4}\pm\varepsilon_{6}\}$. Direct computation shows that any simple root $\alpha^{\prime}<\alpha$ moves some crossing in $R$ other than $Q$ to a nesting.
Finally, (iv) follows by combining (ii) and (iii), since $R^{\prime}$ can be obtained from by first replacing $Q$ in $R$ with its corresponding crossing $Q^{\prime\prime}$ and then replacing $Q^{\prime\prime}$ in the result with $Q^{\prime}$. ∎
Proof of Proposition 4.2.2.
If $W$ has type $D_{4}$ then there are only three possibilities for $R$, and all the assertions follow by direct verification. By Lemma 4.2.3, we may therefore assume that the rank of $W$ is at least $7$.
For each coplanar quadruple $Q\subseteq R$, we define $H_{Q}$ to be the set of all 6-subsets $H$ of $R$ such that there exists a $D_{6}$-subsystem of $\Phi$ containing both $\Psi_{Q}$ and $H$, where $\Psi_{Q}$ is the $D_{4}$-subsystem associated with $Q$. By Proposition 3.3.6, if $Q_{1}$ is any coplanar quadruple in $R$, then either $Q_{1}=Q$, or $Q_{1}\cap Q=\emptyset$, or $\lvert Q_{1}\cap Q\rvert=2$ and there is a unique element of $H_{\Psi}$ that contains both $Q$ and $Q_{1}$.
We now prove (iii). By tracking the contributions towards the crossing number made by the three types of coplanar quadruples $Q_{1}\subseteq R$ just mentioned, we note that
(4.2.1) $C(R^{\prime})-C(R)=-1+\sum_{H\in H_{Q}}\big{(}C((H\backslash Q)\cup Q^{\prime})-C(H)+1\big{)}.$
Here, the term $-1$ comes from the case $Q_{1}=Q$, since the crossing $Q$ in $R$ is replaced by the non-crossing feature $Q^{\prime}$ as we change $R$ to $R^{\prime}$. In the second case where $Q_{1}\cap Q=\emptyset$, the quadruple $Q_{1}$ lies in both $R$ and $R^{\prime}$ and thus does not contribute to the difference $C(R^{\prime})-C(R)$. Finally, every $Q_{1}$ with $\lvert Q_{1}\cap Q\rvert=2$ appears together with $Q$ in a unique element $H$ of $H_{Q}$ and contributes a term in the sum over $H_{Q}$, where we have added 1 to the difference $C((H\backslash Q)\cup Q^{\prime})-C(H)$ to account for the fact that the change from $Q$ to $Q^{\prime}$ in $H$ has been recorded by the term $-1$ in the first case.
Similar arguments based on the facts that $N(Q^{\prime})=N(Q)+1$ and $A(Q^{\prime})=A(Q)$ show that
(4.2.2) $N(R^{\prime})-N(R)=1+\sum_{H\in H_{Q}}\big{(}N((H\backslash Q)\cup Q^{\prime})-N(H)-1\big{)}$
and
(4.2.3) $A(R^{\prime})-A(R)=\sum_{H\in H_{Q}}\big{(}A((H\backslash Q)\cup Q^{\prime})-A(H)\big{)}.$
Since $\lambda=C+2N$, it follows from Equations (4.2.1) and (4.2.2) that
(4.2.4) $\lambda(R^{\prime})-\lambda(R)=1+\sum_{H\in H_{Q}}\big{(}\lambda((H\backslash Q)\cup Q^{\prime})-\lambda(H)-1\big{)}.$
By Lemma 4.2.3.(iii), each summand in the sum over $H_{Q}$ is zero in Equation (4.2.3) and is a nonnegative even number in Equation (4.2.4), therefore $A(R^{\prime})=A(R)$ and the number $d=\lambda(R^{\prime})-\lambda(R)$ is a positive odd number.
To prove the last assertion in (ii), suppose that $s_{\alpha}(Q)=Q^{\prime}$ for some positive root $\alpha$, but no positive root $\alpha^{\prime}<\alpha$ moves a $C$ to an $N$ in $R$. The same minimality condition then applies if $R$ is replaced by an element of $H_{Q}$, so every summand in the sums over $H_{Q}$ in Equations (4.2.1) and (4.2.2) is zero by Lemma 4.2.3 (iii). It follows that $R^{\prime}$ has type $A^{p}C^{q-1}N^{r+1}$, which proves Proposition 4.2.2 (iii).
The proof of Proposition 4.2.2 (ii) follows by a similar but shorter argument, and the proof of Proposition 4.2.2 (iv) follows by combining parts (ii) and (iii).
Finally, to prove Proposition 4.2.2 (i), assume that $\alpha_{i}$ is a simple root such that $\alpha_{i}\not\in R$. We have already proved part (i) if $W$ has type $D_{4}$, and this implies that either $\alpha_{i}$ moves an $A$ to a $C$ or a $C$ to an $N$. In the former case, the conclusions follow from part (ii), and in the latter case, they follow from part (iii), in each case because the simple root $\alpha_{i}$ is minimal in the order $\leq$. ∎
Proof of Theorem 4.1.5.
We first prove (i). Proposition 4.2.2 (i) proves that $(X,\lambda)$ is a scaled $W$-set, so it suffices to show $(X,\lambda)$ satisfies the axioms (QP1) and (QP2). We do so by induction on $n$.
If $n=4$ or $n=6$, then $W$ has type $D$, and the axioms (QP1) and (QP2) can be proved by direct verification or as follows. Suppose $n=2k$. By Remark 3.1.7, it suffices to show that $(X,\lambda)$ is a quasiparabolic set for the symmetric group $S_{n}$. We may identify the set $X$ with the set $X^{\prime}$ of fixed-point free involutions in $S_{n}=W(A_{n-1})$, with each collinear pair $\{\varepsilon_{i}\pm\varepsilon_{j}\}$ in a maximal set $R\in X$ correponding to a factor $(i,j)$ in an involution $\iota\in X^{\prime}$. Under this identification, the actions of $S_{n}$ on $X^{\prime}$ and $X$ coincide with each other, so it suffices to show that $X^{\prime}$ is a quasiparabolic set for $S_{n}$ under the level function $\lambda$. Rains and Vazirani [37, Section 4] proved that $X^{\prime}$ is a quasiparabolic set for $S_{n}$ under the level function $h$ given by $h(\iota)=(\ell(\iota)-k)/2$, where $\ell$ denotes Coxeter length, so it further suffices to show that whenever an involution $\iota\in S_{n}$ corresponds to a maximal set $R$ of positive orthogonal roots, we have $(\ell(\iota)-k)/2=\lambda(R)$. This can be proved by an exhaustive check or by induction on $\lambda(R)$ by using the first three cases in the second paragraph of the proof of Lemma 4.2.3. For example, the involution
$\iota=(13)(26)(45)=s_{2}s_{1}s_{3}s_{5}s_{4}s_{3}s_{5}s_{4}s_{2}\in S_{6}$
corresponds to the set $R=\{13,26,45\}$ of type $ACN$. In this case, we have $(\ell(\iota)-k)/2=(9-3)/2$ and $\lambda(R)=3$, as required.
Now assume $n\geq 7$. Let $r=s_{\alpha}\in T$ be the reflection corresponding to a root $\alpha$, and let $x\in X$. If $rx\neq x$, then ±$\pm\alpha$ are not in $x$, so $r$ moves an $A$ to a $C$, or a $C$ to an $N$, or an $A$ to an $N$, or vice versa by Remark 3.2.4. It follows from Proposition 4.2.2 (ii), (iii), and (iv) that $\lambda(rx)>\lambda(x)$ in all the first three cases and $\lambda(rx)<\lambda(x)$ in the last three cases, threfore axiom (QP1) holds.
To prove axiom (QP2), assume that we have $r\in T$, $x\in X$, $s\in S$, $\lambda(rx)>\lambda(x)$ and $\lambda(srx)<\lambda(sx)$. Then the definition of scaled $W$-sets forces $\lambda(rx)=\lambda(sx)=\lambda(x)+1$, so each of $r$ and $s$ must be an $AC$ or $CN$ move by Proposition 4.2.2 (ii), (iii), and (iv). Let $Q_{1}$ and $Q_{2}$ be the coplanar quadruples of roots in $R$ moved by $r$ and $s$, respectively. Then $Q_{1}$ and $Q_{2}$ are disjoint, or coincide with each other, or intersect in two elements by Proposition 3.3.6. If $Q_{1}$ and $Q_{2}$ were disjoint, we would have $Q_{2}\subseteq r(R)$, so that $s$ would move $Q_{2}$ in $r(R)$, and Proposition 4.2.2 (ii) and (iii) would imply that $\lambda(srx)=\lambda(rx)+1>\lambda(rx)=\lambda(sx),$ contradicting the assumption that $\lambda(srx)<\lambda(sx)$. If $Q_{1}=Q_{2}$, then the fact that $\lambda(rx)=\lambda(sx)=\lambda(x)+1$ implies that $s$ and $r$ must both be $AC$ moves or both be $CN$ moves, according as $Q_{1}=Q_{2}$ is an alignment or crossing, by Proposition 4.2.2. It follows from Remark 3.2.4 that $rx=sx$. Finally, if $\lvert Q_{1}\cap Q_{2}\rvert=2$, then Proposition 3.3.6 implies that there is a subsystem $\Sigma$ of type $D_{6}$ containing both $Q_{1}$ and $Q_{2}$. Applying the inductive hypothesis to $\Sigma$ proves that $rx=sx$, which completes the proof. ∎
We end this subsection by recording some useful consequences of Proposition 4.2.2 concerning sums of $n$-roots:
Corollary 4.2.4.
Let $W$ be a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ with $n$ even.
  • (i)
    If $\beta\leq_{Q}\gamma$ are two positive $n$-roots that are comparable in the quasiparabolic order, then we have $\sigma(\beta)\leq\sigma(\gamma)$, with equality if and only if we have $A(\beta)=A(\gamma)$.
  • (ii)
    If $\alpha_{i}$ is a simple root and $R$ is a maximal orthogonal set of positive roots, then we have
    $\sigma(s_{i}(R))=\begin{cases}\sigma(R)-2\alpha_{i}&\text{\ if\ }\alpha_{i}\in R\text{\ or\ }\alpha_{i}\text{\ moves\ a\ }C\text{\ to\ an\ }A,\\ \sigma(R)+2\alpha_{i}&\text{\ if\ }\alpha_{i}\text{\ moves\ an\ }A\text{\ to\ a\ }C,\\ \sigma(R)&\text{\ otherwise}.\end{cases}$
  • (iii)
    If $\alpha_{i}$ is a simple root and $\beta$ is a positive $n$-root of type $A^{p}C^{q}N^{r}$ such that $\sigma(s_{\alpha_{i}}(\beta))>\sigma(\beta)$, then $s_{\alpha_{i}}(\beta)$ is a positive $n$-root of type $A^{p-1}C^{q+1}N^{r}$, and we have $\sigma(s_{\alpha_{i}}(\beta))-\sigma(\beta)=2\alpha_{i}$. If $\beta$ is nonnesting, then so is $s_{\alpha_{i}}(\beta)$.
Proof.
All the claims can be proved by examining the effects of various types of (simple) reflections on sums and levels of $n$-roots recorded in Theorem 3.2.2 and Proposition 4.2.2. We first prove (i). By the definition of $\leq_{Q}$, it is enough to consider the case where $r\in T$ is a reflection and $\gamma=r(\beta)$ satisfies $\lambda(\beta)<\lambda(\gamma)$. By Proposition 4.2.2, this implies that $r$ is an $AC$, $CN$, or $AN$ move. The proof of (i) then follows from Theorem 3.2.2 (vi).
We now prove (ii). If $\alpha_{i}\in R$, then $\sigma(s_{i}(R))=\sigma(R)-2\alpha_{i}$ because $s_{i}(\alpha_{i})=-\alpha_{i}$. If $\alpha_{i}$ is a $CN$ or $NC$ move, Theorem 3.2.2 (vi) implies that $\sigma(s_{\alpha_{i}}(R))=\sigma(R)$. It follows from Proposition 4.2.2 (i) that the only other possibility is that $\alpha_{i}$ moves an $A$ to a $C$ or vice versa. By Theorem 3.2.2 (vi), this can only happen if $\alpha_{i}$ is the unique simple root in the corresponding nesting, i.e., the root $\alpha_{2}$ in the explicitly constructed sets $\Psi_{2}^{+}=\Psi^{+}_{N}$ in the proof of that theorem. Explicit computation shows that root sums in this case differ by $2\alpha_{i}$ in the precise manner described in the statement, which completes the proof of (ii).
The first part of (iii) follows immediately from (ii) and Proposition 4.2.2 (i). The assertion about nonnesting $n$-roots follows from the special case $r=0$. ∎
Remark 4.2.5.
Burns and Pfeiffer [9, Theorem 1.2] prove that if $T$ is a maximal order abelian subgroup of one of the groups $W$ in Theorem 4.1.5, then $T$ is elementary abelian of order $2^{n}$, where $n$ is the rank of $W$. They also prove that the set of all such subgroups forms a single conjugacy class [9, Theorem 3.1]. It follows that the stabilizers of the elements $x\in X$ can be defined abstractly from the group structure of $W$: they are the normalizers of the maximal order abelian subgroups of $W$.
4.3. Extremal elements
In this subsection, we identify $X$ with the set $\Phi^{+}_{n}$ of positive $n$-roots as usual and discuss an application of Theorem 4.1.5 concerning the maximally aligned and maximally nested $n$-roots, which turn out to be the unique $W$-minimal and $W$-maximal elements of the set $X$. The uniqueness of the maximally aligned and maximally nested $n$-root is not a priori clear, but it will follow conveniently from the general theory of quasiparabolic sets.
Proposition 4.3.1.
Let $W$ be a Weyl group of rank $n$ of types $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even, and let $M$ be the number of coplanar quadruples in a positive $n$-root.
  • (i)
    There is a unique positive $n$-root, $\theta_{A}$, of type $A^{M}$, and it corresponds to the unique $W$-minimal element of the quasiparabolic set $X$ of Theorem 4.1.5.
  • (ii)
    There is a unique positive $n$-root, $\theta_{N}$, of type $N^{M}$, and it corresponds to the unique $W$-maximal element of the quasiparabolic set $X$ of Theorem 4.1.5.
Proof.
We recall that by Theorem 2.8, Remark 2.9, and Corollary 2.10 of [37], every orbit of a quasiparabolic set contains at most one $W$-minimal and at most one $W$-maximal element. If such a $W$-minimal or $W$-maximal element exists, then it can be identified as the unique element with the minimal or maximal possible level, respectively.
The set $X$ is finite, so it has at least one $W$-maximal and one $W$-minimal element. Since $X$ consists of a single $W$-orbit by Lemma 3.1.2, it follows from the paragraph preceding this proposition that $X$ has a unique $W$-maximal and a unique $W$-minimal element, and that they are the unique elements with the minimal and maximal possible level.
Let $R$ be a maximal orthogonal set of positive roots. If $R$ contains any coplanar quadruple $Q$ that is a crossing or a nesting, then by Remark 3.2.4 we can find a reflection $s_{\alpha}$ that moves $Q$ to an alignment or a crossing, respectively, and in both cases Proposition 4.2.2 implies that $\lambda(s_{\alpha}(R))<\lambda(R)$. Iterating this procedure proves the existence of an $n$-root of type $A^{M}$, which achieves the lowest possible value of $\lambda$. The uniqueness property of the previous paragraph then completes the proof of (i). Part (ii) can be proved similarly, by using the fact that any alignment or crossing in $R$ would induce a level-increasing $AC$ or $CN$ move. ∎
We will prove shortly, in Proposition 5.1.2, that $W$ also has a unique maximally crossing element, $\theta_{C}$. The element $\theta_{C}$ will be the unique minimal element in a quasiparabolic set of a parabolic subgroup of $W$.
Remark 4.3.2.
In Theorem 5.3.2 below, we will introduce another partial order, $\leq_{\mathcal{B}}$, on the positive $n$-roots. The argument of [24, Section 6] can be adapted to show that, under suitable identifications, $\leq_{\mathcal{B}}$ refines the monoidal order introduced by Cohen, Gijsbers, and Wales [16, Section 3]. The quasiparabolic order has $\theta_{N}$ and $\theta_{A}$ as its unique maximal and unique minimal elements, whereas the monoidal order (and $\leq_{\mathcal{B}}$) has $\theta_{C}$ as its unique maximal element and has multiple minimal elements.
5. Feature-avoiding elements
In Section 5, we develop the properties of $n$-roots that avoid features of a given type: the alignment-free, noncrossing and nonnesting elements. We show that the alignment-free elements form a quasiparabolic set $X_{I}$ of a maximal standard parabolic subgroup $W_{I}$ of $W$, and that the unique maximally crossing possible $n$-root $\theta_{C}$ is the unique $W_{I}$-minimal element of $X_{I}$. We also show that the sets of noncrossing elements and of nonnesting elements both form bases of the Macdonald representation. Moreover, the basis of noncrossing elements may be viewed as a canonical basis and behaves in a way that is reminiscent of the set of simple roots of a root system (Theorem 5.3.2). The basis of nonnesting elements admits an interesting combinatorial characterization: it is a distributive lattice induced by a suitable Bruhat order (Theorem 5.4.5). Finally, we introduce the notion of $\sigma$-equivalence classes to tie together the alignment-free, noncrossing, and nonnesting elements. These equivalence classes turn out to be intervals with respect to the quasiparabolic order on $X$, and the set $X_{I}$ of alignment-free elements form the top class with respect to a natural partial order. Any set of $\sigma$-equivalence class representatives forms a basis of the Macdonald representation, and the change of basis matrices between any pair of such bases, including the noncrossing and nonnesting bases, are unitriangular (Theorem 5.5.2).
Throughout the rest of Section 5, we assume that we are working with a Weyl group $W$ of rank $n$ and type $E_{7},E_{8}$ or $D_{n}$ for $n$ even. All results hold independently of the rank and type of $W$, and we shall omit the statement of the above assumption except in the main theorems. We define the sum of each positive $n$-root $\gamma$ to be the sum of the components of $\gamma$ in normal form, and we denote it by $\sigma(\gamma)$.
5.1. Alignment-free elements
Recall from Proposition 4.3.1 that $W$ has a unique positive $n$-root $\theta_{N}$ that avoids both alignments and crossings. We will use $\theta_{N}$ to help study general alignment-free elements.
Proposition 5.1.1.
Let $\theta_{N}$ be the unique positive $n$-root of type $N^{M}$.
  • (i)
    A noncrossing $n$-root (i.e., one of type $A^{p}N^{r}$) has a simple component.
  • (ii)
    The $n$-root $\theta_{N}$ has a unique simple component, $\alpha_{x}$.
  • (iii)
    If $\alpha_{i}$ is a simple root, then $B(\sigma(\theta_{N}),\alpha_{i})\geq 0$, where equality holds if and only if $\alpha_{i}\neq\alpha_{x}$.
  • (iv)
    Let $W_{I}$ be the parabolic subgroup of $W$ generated by the set $S\backslash\{\alpha_{x}\}$. Then the stabilizer of $\sigma(\theta_{N})$ is precisely $W_{I}$, and we have $B(\sigma(\theta_{N}),\alpha_{i})\geq 0$ for all simple roots $\alpha_{i}$.
Proof.
Let $\gamma$ be a noncrossing $n$-root of type $A^{p}N^{r}$, and let $R$ be the set of components of $\gamma$. Assume for a contradiction that $R$ contains no simple root, and let $\beta$ be a root of minimal height in $R$. The bilinear form $B$ has the property that $B(\alpha,\alpha^{\prime})\in\{-2,-1,0,1,2\}$ for any roots $\alpha,\alpha^{\prime}$, with $B(\alpha,\alpha^{\prime})=2$ if and only if $\alpha=\alpha^{\prime}$. Furthermore, for any positive root $\alpha$, there exists a simple root $\alpha_{i}$ such that $B(\alpha,\alpha_{i})>0$ [26, Theorem 1.5]. It follows that there exists a simple root $\alpha_{i}$ such that $B(\beta,\alpha_{i})=1$.
Since $\alpha_{i}$ is not in $R$, it moves a coplanar quadruple $Q\subseteq R$, and we have $\beta\in Q$ since $B(\beta,\alpha_{i})\neq 0$. Let $\Psi$ be the $D_{4}$-subsystem associated to $Q$. By hypothesis, $Q$ is either an alignment or a nesting, and $\beta$ is an element of $Q$ of minimal height. It follows from the explicit description of the sets $\Psi^{+}_{A}=\Psi^{+}_{3}$ and $\Psi^{+}_{N}=\Psi^{+}_{2}$ in the proof of Theorem 3.2.2 that $\beta$ is a root in the induced simple system of $\Psi$. On the other hand, since $\alpha_{i}$ is simple root of $W$, $\alpha_{i}$ is also in this induced simple system. This is a contradiction, because we cannot have $B(\gamma_{1},\gamma_{2})>0$ for two simple roots $\gamma_{1}$ and $\gamma_{2}$ in a root system. This completes the proof of (i).
Now suppose further that $\gamma=\theta_{N}$. It follows from (i) that $R$ contains a simple root, so assume for a contradiction that $R$ contains two simple roots, $\alpha_{x}$ and $\alpha_{y}$. Let $P$ be a path from $x$ to $y$ in the Dynkin diagram $\Gamma$, and let $\beta$ be the root $\sum_{p\in P}\alpha_{p}$. Note that $B(\beta,\alpha_{x})=B(\beta,\alpha_{y})=1$, so that $\alpha_{x}$ and $\alpha_{y}$ are both elements of the coplanar quadruple $Q$ consisting of the roots moved by $\beta$. Let $\Psi$ be the $D_{4}$-subsystem associated with $Q$. Then $\alpha_{i},\alpha_{j}$ are both induced simple roots of $\Psi$ since they are simple roots of $W$. On the other hand, the type of $R$ is $N^{M}$, so $Q$ is a nesting and thus contains a unique minimal root by the description of the set $\Psi^{+}_{N}=\Psi^{+}_{2}$ in the proof of Theorem 3.2.2. This is a contradiction, and (ii) follows.
To prove (iii), let ${\alpha_{i}}$ be a simple root. If $\alpha_{i}=\alpha_{x}$, then $\alpha_{x}$ is a component of $\theta_{N}$ and we have
$s_{i}(\sigma(\theta_{N}))=\sigma(\theta_{N})-2\alpha_{i}.$
If $\alpha_{i}\neq\alpha_{x}$, then $\alpha_{i}$ is not a component of $\theta_{N}$, and Proposition 4.2.2 (i) implies that $\alpha_{i}$ moves an $N$ to a $C$. Corollary 4.2.4 (ii) then implies that
$s_{i}(\sigma(\theta_{N})=\sigma(\theta_{N}).$
Equation 2.1.1 implies that $B(\alpha_{i},\sigma(\theta_{N}))\geq 0$ for all $i$, with equality holding if and only if $\alpha_{i}\neq\alpha_{x}$.
Part (iv) follows from (iii) by [26, Theorem 1.12 (a)], which says that the stabilizer of $\sigma(\theta_{N})$ in $W$ is generated by the simple reflections it contains. ∎
Proposition 5.1.2.
Let $\theta_{N}$ be the unique positive $n$-root of type $N^{M}$, let $\alpha_{x}$ be the unique simple component of $\theta_{N}$, and let $W_{I}$ be the parabolic subgroup of $W$ generated by the set $S\backslash\{\alpha_{x}\}$.
  • (i)
    The $W_{I}$-orbit of positive $n$-roots that contains $\theta_{N}$ is a quasiparabolic set $(X_{I},\lambda_{I})$ for $W_{I}$, where $\lambda_{I}$ is the restriction of $\lambda$ to $X_{I}$.
  • (ii)
    The following are equivalent for a positive $n$-root $\beta$:
    • (1)
      $\beta$ has type $C^{q}N^{r}$ for some $q$ and $r$;
    • (2)
      $\sigma(\beta)=\sigma(\theta_{N})$;
    • (3)
      $\beta$ is an element of $X_{I}$.
    In particular, the elements of the quasiparabolic set $X_{I}$ are precisely the alignment-free positive $n$-roots.
  • (iii)
    There is a unique positive $n$-root, $\theta_{C}$, of type $C^{M}$, and it corresponds to the unique $W$-minimal element of the quasiparabolic set $X_{I}$.
  • (iv)
    If $\alpha$ is a root of $W_{I}$ and $\beta$ is an $n$-root in $X_{I}$ whose components do not contain ±$\pm\alpha$, then $\lambda(s_{\alpha}(\beta))=\lambda(\beta)+1\mod 2$.
Proof.
Part (i) follows from Theorem 4.1.5 by restriction.
We now prove the implication (1) $\Rightarrow$ (2) of part (ii). Let $\beta$ be an $n$-root of type $C^{q}N^{r}$. If $q=0$, then $\beta=\theta_{N}$ by Proposition 4.3.1 (ii) and (2) follows immediately, so suppose that $q>0$. By Remark 3.2.4 and Proposition 4.2.2 (iii), there exists a reflection $\alpha$ that moves a crossing to a nesting in such a way that the $n$-root $\beta^{\prime}=s_{\alpha}(\beta)$ has type $C^{q-1}N^{r+1}$. Corollary 4.2.4 (ii) proves that $\sigma(\beta^{\prime})=\sigma(\beta)$. It now follows from induction on $q$ that $\sigma(\beta)=\sigma(\theta_{N})$, proving (2).
To prove (2) $\Rightarrow$ (3), assume that $\sigma(\beta)=\sigma(\theta_{N})$. By Lemma 3.1.2, there exists $w\in W$ such that $w(\theta_{N})=\beta$. It follows that
$w\sigma(\theta_{N})=\sigma(w(\theta_{N}))=\sigma(\beta)=\sigma(\theta_{N}).$
It follows that $w\in W_{I}$ and $\beta\in X_{I}$, which proves (3).
To prove (3) $\Rightarrow$ (1), let $\beta=w(\theta_{N})$ for some $w\in W_{I}$ and let $w=s_{i_{1}}s_{i_{2}}\cdots s_{i_{k}}$ be a reduced word of $w$. Then each simple reflection $s_{i_{j}}$ fixes $\sigma(\theta_{N})$, so it follows from Corollary 4.2.4 (ii) proves that $s_{i_{j}}$ is a $CN$ or $NC$ move. It follows from Proposition 4.2.2 (i) that $\beta$ has type $C^{q}N^{r}$, which implies (1) and completes the proof of (ii).
To prove (iii), note that the quasiparabolic set $X_{I}$ is finite and transitive, so it follows, as in the proof of Proposition 4.3.1 (i), that there $X_{I}$ has a unique minimal element, namely, the unique element having minimal level in $X_{I}$. The elements of $X_{I}$ are all of type $C^{q}N^{r}$ by (ii) where $q+r=M$ and $M$ is as in Corollary 3.3.5, so we have $\lambda_{I}(\gamma)\geq M$ for any element $\gamma\in X_{I}$, with equality holding if and only if $\gamma$ has type $C^{M}$. To prove (iii) it now remains to show that such an element exists.
Let $\beta$ be an $n$-root in $X_{I}$, and suppose $\beta$ has type $C^{q}N^{r}$ where $r>0$. Then $\sigma(\beta)=\sigma(\theta_{N})$, and $\beta$ admits an $NC$ move by a reflection $s_{\alpha}$ correponding to some root $\alpha$. Theorem 3.2.2 (vi) implies that $\sigma(s_{\alpha}(\beta))=\sigma(\beta)=\sigma(\theta_{N})$, so $s_{\alpha}(\beta)$ is in $X_{I}$ and has type $C^{a}N^{b}$ by (ii). Proposition 4.2.2 (iii) implies that $a>q$ and $b<r$, and that $s_{\alpha}(\beta)$ has a lower level than $\beta$. It follows that $X_{I}$ has an element of type $C^{M}$, which completes the proof of (iii).
Suppose that $\alpha$ and $\beta$ are as in the statement of (iv). Since neither of ±$\pm\alpha$ is a component of $\beta$, the reflection $s_{\alpha}$ must move a $C$ to an $N$ or vice versa by Remark 3.2.4 and (ii). Proposition 4.2.2 (iii) then implies that $\lambda(\beta)$ and $\lambda(s_{\alpha}(\beta))$ have opposite parities. ∎
Remark 5.1.3.
With some work, it can be shown that the positive $n$-roots in $X_{I}$ are also exactly the positive $n$-roots $\beta$ with the property that every component $\alpha$ of $\beta$ has $x$-height 1, in the sense that $\alpha$ expands into a linear combination of simple roots where the simple root $\alpha_{x}$ appears with coefficient $1$. Since the simple roots of $W_{I}$ do not include $\alpha_{x}$, it follows that no root of $W_{I}$ divides any $n$-root in $X_{I}$. In other words, the “if” condition in Proposition 5.1.2 (iv) in fact holds for every root $\alpha$ of $W_{I}$ and every $n$-root $\beta\in X_{I}$. This implies that the sets of all $n$-roots in $X_{I}$ with even levels and of all $n$-roots in $X_{I}$ with odd levels are interchanged by $s_{\alpha}$ for every $\alpha\in W_{I}$. In particular, these two sets have the same cardinality.
5.2. Two bases
The goal of this subsection is to prove that the noncrossing $n$-roots and the nonnesting $n$-roots each form a basis for the Macdonald reprsentation. The proof is based on a commutative version of Bergman’s diamond lemma [5], which is a special case of Newman’s diamond lemma [35]. We define the crossing order, $\leq_{C}$, on the set of positive $n$-roots by declaring that $\beta\leq_{C}\gamma$ if either $\sigma(\beta)<\sigma(\gamma)$, or both $\sigma(\beta)=\sigma(\gamma)$ and $\beta\geq_{Q}\gamma$, where $\leq_{Q}$ is the quasiparabolic order. Similarly, we define the nesting order, $\leq_{N}$, on the set of positive $n$-roots by declaring that $\beta\leq_{N}\gamma$ if either $\sigma(\beta)<\sigma(\gamma)$, or both $\sigma(\beta)=\sigma(\gamma)$ and $\beta\leq_{Q}\gamma$.
Consider the relations $\gamma\gamma_{C}=\gamma\gamma_{N}+\gamma\gamma_{A}$ of Theorem 3.2.2 (vi). We have $\gamma\gamma_{A}<_{C}\gamma\gamma_{C}$ and $\gamma\gamma_{A}<_{N}\gamma\gamma_{N}$ because $\sigma(\gamma_{A})<\sigma(\gamma_{C})=\sigma(\gamma_{N})$ by Theorem 3.2.2 (vi). We also have $\gamma_{C}<_{N}\gamma_{N}$ and $\gamma_{N}<_{C}\gamma_{C}$ by Proposition 4.2.2 (iii) and the definition of $\leq_{Q}$, because for any component $\alpha$ of $\theta_{A}$, the reflection $s_{\alpha}$ moves the components of $\gamma_{C}$ to those of $\gamma_{N}$ by Theorem 3.2.2 (vi). It also follows that $\lambda(\gamma\gamma_{C})<\lambda(\gamma\gamma_{N})$ and $\gamma\gamma_{C}\leq_{Q}\gamma\gamma_{N}$. We may therefore regard the relations $\gamma\gamma_{C}=\gamma\gamma_{N}+\gamma\gamma_{A}$ as directed reduction rules, each of which operates on a single term $\lambda_{i}\beta_{i}$ in a linear combination $\sum_{i}\lambda_{i}\beta_{i}$, where the $\beta_{i}$ are positive $n$-roots. Each reduction rule can be used either (a) to express a positive $n$-root $\gamma\gamma_{C}$ containing a crossing as the sum of two other positive $n$-roots $\gamma\gamma_{A},\gamma\gamma_{N}$ that are strictly lower than it in the crossing order, or (b) to express a positive $n$-root $\gamma\gamma_{N}$ containing a nesting as a linear combination of two other positive $n$-roots $\gamma\gamma_{A},\gamma\gamma_{C}$ that are strictly lower than it in the nesting order.
In order to apply the diamond lemma, we need to know (a) that it is never possible to apply an infinite sequence of reduction rules to a linear combination of $n$-roots, and (b) that the reduction rules are confluent. The latter condition means that if $m$ is a linear combination of $n$-roots and if $f_{1}$ and $f_{2}$ are two different reductions that can be applied to $m$, then the linear combinations $f_{1}(m)$ and $f_{2}(m)$ themselves have a common reduction, $m^{\prime}$. In other words, it is possible to reduce $f_{1}(m)$ to $m^{\prime}$ by applying a suitable sequence of reductions, and it is possible to reduce $f_{2}(m)$ to the same $m^{\prime}$ by applying a possibly different sequence of reductions. If these two conditions hold, the conclusion of the diamond lemma is that every element of the module may be uniquely expressed as an element to which no reduction rules may be applied; in other words, a unique linear combination of noncrossing $n$-roots, or a unique linear combination of nonnesting $n$-roots.
Conversely, the diamond lemma guarantees that if each element $m$ can be uniquely expressed as a linear combination of nonnesting (or noncrossing) $n$-roots, then the reduction relations are confluent.
Lemma 5.2.1.
There are $2$ nonnesting positive $4$-roots in type $D_{4}$, and $5$ nonnesting positive $6$-roots in type $D_{6}$. The nonnesting positive $n$-roots are linearly independent in each case.
Proof.
In type $D_{4}$, the set in question is $\{(\varepsilon_{1}^{2}-\varepsilon_{2}^{2})(\varepsilon_{3}^{2}-\varepsilon_{4}^{2}),\ (\varepsilon_{1}^{2}-\varepsilon_{3}^{2})(\varepsilon_{2}^{2}-\varepsilon_{4}^{2})\}$, which is clearly linearly independent. In type $D_{6}$, the nonnesting positive $6$-roots correspond to the matchings
$\{14,25,36\},\ \{13,25,46\},\ \{13,24,56\},\ \{12,35,46\},\,\text{and}\;\{12,34,56\}.$
One can check that this set is linearly independent by comparing coefficients of $\varepsilon_{1}^{2}\varepsilon_{2}^{2}\varepsilon_{3}^{2}$, $\varepsilon_{1}^{2}\varepsilon_{2}^{2}\varepsilon_{4}^{2}$, $\varepsilon_{1}^{2}\varepsilon_{2}^{2}\varepsilon_{5}^{2}$, $\varepsilon_{1}^{2}\varepsilon_{3}^{2}\varepsilon_{4}^{2}$, and $\varepsilon_{1}^{2}\varepsilon_{3}^{2}\varepsilon_{5}^{2}$. ∎
Theorem 5.2.2.
Let $W$ be a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even. Let $j^{\Phi}_{nA_{1}}({\rm{sgn}})$ be the Macdonald representation of $W$.
  • (i)
    The nonnesting positive $n$-roots form a ${\mathbb{Q}}$-basis for $j^{\Phi}_{nA_{1}}({\rm{sgn}})$.
  • (ii)
    The noncrossing positive $n$-roots form a ${\mathbb{Q}}$-basis for $j^{\Phi}_{nA_{1}}({\rm{sgn}})$.
  • (iii)
    The alignment-free positive $n$-roots span $j^{\Phi}_{nA_{1}}({\rm{sgn}})$.
Proof.
We first prove (i) by using the reduction rule $\gamma\gamma_{N}=\gamma\gamma_{C}-\gamma\gamma_{A}$ of Theorem 3.2.2 (vi) to express an $n$-root that contains a nesting as a linear combination of $n$-roots that are strictly lower in the nesting order. There are no infinite descending chains in the crossing order because there are only finitely many $n$-roots. It remains to show that the reductions $f_{i}$ are confluent, by induction on the rank $n$. By Lemma 5.2.1, this is already known to be the case in types $D_{4}$ and $D_{6}$, so we assume from now on that we have $n>6$.
If two reductions, $f_{i}$ and $f_{j}$, affect different terms $\lambda_{i}\beta_{i}$ in the linear combination $m=\sum_{i}\lambda_{i}\beta_{i}$, or if $f_{i}$ and $f_{j}$ affect disjoint components of the same term $\lambda_{i}\beta_{i}$, then $f_{i}$ and $f_{j}$ commute. It is then immediate that $f_{i}(m)$ and $f_{j}(m)$ have a common reduction, namely, $f_{i}f_{j}(m)=f_{j}f_{i}(m)$. The proof of confluence now reduces to proving that if $f_{i}$ and $f_{j}$ change at least one component in the same $n$-root $\beta$, then $f_{i}(\beta)$ and $f_{j}(\beta)$ have a common reduction. In this case, if $Q_{i}$ and $Q_{j}$ are the sets of components of $\beta$ that are moved by $f_{i}$ and $f_{j}$ respectively, then Proposition 3.3.6 (i) implies that either $Q_{i}=Q_{j}$, or $|Q_{i}\cap Q_{j}|=2$. In the first case, we have $f_{i}=f_{j}$, and there is nothing to prove. In the latter case, Proposition 3.3.6 (ii) implies that there is a root subsystem $\Psi$ of type $D_{6}$ that contains $Q_{i}$ and $Q_{j}$ as coplanar quadruples. Confluence now follows by applying the inductive hypothesis to $\Psi$, which completes the proof of (i).
We now prove (ii) by using the reduction rule $\gamma\gamma_{C}=\gamma\gamma_{N}+\gamma\gamma_{A}$ of Theorem 3.2.2 (vi) to express an $n$-root that contains a crossing as a sum of $n$-roots that are strictly lower in the crossing order, $\leq_{C}$. It follows that the noncrossing positive $n$-roots form a spanning set. There are $2$ noncrossing positive $4$-roots in type $D_{4}$, corresponding to the matchings $\{12,34\}$ and $\{14,23\}$, and $5$ noncrossing positive $6$-roots in type $D_{6}$, corresponding to the matchings
$\{16,25,34\},\ \{16,23,45\},\ \{14,23,56\},\ \{12,36,45\},\,\text{and}\;\{12,34,56\}.$
These spanning sets are bases of $j^{\Phi}_{nA_{1}}({\rm{sgn}})$ by (i), and the rest of the argument used to prove (i) now applies mutatis mutandis.
Part (iii) follows by expressing the reduction rule in the form $\gamma\gamma_{A}=\gamma\gamma_{C}-\gamma\gamma_{N}.$ By Theorem 3.2.2 (vi), we have $\sigma(\gamma\gamma_{A})<\sigma(\gamma\gamma_{N})=\sigma(\gamma\gamma_{C})$. This implies that the relation can only be applied finitely many times before the procedure terminates, and (iii) follows. ∎
We will refer to the bases of nonnesting and noncrossing positive $n$-roots as the nonnesting basis and noncrossing basis of the Macdonald representation.
5.3. The noncrossing basis is the canonical basis.
In this subsection, we show that the noncrossing basis behaves in the Macdonald representation in many ways like a simple system in the reflection representation. In particular, every $n$-root decomposes into the noncrossing basis with coefficients of like sign, and the noncrossing $n$-roots are precisely the minimal ones that are minimal in the sense that they are not further decomposable. This minimality property yields an elementary algebraic characterization. We also show that the maximally crossing $n$-root $\theta_{C}$ has a maximal decomposition into the noncrossing basis in a natural sense, and that simple reflections act on the noncrossing basis in a way reminiscent of the way they act on a simple system in the reflection representation. For these reasons, we may think of the noncrossing basis as the canonical basis of the Macdonald representation.
Lemma 5.3.1.
If $\beta$ and $\lambda\beta$ are both $n$-roots for some scalar $\lambda$, then we must have $\lambda=\pm 1$.
Proof.
Theorem 5.2.2 implies that the scalar $\lambda$ in (ii) lies in ${\mathbb{Q}}$. Lemma 3.1.2 implies that there exists $w\in W$ such that $w(\beta)=\lambda\beta$. Because $w$ has finite order, it follows that $\lambda$ is a root of unity, and this forces $\lambda=\pm 1$. ∎
Theorem 5.3.2.
Let $W$ be a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even, and let $\mathcal{B}$ be the set of noncrossing positive $n$-roots.
  • (i)
    Every $n$-root is a ${\mathbb{Z}}$-linear combination of elements of $\mathcal{B}$, with coefficients of like sign. This sign is positive if the $n$-root is positive, and is negative if the $n$-root is negative.
  • (ii)
    A positive $n$-root is noncrossing if and only if it is not a positive linear combination of other positive $n$-roots.
  • (iii)
    Define $\gamma\leq_{\mathcal{B}}\gamma^{\prime}$ for positive $n$-roots $\gamma=\sum_{\beta\in\mathcal{B}}c_{\beta}\beta$ and $\gamma^{\prime}=\sum_{\beta\in\mathcal{B}}c^{\prime}_{\beta}\beta$ whenever $c_{\beta}\leq d_{\beta}$ for all $\beta\in\mathcal{B}$. Then $\leq_{\mathcal{B}}$ is a partial order on the set $X$ of positive $n$-roots. The maximally crossing element $\theta_{C}$ is the unique maximal element of $X$ with respect to $\leq_{\mathcal{B}}$.
  • (iv)
    If $\gamma\in\mathcal{B}$ and $\alpha_{i}$ is a simple root, then we have
    $s_{\alpha_{i}}(\gamma)=\begin{cases}-\gamma&\text{\ if\ }\alpha_{i}|\gamma;\\ \gamma+\gamma^{\prime}&\text{\ otherwise,\ for\ some\ }\gamma^{\prime}\in\mathcal{B}{\text{\ such\ that\ }}\alpha_{i}|\gamma^{\prime}.\\ \end{cases}$
Proof.
Let $\beta$ be a positive $n$-root. By the proof of Theorem 5.2.2, the result of applying reductions of the form $\gamma\gamma_{C}=\gamma\gamma_{N}+\gamma\gamma_{A}$ to $\beta$ until this is no longer possible has the effect of expressing $\beta$ as a positive integer linear combination of noncrossing $n$-roots, and this procedure will always terminate after finitely many steps. This proves (i) for positive $n$-roots, and the statement for negative $n$-roots follows because $n$-roots occur in positive-negative pairs.
If $\beta$ is a positive $n$-root that contains a crossing, then $\beta$ is a positive linear combination of other positive $n$-roots by applying the reduction rule in the first paragraph. Conversely, suppose that $\beta$ is a noncrossing $n$-root and that $\beta=\sum_{i}\lambda_{i}\beta_{i}$, where $\lambda_{i}>0$ and $\beta_{i}$ is a positive $n$-root that is different from $\beta$ for each $i$. Part (i) implies that each of the $\beta_{i}$ is a positive linear combination of noncrossing $n$-roots. Because no cancellation can occur in the sum, Theorem 5.2.2 (ii) implies that this is only possible if each $\beta_{i}$ is a multiple of $\beta$. Collecting terms, we then have $\beta=\lambda\beta_{i}$. Lemma 5.3.1 and the assumption that $\beta_{i}$ is positive then imply that $\lambda=1$ and $\beta=\beta_{i}$, which is a contradiction.
The relation $\leq_{\mathcal{B}}$ in (iii) is clearly a partial order on $X$. Since $X$ is finite, it contains at least one maximal positive $n$-root with respect to $\leq$. To prove (iii), it then suffices to show that for every $n$-root $\gamma\in X$ not equal to $\theta_{C}$, there is an element $\gamma^{\prime}\in X$ such that $\gamma<_{\mathcal{B}}\gamma^{\prime}$. Let $\gamma\in X$ be an $n$-root not equal to $\theta_{C}$, so that we can factorize $\gamma$ as $\gamma_{1}\gamma^{\prime}$, where $\gamma^{\prime}$ is either an alignment or a nesting. In either case, Theorem 3.2.2 (vi) implies that there exists a reflection $s_{\alpha}$ such that $s_{\alpha}(\gamma^{\prime})$ is a crossing, and that we have $s_{\alpha}(\gamma^{\prime})=\gamma^{\prime}+\gamma^{\prime\prime}$, where $\gamma^{\prime\prime}$ is a nesting if $\gamma^{\prime}$ is an alignment, or vice versa. We then have
$s_{\alpha}(\gamma)=\gamma+\gamma_{1}\gamma^{\prime\prime},$
where $\gamma_{1}\gamma^{\prime\prime}$ is also a positive $n$-root. If we write the $n$-root $s_{\alpha}(\gamma)$ as $s_{\alpha}(\gamma)=\sum_{\beta\in\mathcal{B}}e_{\beta}\beta$, then it follows from (i) that $c_{\beta}\leq e_{\beta}$ for all $\beta\in\mathcal{B}$. It follows that $\gamma<_{\mathcal{B}}s_{\alpha}(\gamma)$, proving (iii).
In the situation of (iv), it is immediate that if $\alpha_{i}$ is a component of $\gamma$, then $s_{i}(\gamma)=-\gamma$. Suppose from now on that this is not the case, and let $A^{p}N^{r}$ be the type of $\gamma$. Let $R$ be the set of components of $\gamma$, let $Q\subseteq R$ be the coplanar quadruple moved by $\alpha_{i}$, and let $\Psi$ be the $D_{4}$-subsystem of $Q$. Then the sets $Q,Q^{\prime}=s_{i}(Q)$ and $Q^{\prime\prime}=\Psi^{+}\setminus(Q\cup Q^{\prime})$ are the three distinct coplanar quadruples partitioning the induced positive system by Proposition 3.1.8 (ii) and Theorem 3.2.2. Let $\Psi^{+}_{A},\Psi^{+}_{C}$ and $\Psi^{+}_{N}$ be the alignment, crossing, and nesting in $\Psi^{+}$, respectively. Then since $\alpha_{i}$ is a simple root, Proposition 4.2.2 (i) implies that we must have one of the following two situations:
  1. (1)
    $Q=\Psi^{+}_{A}$, $Q^{\prime}=\Psi^{+}_{C}$, $s_{i}(\gamma)$ has type $A^{p-1}CN^{r}$, and $Q^{\prime\prime}=\Psi^{+}_{N}$;
  2. (2)
    $Q=\Psi^{+}_{N}$, $Q^{\prime}=\Psi^{+}_{C}$, $s_{i}(\gamma)$ has type $A^{p}CN^{r-1}$, and $Q^{\prime\prime}=\Psi^{+}_{A}$.
Let $\gamma_{x}=\prod_{\beta\in x}\beta$ for all $x\in\{Q,Q^{\prime},Q^{\prime\prime}\}$, and write $\gamma=\gamma_{1}\gamma_{Q}$. Then we have $\gamma_{Q^{\prime}}=\gamma_{Q}+\gamma_{Q^{\prime\prime}}$ and thus
$s_{i}(\gamma)=\gamma_{1}\gamma_{Q^{\prime}}=\gamma+\gamma_{1}\gamma_{Q^{\prime\prime}}$
in both of the above cases. We have $\alpha_{i}\in\Psi$ by Proposition 3.1.8 (ii), and $\alpha_{i}$ must lie in the induced simple system of $\Psi$ since it is simple. Theorem 3.2.2 (v) then implies that $\alpha\notin Q^{\prime}$, and we have $\alpha_{i}\notin Q$ by assumption, so we have $\alpha\in Q^{\prime\prime}$. It follows that $\alpha$ divides the $n$-root $\gamma^{\prime\prime}=\gamma_{1}\gamma_{Q^{\prime\prime}}$.
It remains to prove that $\gamma^{\prime\prime}$ is noncrossing. We treat case (1) first. If $\alpha$ is any root that is minimal in $Q$, then $\alpha$ moves the crossing $Q^{\prime}$ to the nesting $Q^{\prime\prime}$ by Theorem 3.2.2 (vi), so that $\gamma^{\prime\prime}=s_{\alpha}(s_{i}(\gamma))$. Since $s_{i}(\gamma)$ has type $A^{p-1}CN^{r}$, $Q^{\prime}$ is the unique crossing in $s_{i}(\gamma)$, therefore any root that moves a $C$ in $s_{i}(\gamma)$ to an $N$ must move $Q^{\prime}$, and it must move $Q^{\prime}$ to $Q^{\prime\prime}$ by Proposition 3.1.8 (ii). Together with Theorem 3.2.2 (vi), this further implies that any root moving $Q^{\prime}$ to $Q^{\prime\prime}$ must come from $Q$, so it follows that $\alpha$ is minimal among roots moving a $C$ in $s_{i}(\gamma)$ to an $N$. It follows from Proposition 4.2.2 (iii) that $\gamma^{\prime\prime}=s_{\alpha}(s_{i}(\gamma))$ has type $A^{p-1}N^{r+1}$, therefore $\gamma^{\prime\prime}$ is noncrossing. A similar argument shows that $\gamma^{\prime\prime}$ has type $A^{p+1}N^{r-1}$ in case (2), so $\gamma^{\prime\prime}$ is noncrossing in both cases. ∎
Remark 5.3.3.
  • (i)
    There are other constructions of the basis of noncrossing $n$-roots; for example, one can modify the monomial bases of [20] by specializing the parameter to $1$ and twisting by the sign representation. However, the $n$-root approach has the significant advantage that it is relatively easy, given an arbitrary group element $w$ and an arbitrary $n$-root $\alpha$, to express $w(\alpha)$ as a linear combination of basis elements. The bases in type $D_{n}$ may be constructed diagrammatically in terms of perfect matchings, as we explain at the end of Section 6.1. There is also a diagrammatic construction in types $E_{7}$ and $E_{8}$, as described in [49] and [22], but we do not pursue this here because it is not easy to recover the components of a basis $n$-root by inspection of the corresponding diagram.
  • (ii)
    With some more work, it can be shown that every component of a noncrossing $n$-root has odd height, and conversely, that every root of odd height occurs as a component of some noncrossing $n$-root.
5.4. The nonnesting basis forms a distributive lattice.
In this subsection, we show that the nonnesting basis is naturally indexed by a distributive lattice whose unique maximal and minimal elements are given by the maximally crossing and aligned $n$-roots $\theta_{C}$ and $\theta_{A}$, respectively. This lattice is induced by the left weak Bruhat order $\leq_{L}$ of $W$ and is isomorphic to a lattice consisting of certain fully commutative elements. We recall that $\leq_{L}$ is defined by the condition that $v\leq_{L}w$ if $w=uv$ for some $u\in W$ such that $\ell(w)=\ell(u)+\ell(v)$, or, equivalently, by the condition that $\ell(wv^{-1})+\ell(v)=\ell(w)$. An element $w$ in a simply laced Weyl group is fully commutative precisely when no reduced word for $w$ contains a factor of the form $s_{i}s_{j}s_{i}$ [47].
Definition 5.4.1.
Let $\theta,\theta^{\prime}$ be two nonnesting positive $n$-roots. A nonnesting sequence from $\theta$ to $\theta^{\prime}$ is a (possibly trivial) sequence $(\theta_{i})=(\theta_{0}=\theta,\theta_{1},\ldots,\theta_{r}=\theta^{\prime})$ of positive nonnesting $n$-roots such that for all $1\leq j\leq r$ there exists a simple root $\alpha_{i_{j}}$ such that
(5.4.1) $s_{i_{j}}(\theta_{j-1})=\theta_{j}\;\text{ and }\;\sigma(\theta_{j})>\sigma(\theta_{j-1}).$
If $s_{i_{1}},s_{i_{2}},\cdots,s_{i_{r}}$ are simple reflections satisfying the condition in (5.4.1), we say that $\mathbf{w}=s_{i_{1}}s_{i_{2}}\cdots s_{i_{r}}$ is a $(\theta,\theta^{\prime})$-word, and we call the element $w$ expressed by $\mathbf{w}$ a $(\theta,\theta^{\prime})$-element. Note that we have $w(\theta^{\prime})=\theta$.
Remark 5.4.2.
Let $\theta$ be a nonnesting positive $n$-root of type $A^{p}C^{q}$ and let $\alpha_{i}$ be a simple root. The condition that $\sigma(s_{i}(\theta))>\sigma(\theta)$ is equivalent to the condition that $B(\sigma(\theta),\alpha_{i})<0$ by Equation (2.1.1), because $\sigma(s_{i}(\theta))=s_{i}(\sigma(\theta))$. In addition, by Corollary 4.2.4 (ii) and (iii), the condition that $\sigma(s_{i}(\theta))>\sigma(\theta)$ is also equivalent to the condition that $\sigma(s_{i}(\theta))=\sigma(\theta)+2\alpha_{i}$, or the condition that $s_{i}$ moves $\theta$ to an $n$-root of type $A^{p-1}C^{q+1}$. It follows that if $(\theta_{0},\theta_{1},\cdots,\theta_{r})$ is a nonnesting sequence, then we have $\lambda(\theta_{j})=\lambda(\theta_{j-1})+1$ for all $1\leq j\leq r$. In particular, every nonnesting sequence is a saturated chain with respect to the quasiparabolic order $\leq_{Q}$.
Remark 5.4.3.
Let $\theta$ and $\theta^{\prime}$ be two positive $n$-roots with $\lambda(\theta^{\prime})>\lambda(\theta)$, and let $w\in W$ be an element such that $w(\theta^{\prime})=\theta$. Let $\mathbf{w}=s_{i_{1}}\cdots s_{i_{r}}$ be an arbitrary word for $w$. By the definition of quasiparabolic sets, applying a simple reflection decreases the level by at most 1, so any element taking $\theta^{\prime}$ to $\theta$ has length at least $\lambda(\theta^{\prime})-\lambda(\theta)$. It follows that $r\geq\lambda(\theta^{\prime})-\lambda(\theta)$. It also follows that if $r=\lambda(\theta^{\prime})-\lambda(\theta)$, then $\mathbf{w}$ is reduced and successively applying the simple reflections $s_{i_{r}},\cdots,s_{i_{2}},s_{i_{1}}$ starting from $\theta^{\prime}$ must reduce the level by 1 at each step. In particular, if $r=\lambda(\theta^{\prime})-\lambda(\theta)$ and $\theta^{\prime}$ is nonnesting, then it follows from Proposition 4.2.2 (i) that each of these simple reflections is a $CA$ move, so that conversely the sequence $\theta_{\mathbf{w}}:=(\theta_{0},\cdots,\theta_{r})$ defined by $\theta_{0}=\theta,\theta_{j}=s_{i_{j}}(\theta_{j-1})$ for $1\leq j\leq r$ must be a nonnesting sequence by Remark 5.4.2.
Proposition 5.4.4.
Let $M$ be the number of quadruples in each $n$-root of $W$, let $W_{I}$ be the parabolic subgroup of Proposition 5.1.2 (i), and let $\theta$ be a nonnesting positive $n$-root of type $A^{p}C^{q}$.
  • (i)
    If $\theta$ is the maximally crossing element $\theta_{C}$, then we have $B(\sigma(\theta),\alpha_{i})\geq 0$ for every simple root $\alpha_{i}$. Otherwise, there exists a simple root $\alpha_{i}$ such that $B(\sigma(\theta),\alpha_{i})<0$.
  • (ii)
    There exists a nonnesting sequence from $\theta$ to $\theta_{C}$, and we have $\text{\rm ht}(\sigma(\theta))=\text{\rm ht}(\sigma(\theta_{C}))-2p$.
  • (iii)
    Every $(\theta,\theta_{C})$-word is reduced. Every $(\theta,\theta_{C})$-element is fully commutative and has length $p$. Every shortest element taking $\theta_{C}$ to $\theta$ has length $p$.
  • (iv)
    There is a unique $(\theta,\theta_{C})$-element $w$. It is the unique shortest element in the coset $wW_{I}$, and is also the unique shortest element in $W$ taking $\theta_{C}$ to $\theta$.
  • (v)
    There exists a nonnesting sequence from the maximally aligned element $\theta_{A}$ to $\theta_{C}$ that includes $\theta$.
Proof.
The first assertion of (i) follows from Proposition 5.1.1 (iii) and the fact that $\sigma(\theta_{C})=\sigma(\theta_{N})$ by Proposition 5.1.2. Let $V$ be the reflection representation of $W$ and let
$D=\{v\in V:B(v,\alpha_{i})\geq 0\text{\ for\ all\ simple\ roots\ }\alpha_{i}\}.$
The set $D$ is a fundamental domain for the action of $W$ on $V$ by [26, Theorem 1.12 (a)], and we have $\theta_{C}\in D$. If $\theta$ is a nonnesting $n$-root different from $\theta_{C}$, then $\theta$ and $\theta_{C}$ are conjugate under the action of $W$ by Lemma 3.1.2, therefore so are $\sigma(\theta)$ and $\sigma(\theta_{C})$. It follows that $\theta\not\in D$, therefore $B(\sigma(\theta),\alpha_{i})<0$ for some simple root $\alpha_{i}$.
We prove (ii) by induction on $p$. If $p=0$, then $\theta=\theta_{C}$ by Proposition 5.1.2 (iii) and the conclusion of (ii) holds trivially. If $p>0$, then $\theta\neq\theta_{C}$ and there exists a simple root $\alpha_{i}$ with $B(\sigma(\theta),\alpha_{i})<0$ by (i). The simple reflection $s_{i}$ satifies the condition (5.4.1), adds 2 to the height of the sum, and sends $\theta$ to an $n$-root of type $A^{p-1}C^{q+1}$ by Remark 5.4.2, so (ii) follows by induction.
Let $\mathbf{w}=s_{i_{1}}\cdots s_{i_{r}}$ be a $(\theta,\theta_{C})$-word expressing a $(\theta,\theta_{C})$ element $w$. Then $w$ takes $\theta_{C}$ to $\theta$, and we have $r=\lambda(\theta_{C})-\lambda(\theta)=p$ by by Remark 5.4.2. Remark 5.4.3 then implies that $\mathbf{w}$ is reduced. It follows that $\ell(w)=r=p$. Since an element taking $\theta_{C}$ to $\theta$ has length at least $\lambda(\theta_{C})-\lambda(\theta)=p$ by Remark 5.4.3 and $w$ is such a shortest element, it also follows that every shortest element taking $\theta_{C}$ to $\theta$ has length $p$.
To prove (iii), it remains to show that $\mathbf{w}$ cannot contain a factor of the form $s_{\alpha_{i}}s_{\alpha_{j}}s_{\alpha_{i}}$. By Remark 5.4.2 and direct computation, such a factor would imply the existence of a subsequence $(\theta_{a},\theta_{a+1},\theta_{a+2},\theta_{a+3})$ such that
$s_{\alpha_{i}}(\theta_{a+2})=s_{\alpha_{i}}(\theta_{a}+2\alpha_{i}+2\alpha_{j})=\theta_{a}+2\alpha_{i}+2\alpha_{j}=\theta_{a+2}.$
This contradicts the fact that $s_{\alpha_{2}}(\theta_{a+2})=\theta_{a+3}$, which completes the proof of (iii).
Let $w$ be a shortest element taking $\theta_{C}$ to $\theta$. Then $w$ has length $p=\lambda(\theta_{C})-\lambda(\theta)$ by (iii). Let $\mathbf{w}=s_{i_{1}}\dots s_{i_{p}}$ be a reduced word for $w$. Remark 5.4.3 implies that if we start from $\theta_{C}$ and apply $s_{i_{p}},\cdots,s_{i_{2}},s_{i_{1}}$ successively, each simple reflection must be a $CA$ move. In particular, $s_{i_{p}}$ must perform a $CA$ move on $\theta_{C}$, so we have
$s_{i_{p}}(\sigma(\theta_{C}))=\sigma(s_{i_{p}}(\theta_{C}))=\sigma(\theta_{C})-2\alpha_{i_{p}}$
by Corollary 4.2.4 (ii). This implies that $B(\sigma(\theta_{N}),\alpha_{i_{p}})=B(\sigma(\theta_{C}),\alpha_{i_{p}})>0$, so it follows from Proposition 5.1.1 (iii) that $\alpha_{i_{p}}=\alpha_{x}$ where $\alpha_{x}$ is the unique Coxeter generator of $W$ not in $I$. In other words, every reduced word for $w$ ends in $s_{\alpha_{x}}$. It follows from [26, Proposition 1.10 (c)] that $w$ is the unique shortest element in $wW_{I}$.
If $w^{\prime}$ is another shortest element taking $\theta_{C}$ to $\theta$, then $w^{\prime}w^{-1}(\theta_{C})=\theta_{C}$, so $w^{\prime}w^{-1}\in W_{I}$ by Proposition 5.1.1 (iv). It follows that the cosets $wW_{I}$ and $w^{\prime}W_{I}$ are equal, therefore we have $w=w^{\prime}$, because $w$ and $w^{\prime}$ are both the unique shortest element in the common coset $wW_{I}=w^{\prime}W_{I}$ by the last paragraph. This proves the uniqueness of the shortest element taking $\theta_{C}$ to $\theta$. Part (iii) says that each $(\theta,\theta_{C})$-element is such a shortest element, and (iv) now follows.
Finally, to prove (v), we recall from Proposition 4.3.1 (i) that $\theta_{A}$ is the unique minimal element of the set $X$. It follows from [37, Theorem 2.8] that there exists an element $u\in W$ such that $u(\theta_{A})=\theta$ and $\ell(u)=\lambda(\theta)-\lambda(\theta_{A})=q$. Let $v$ be the unique $(\theta,\theta_{C})$-element, let $s_{i_{1}}\cdots s_{i_{p}}$ be a reduced word for $v$, and let $s^{\prime}_{i_{1}}\cdots s^{\prime}_{i_{q}}$ be a reduced word for $u^{-1}$. Then $u^{-1}v$ takes $\theta_{C}$ to $\theta_{A}$ and has length at most $p+q=M=\lambda(\theta_{C})-\lambda(\theta_{A})$, so $u^{-1}v$ must be the unique $(\theta_{A},\theta_{C})$-element by (ii) and (iii), and the word $\mathbf{w}=s^{\prime}_{i_{1}}\cdots s^{\prime}_{i_{q}}s_{i_{1}}\cdots s_{i_{p}}$ must be a reduced word for $u^{-1}v$. Remark 5.4.3 now implies that starting from $\theta_{A}$ and appling $s^{\prime}_{i_{1}},\cdots,s^{\prime}_{i_{q}},s_{i_{1}},\cdots,s_{i_{p}}$ successively yields a nonnesting sequence $\theta_{\mathbf{w}}$ from $\theta_{A}$ to $\theta_{C}$ that reaches $\theta=v\theta_{C}$ after the first $q$ steps, and (v) follows. ∎
Theorem 5.4.5.
Let $W$ be a Weyl group of rank $n$ of types $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even. Let $M$ be the number of coplanar quadruples in a maximal orthogonal set, and let $W_{I}$ be the parabolic subgroup of Proposition 5.1.2 (i).
  • (i)
    There is a unique element $w_{N}\in W$ of minimal length such that $w_{N}(\theta_{C})=\theta_{A}$. The element $w_{N}$ is fully commutative and has length $\ell(w_{N})=M$, and is the unique element of minimal length in the coset $w_{N}W_{I}$.
  • (ii)
    The set
    $L=\{v(\theta_{C}):v\leq_{L}w_{N}\}$
    is a complete, irredundantly described set of nonnesting positive $n$-roots. The set $L$ has the structure of a distributive lattice, induced by the weak Bruhat order $\leq_{L}$.
  • (iii)
    If $\gamma_{1}$ and $\gamma_{2}$ are positive $n$-roots satisfying $\text{\rm ht}(\sigma(\gamma_{1}))-\text{\rm ht}(\sigma(\gamma_{2}))=2M$, and $w$ is an element expressed by a word $\mathbf{w}$ of length $M$ satisfying $w(\gamma_{1})=\pm\gamma_{2}$, then $\mathbf{w}$ is reduced, and we must have $\gamma_{1}=\theta_{C}$, $\gamma_{2}=\theta_{A}$, $w=w_{N}$, and $w(\gamma_{1})=\gamma_{2}$.
Proof.
Part (i) follows from Proposition 5.4.4 in the case where $\theta_{0}=\theta_{A}$.
Let $w_{N}\in W$ be as in (i) and let $v\leq_{L}w_{N}$. We may complete a reduced word $\mathbf{v}$ for $v$ to a reduced word of the form $\mathbf{w}=\mathbf{u}\cdot\mathbf{v}$ for $w_{N}$. Remark 5.4.3 implies that $\mathbf{w}$ gives rise to a nonnesting sequence $\theta_{\mathbf{w}}$ from $\theta_{A}$ to $\theta_{C}$ that passes $v(w_{N})$. It follows that the elements of $L$ are indeed all nonesting positive $n$-roots. Conversely, for every nonnesting positive root $\theta$, it follows from Proposition 5.4.4 (v) and its proof that $\theta=v(\theta_{C})$ for some element $v\leq_{L}w_{N}$, so the list $L$ is complete. Finally, if $v\leq_{L}w_{N}$ and $v^{\prime}\leq_{L}w_{N}$ are elements such that $v(\theta_{C})=v^{\prime}(\theta_{C})$, then $v^{\prime}v^{-1}$ stabilizes $\theta_{C}$ and hence $\sigma(\theta_{C})$, so we have $v^{\prime}v\in W_{I}$ and $v^{\prime}W_{I}=vW_{I}$. Since $w_{N}$ is the shortest element in $w_{N}W_{I}$, the elements $v^{\prime}$ and $v$ must be the shortest elements in $vW_{I}$ and $v^{\prime}W_{I}$ as well, which implies that $v=v^{\prime}$ as well. It follows that the list $L$ irredundantly describes the positive nonnesting $n$-roots, proving the first statement of (ii).
By [47, Theorem 3.2], the fact that $w_{N}$ is fully commutative implies that the poset $\{x:x\leq_{L}w\}$ is a distributive lattice. This completes the proof of (ii).
Suppose that the conditions of (iii) hold, and let $s_{i_{1}}s_{i_{2}}\cdots s_{i_{M}}$ be a reduced expression for $w$. Since $\text{\rm ht}(\sigma(\gamma_{1}))-\text{\rm ht}(\sigma(\gamma_{2}))=2M$, as we start from $\theta_{C}$ and successively apply the simple reflection $s_{i_{M}},\dots,s_{i_{2}},s_{i_{1}}$, the application of each simple reflection $s_{i_{j}}$ must subtract $2$ from the height of the sum and change a $C$ to an $A$ by Corollary 4.2.4 (ii). It is therefore not possible at any stage for a simple reflection to negate a component of an $n$-root, which implies that we have $w(\gamma_{1})=\gamma_{2}$. Each simple reflection $s_{i_{j}}$ also causes no change in the number of nestings by Proposition 4.2.2 (i), so the fact that $\text{\rm ht}(\sigma(\gamma_{1}))-\text{\rm ht}(\sigma(\gamma_{2}))=2M$ implies that $\gamma_{1}$ and $\gamma_{2}$ have types $C^{M}$ and $A^{M}$, respectively, so we have $\gamma_{1}=\theta_{C}$ and $\gamma_{2}=\theta_{A}$. We then have $w=w_{N}$ by (i), which completes the proof of (iii). ∎
We call the element $w_{N}$ from Theorem 5.4.5 the nonnesting element of $W$. In Section 6, we will compute $w_{N}$ explicitly with the help of Theorem 5.4.5 (iii).
5.5. Sum equivalence
We say that two positive $n$-roots $\beta$ and $\gamma$ of $W$ are sum equivalent, or $\sigma$-equivalent, if $\sigma(\beta)=\sigma(\gamma)$. If $C$ and $C^{\prime}$ are two $\sigma$-equivalence classes, then we write $C\leq_{\sigma}C^{\prime}$ if $\sigma(\beta)\leq\sigma(\gamma)$ for any $\beta\in C$ and $\gamma\in C^{\prime}$. The goal of this subsection is to show that the $\sigma$-equivalence classes of $X$ are highly compatible with the quasiparabolic order $\leq_{Q}$ and the feature-avoiding $n$-roots.
Proposition 5.5.1.
Let $\mathcal{B}$ be the set of nonnesting positive $n$-roots of $W$.
  • (i)
    If $\beta,\beta^{\prime}\in\mathcal{B}$ are nonnesting positive $n$-roots with $\sigma(\beta)=\sigma(\beta^{\prime})$, then we have $\beta=\beta^{\prime}$.
  • (ii)
    Each positive $n$-root $\gamma$ is $\sigma$-equivalent to a unique nonnesting $n$-root $f(\gamma)$ and a unique noncrossing $n$-root $g(\gamma)$. We have $f(\gamma)\leq_{Q}\gamma$, and
    $\gamma=f(\gamma)+\sum_{\beta\in\mathcal{B}:\sigma(\beta)<\sigma(\gamma)}\lambda_{\beta,\gamma}\beta$
    for suitable integers $\lambda_{\beta,\gamma}$.
  • (iii)
    Every $\sigma$-equivalence class contains a unique nonnesting $n$-root, $\beta_{1}$, and a unique noncrossing $n$-root, $\beta_{2}$. The $\sigma$-equivalence class containing $\beta_{1}$ and $\beta_{2}$ is the interval
    $[\beta_{1},\beta_{2}]=\{\gamma\in X:\beta_{1}\leq_{Q}\gamma\leq_{Q}\beta_{2}\}$
    in the quasiparabolic set $X$.
  • (iv)
    The set of alignment-free positive $n$-roots is a $\sigma$-equivalence class and is equal to the interval $[\theta_{C},\theta_{N}]$ in the quasiparabolic set $X$. It is the top $\sigma$-equivalence class with respect to the partial order $\leq_{\sigma}$.
Proof.
Suppose that $\beta$ and $\beta^{\prime}$ are nonnesting positive $n$-roots with $\sigma(\beta^{\prime})=\sigma(\beta)$. It follows from Proposition 5.4.4 (ii) that $\beta$ and $\beta^{\prime}$ have the same number of alignments, namely, the number $p=(\text{\rm ht}(\sigma(\theta_{C}))-\text{\rm ht}(\sigma(\theta)))/2$. If $p=0$, then we have $\beta=\theta_{C}=\beta^{\prime}$ by Proposition 5.1.2 (iii). If $p>0$, then neither $\beta$ nor $\beta^{\prime}$ equals $\theta_{C}$, so there is a simple root $\alpha_{i}$ satisfying $B(\sigma(\beta),\alpha_{i})=B(\sigma(\beta^{\prime}),\alpha_{i})<0$ by Proposition 5.4.4 (i). By Remark 5.4.2, both $s_{i}(\beta)$ and $s_{i}(\beta^{\prime})$ are nonnesting positive $n$-roots with $p-1$ alignments, and we have
$\sigma(s_{i}(\beta))=s_{i}(\sigma(\beta))=\sigma(\beta)+2\alpha_{i}=\sigma(\beta^{\prime})+2\alpha_{i}=s_{i}(\sigma(\beta^{\prime}))=\sigma(s_{i}(\beta^{\prime})),$
therefore (i) follows by induction on $p$.
Let $\gamma$ be a positive $n$-root, and let $\leq_{N}$ be the nesting order defined in Section 5.2. If $\gamma$ contains no nesting, we can simply take $f(\gamma)=\gamma$. Otherwise, we can factorize $\gamma=\gamma^{\prime}\gamma_{N}$ where $\gamma_{N}$ is a nesting. By the second paragraph of Section 5.2, we can write $\gamma=\gamma^{\prime}\gamma_{C}-\gamma^{\prime}\gamma_{A}$ where we have (a) $\gamma^{\prime}\gamma_{C}\leq_{N}\gamma$, because $\sigma(\gamma^{\prime}\gamma_{C})=\sigma(\gamma)$ and $\gamma^{\prime}\gamma_{C}<_{Q}\gamma$, and (b) $\gamma^{\prime}\gamma_{A}\leq_{N}\gamma$, because $\sigma(\gamma^{\prime}\gamma_{A})<\sigma(\gamma)$. Taking $f(\gamma)=f(\gamma^{\prime}\gamma_{C})$ proves the existence of $f(\gamma)$ and the required expression for $\gamma$ by induction on the order $\leq_{N}$. The uniqueness of $f(\gamma)$ follows from (i). We can use a similar induction using the crossing order $\leq_{C}$ to show that $\gamma$ is $\sigma$-equivalent to a noncrossing $n$-root $g(\gamma)$ such that $\gamma\leq_{Q}g(\gamma)$, and this completes the proof of (ii).
It follows from (ii) that every $\sigma$-equivalence class contains a unique nonnesting $n$-root, and that the number of $\sigma$-equivalence classes equals the number of nonnesting $n$-roots. The latter number is the dimension of the Macdonald representation and also the number of noncrossing roots by Theorem 5.2.2 (i) and (ii). Since each $\sigma$-equivalence class contains at least one noncrossing $n$-root by (ii), it follows that each $\sigma$-equivalence class must contain exactly one nonnesting element, and exactly one noncrossing element. This proves the first sentence of (iii).
Let $E$ be a $\sigma$-equivalence class with unique nonnesting element $\beta_{1}$ and unique noncrossing element $\beta_{2}$. Then we have $[\beta_{1},\beta_{2}]\subseteq E$ by Corollary 4.2.4 (i). Conversely, if $\gamma$ is an $n$-root in $E$, then (ii) and its proof imply that we may find a nonnesting $n$-root $f(\gamma)\in E$ and a noncrossing $n$-root $g(\gamma)\in E$ such that $f(\gamma)\leq_{Q}\gamma\leq_{Q}g(\gamma)$. We must have $f(\gamma)=\beta_{1}$ and $g(\gamma_{2})=\beta_{2}$ the uniqueness of the nonnesting and noncrossing elements in $E$, therefore $\gamma\in[\beta_{1},\beta_{2}]$. It follows that $E=[\beta_{1},\beta_{2}]$.
For every nonnesting root $\theta$ not equal to $\theta_{C}$, there is a nontrivial nonnesting sequence from $\theta$ to $\theta_{C}$ by by Proposition 5.4.4 (ii), so $\sigma(\theta)<\sigma(\theta_{C})$ by Definition 5.4.1. Part (iv) now follows from (iii) and Proposition 5.1.2 (ii)–(iii). ∎
Theorem 5.5.2.
Let $W$ be a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even. If $\mathcal{B}$ is any set of $\sigma$-equivalence class representatives, then $\mathcal{B}$ is a basis for the Macdonald representation $j^{\Phi}_{nA_{1}}({\rm{sgn}})$. The change of basis matrix between any two such bases is unitriangular with integer entries. In particular, this is true for the change of basis matrix between the nonnesting basis and the noncrossing basis.
Proof.
By Proposition 5.5.1 (ii), each element $\gamma\in\mathcal{B}$ is the sum of the nonnesting element $\gamma^{\prime}$ that is $\sigma$-equivalent to $\gamma$ and a ${\mathbb{Z}}$-linear combination of nonnesting elements with strictly lower sums. The nonnesting elements form a basis for $j^{\Phi}_{nA_{1}}({\rm{sgn}})$ by Theorem 5.2.2 (i), from which it follows that the set $\mathcal{B}$ is also a basis, and that the change of basis from $\mathcal{B}$ to the nonnesting basis is unitriangular with integer entries. If $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ are two such bases, then the change of basis matrix from $\mathcal{B}_{1}$ to $\mathcal{B}_{2}$ is unitriangular with integer entries because it is the product of the matrix changing $\mathcal{B}_{1}$ to $\mathcal{B}^{\prime}$ with the inverse of the matrix changing $\mathcal{B}_{2}$ to $\mathcal{B}^{\prime}$, both of which are unitriangular with integer entries. Finally, the last assertion follows because Proposition 5.5.1 (iii) implies that both the nonnesting and noncrossing bases are sets of $\sigma$-equivalence class representatives. ∎
Remark 5.5.3.
Recall that the Möbius function, $\mu$, of a partially ordered set $P$ is defined to satisfy $\mu(x,x)=1$, $\mu(x,y)=0$ if $x\not\leq y$, and
$\sum_{z:x\leq z\leq y}\mu(x,z)=0$
if $x<y$. A poset is Eulerian if we have $\mu(x,y)=(-1)^{\lambda(y)-\lambda(x)}$ whenever $x\leq y$. It can be shown that if $x,y\in X$ correspond to a nonnesting and noncrossing element, respectively, then the interval $I=[x,y]$ corresponds to a $\sigma$-equivalence class if and only if $I$ is Eulerian.
Remark 5.5.4.
Reading [38] defines a poset congruence to be an equivalence relation on a poset $X$ such that
  • (i)
    each equivalence class is an interval;
  • (ii)
    the projection mapping $x\in X$ to the maximal element in its equivalence class is order preserving; and
  • (iii)
    the projection mapping $x\in X$ to the minimal element in its equivalence class is order preserving.
It can be shown using [51, Proposition 42] that, in type $D_{n}$, the equivalence relation induced on $X$ by $\sigma$ is a poset congruence. It can also be shown that the same is true in types $E_{7}$ and $E_{8}$.
6. Examples
In this section, we give type-specific details about the $n$-roots in types $D_{n}$ for $n$ even, $E_{7}$, and $E_{8}$. In all types, we explicitly describe the maximally aligned, crossing, and nesting $n$-roots $\theta_{A},\theta_{C}$ and $\theta_{N}$. We find the nonnesting word $w_{N}$, and we use $w_{N}$ and Theorem 5.4.5 (ii) to deduce the dimension of the Macdonald representation $j^{\Phi}_{nA_{1}}({\rm{sgn}})$. We also discuss type-specific properties of the set $X_{I}$ of alignment-free positive $n$-roots for all types. In addition, we explain precise connections between the Macdonald representation $j^{\Phi}_{nA_{1}}({\rm{sgn}})$ of type $D_{2k}$ and a Specht module of the symmetric group $S_{2k}$ (Proposition 6.1.2),
We note that by Lemma 3.1.1 and Remark 3.2.3, the noncrossing and nonnesting positive $n$-roots of type $D_{2k}$ can be easily recovered from the well-studied noncrossing and nonesting perfect matchings of $[2k]$. More generally, in all types, the nonnesting positive $n$-roots can be computed efficiently via the elements the elements $\theta_{C}$ and $w_{N}$ by Theorem 5.4.5 (ii), and it is possible to construct the noncrossing $n$-roots using Fan’s construction of monomial cells in [20]. In the notation of [20], the maximally aligned $n$-root $\theta_{A}$ can be identified with the element $b_{1}b_{3}\cdots b_{2k-1}$ in type $D_{2k}$, with $b_{2}b_{4}b_{6}b_{7}$ in type $E_{7}$ (with the labelling of Figure 1 (d)), and with $b_{2}b_{3}b_{5}b_{7}$ in type $E_{8}$ (with the labelling of Figure 1 (e)). In types $E_{7}$ and $E_{8}$, it is also possible to use a computer program to find all noncrossing and nonnesting $n$-roots by generating all the (finitely many) positive $n$-roots and then removing all $n$-roots where a crossing or nesting can be found. For these reasons, and to save space, we have chosen not to list the noncrossing and nonnesting bases in type $E_{7}$ or $E_{8}$ in this paper (although the complete lists are available upon request).
6.1. Type $D_{2k}$
If $W$ has type $D_{n}$ for an even integer $n=2k$, then the positive $n$-roots can be naturally identified with the perfect matchings of $[n]$, as explained in Lemma 3.1.1 (ii). Under this identification, the actions of $W$ on the $n$-roots and on the matching agree, and the alignments, crossings, and nestings in the $n$-roots correspond to the alignments, crossing, and nestings in the matchings in the obvious way by Remark 3.2.3. We also recall from Section 2.3 and Remark 3.1.7 that the reflection $r=s_{\alpha}\in W$ acts as the transposition $(ij)$ on the $n$-roots for each root $\alpha=\varepsilon_{i}\pm\varepsilon_{j}$ of $W$, so that the action of $W$ factors through the homomorphism $\phi:W\rightarrow S_{2k}$ of Equation (2.3.1) to induce an action of $S_{2k}=W(A_{2k-1})$ on the $n$-roots, giving the Macdonald representation $j^{\Phi}_{nA_{1}}({\rm{sgn}})$ the structure of an $S_{2k}$-module (where the elements of $S_{2k}$ permute the indices of the terms $\varepsilon_{i}^{2}$). The above facts will allow us to connect the theory of $n$-roots in type $D_{n}$ to some widely studied type-$A$ objects and results.
Recall that the number of coplanar quadruples in each positive $n$-root is $M=\binom{k}{2}$, the number of pairs of 2-blocks, by Corollary 3.3.5.
Let $\nu_{A},\nu_{C},$ and $\nu_{N}$ be the positive $n$-roots corresponding to the matchings $\{12,34,\cdots,(n-1)n\},\{1(k+1),2(k+2),\cdots,k(2k)\}$ and $\{1n,2(n-1),\cdots,k(k+1)\}$, respectively. Every pair of 2-blocks in the first matching form an alignment, so the matching contains $\binom{k}{2}=M$ alignments. It then follows from Proposition 4.3.1 (i) that $\nu_{A}$ is the unique maximally aligned $n$-root $\theta_{A}$ in the set $X$. Similar arguments show that $\nu_{C}=\theta_{C}$ and $\nu_{N}=\theta_{N}$ by Proposition 5.1.2 (iii) and Proposition 4.3.1 (ii), respectively. Note that we have $\sigma(\theta_{N})=2\sum_{i=1}^{k}\varepsilon_{i}$.
Let $w$ be the element expressed by the word
$\mathbf{w}=\mathbf{w}_{2,k-2}\mathbf{w}_{3,k-3}\cdots\mathbf{w}_{k,0},$
where $\mathbf{w}_{i,j}:=s_{i}s_{i+2}s_{i+4}\cdots s_{i+2j}$. For example, in type $D_{8}$, we have $w=(s_{2}s_{4}s_{6})(s_{3}s_{5})(s_{4})$, and the heap of $w$ is shown in Figure 2 (a). The word $\mathbf{w}$ has $M$ letters, and it is straightforward to verify that $w(\theta_{C})=\theta_{A}$, so it follows from Theorem 5.4.5 (i) that $w$ is the fully commutative nonnesting element $w_{N}$ and $\mathbf{w}$ is a reduced word for it.
$s_{2}$$s_{4}$$s_{6}$$s_{3}$$s_{5}$$s_{4}$
$s_{1}$$s_{3}$$s_{5}$$s_{2}$$s_{4}$$s_{3}$$s_{7}$
$s_{1}$$s_{4}$$s_{6}$$s_{8}$$s_{3}$$s_{5}$$s_{7}$$s_{4}$$s_{6}$$s_{5}$$s_{2}$$s_{4}$$s_{3}$$s_{1}$
Figure 2. The heaps of the nonnesting elements of types $D_{8}$, $E_{7}$, and $E_{8}$
Since $w_{N}=w$ is fully commutative, the elements in the set $\{v\in W:v\leq_{L}w_{N}\}$ are in bijection with the order filters of the heap poset of $w_{N}$. (See [47, Section 2.2] for the definition of the heap poset; an order filter of a poset $P$ is a subset of $P$ such that $y\in I$ whenever the conditions $y\in P$, $x\in I$, and $x\leq y$ hold.) These filters are in canonical correspondence with Dyck paths of order $k$, i.e., staircase walks from $(0,0)$ to $(k,k)$ that lie strictly below (but may touch) the diagonal $y=x$. It is well known [46, Theorem 1.5.1 (vi)] that the number of such paths is the $k$-th Catalan number, $C_{k}=\frac{1}{k+1}\binom{2k}{k}$. Theorem 5.4.5 (ii) and Theorem 5.2.2 imply that the number of nonnesting positive $n$-roots of $W$ is given by $C_{k}$, as are the number of noncrossing positive $n$-roots and the dimension of the Macdonald representation $j^{\Phi}_{nA_{1}}({\rm{sgn}})$.
The level function $\lambda$ in type $D_{2k}$ has a combinatorial interpretation that is natural in the context of combinatorial game theory [29]. The matching corresponding to an $n$-root $\beta$ can be identified with a Steiner system $S(1,2,2k)$, i.e., a collection of 2-blocks of $[2k]$ with the property that any singleton lies in a unique 2-block. The level $\lambda(\beta)$ then counts the number of 2-element subsets $E$ of $[2k]$ with the property that the matching corresponding to $\beta$ contains no 2-blocks of the form $(E\setminus\{j\})\cup\{i\}$ where $i\leq j$ and $j\in E$ (in particular, the matching cannot contain $E$). With some more work, it can be shown that each crossing gives rise to one such subset $E$, and each nesting gives rise to two such subsets. This gives a combinatorial interpretation of the formula $\lambda(x)=C(x)+2N(x)$, and also explains the appearance of the product of odd quantum integers in [29, Equation (4.2)]. In addition, the quantity $C(m)+2N(m)$ associated to each matching $m$ appears as “$C(m)+2U(m)$” in the context of octabasis Laguerre polynomials in [45, Sections 4 and 5], as the weight “$\omega(m)$” in the context of Gaussian $q$-distributions in [19, Theorem 4], and as “$cov(m)-cro(m)$” in the context of $q$-Bessel numbers in [14, Section 4].
The poset structure on the set $X_{I}$ in type $D_{2k}$ coincides with a familiar one.
Proposition 6.1.1.
Suppose $W$ has type $D_{2k}$. Then as a poset, the interval $X_{I}=[\theta_{C},\theta_{N}]$ in the quasiparabolic set $X$ is canonically isomorphic to the symmetric group $S_{k}$ under the (strong) Bruhat order via the map $\varphi:S_{k}\rightarrow X_{I}$ sending each element $\tau\in S_{n}$ to the $n$-root
(6.1.1) $\varphi(\tau)=\prod_{i=1}^{k}(\varepsilon_{i}^{2}-\varepsilon_{\tau(i)+k}^{2}).$
Under this bijection, we have
$\lambda(\varphi(\tau))=M+\ell(\tau)$
for every $\tau\in S_{k}$, where $M$ is the number of coplanar quadruples in each $n$-root and $\ell$ denotes Coxeter length.
Proof.
By Proposition 5.1.2 (ii), the set $X_{I}$ is the $\sigma$-equivalence class of the $n$-root $\theta_{N}$. We noted earlier that $\sum(\theta_{N})=2\sum_{i=1}^{k}\varepsilon_{i}$, which implies that $n$-roots in $X_{I}$ are precisely the positive $n$-roots whose components are all of the form $\varepsilon_{i}\pm\varepsilon_{j}$, where $1\leq i\leq k<j\leq 2k$. These are precisely the $n$-roots listed in the theorem, so the map $\varphi$ is surjective. It is clear that $\varphi$ is also injective, so $\varphi$ is a bijection.
The Bruhat order on $S_{k}$ is generated by relations of the form $\tau<r\tau$ where we have $\tau\in S_{k}$ and $r$ is a reflection $r=(\tau(i),\tau(j))\in S_{k}$ for some $i<j$ such that $\tau(i)<\tau(j)$ [26, Section 5.9, Example 2]. In this case, the quadruple $\{\varepsilon_{i}\pm\varepsilon_{\tau(i)+k},\varepsilon_{j}\pm\varepsilon_{\tau(j)+k}\}$ contained in $\varphi(\tau)$ is a crossing and is moved to the nesting $\{\varepsilon_{i}\pm\varepsilon_{\tau(j)+k},\varepsilon_{j}\pm\varepsilon_{\tau(i)+k}\}$, so we have $\lambda(\varphi(\tau))<\lambda(r\varphi(\tau))$ in $X_{I}$ by Proposition 4.2.2 (iii). Conversely, if we have $\lambda(\varphi(\tau))<\lambda(r\varphi(\tau))$ in $X_{I}$ for some $\tau\in S_{k}$ and some reflection $r\in W$, then since $\varphi(\tau)$ has no alignments, $r$ must move a crossing in $\varphi(\tau)$ to a nesting by Proposition 4.2.2 (iii). The crossing moved must be of the form $\{\varepsilon_{i}\pm\varepsilon_{\tau(i)+k},\varepsilon_{j}\pm\varepsilon_{\tau(j)+k}\}$ for some $i,j\in[k]$ such that $i<j$ and $\tau(i)<\tau(j)$, and the only possibilities for $r$ are $(ij)$ and $(\tau(i)+k,\tau(j)+k)$. In either case, we have $r\varphi(\tau)=\varphi(r^{\prime}\tau)$ for the reflection $r^{\prime}=(\tau(i),\tau(j))\in S_{k}$, so that we have
$\varphi^{-1}(\varphi(\tau))=\tau<r^{\prime}\tau=\varphi^{-1}(r\varphi(\tau))$
where $<$ denotes the Bruhat order in $S_{k}$. It now follows that $\varphi$ is a poset isomorphism.
To prove the last assertion, we note that each inversion of a permutation $\tau\in S_{k}$ corresponds to a nesting in the corresponding alignment-free $n$-root $\gamma=\varphi(\tau)$, and we recall that $\ell(\tau)$ equals the number of inversions in $\tau$. It follows that $N(\gamma)=\ell(\tau)$, therefore we have
$\lambda(\gamma)=C(\gamma)+2N(\gamma)=(C(\gamma)+N(\gamma))+N(\gamma)=M+N(\gamma)=M+\ell(\tau).\qed$
We now discuss the structure of the space $j^{D_{n}}_{nA_{1}}({\rm{sgn}})$ underlying the Macdonald reprsentation as an $S_{2k}$-module. As a vector space, $j^{D_{n}}_{nA_{1}}({\rm{sgn}})$ is isomorphic to the free vector space on the noncrossing perfect matchings of $[n]=[2k]$, which is denoted by $V(n,k,0)$ in the work of Rhoades [41]. Furthermore, given a simple reflection $s_{i}=(i,i+1)\in S_{n}$ and a noncrossing perfect matching $m$ corresponding to an $n$-root $\gamma$, the reflection $s_{i}$ acts on $m$ in one of the following ways:
  1. (1)
    if $i(i+1)$ is a 2-block in $m$, then the root $\gamma$ contains $\varepsilon_{i}^{2}-\varepsilon_{i+1}^{2}$ as factor, so $s_{i}(m)=-m$;
  2. (2)
    if $i(i+1)$ if not a 2-block in $m$, then $m$ contains two blocks $ia$ and $(i+1)b$ which either form an alignment (if $a<i<i+1<b$) or a nesting (if $b<a<i$ or $i+1<b<a$). In all cases, we have $s_{i}(m)=m^{\prime\prime}$ where $m^{\prime\prime}$ is the matching $(m\setminus\{ia,(i+1)b\})\cup\{(i+1)a,ib\}$. Here, the blocks $(i+1)a$ and $ib$ form a crossing, and the Ptolemy relation $\gamma_{C}=\gamma_{N}+\gamma_{A}$ from Theorem 3.2.2 (vi) implies that
    (6.1.2) $s_{i}(m)=m^{\prime\prime}=m+m^{\prime}$
    where $m^{\prime}$ is the perfect matching $m^{\prime}=(m\setminus\{ia,(i+1)b\})\cup\{i(i+1),ab\}$. Here, the blocks $i(i+1)$ and $ab$ form the nesting in the Ptolemy relation if $\{ia,(i+1)b\}$ is an alignment and form the alignment in the Ptolemy relation if $\{ia,(i+1)b\}$ is a nesting. The matching $m^{\prime}$ is noncrossing by Theorem 5.3.2 (iv).
It follows from the above analysis that the action of $S_{2k}$ on $j^{D_{n}}_{nA_{1}}({\rm{sgn}})$ agrees with the action of $S_{2k}$ on the space $V(n,k,0)$ defined by Rhoades. The precise formula in Equation (6.1.2) appears in the work of Kim [31, Equation (1.3)]. By [41, Proposition 5.2], as an $S_{2k}$ module $V(n,k,0)$ is isomorphic to the Specht module $S^{(k,k)}$ corresponding to the 2-row partition $(k,k)$, so we may summarize our discussion as follows:
Proposition 6.1.2.
If $W$ has type $D_{n}$ for $n=2k$ even, then the action of $W$ on the Macdonald representation $j^{D_{n}}_{nA_{1}}({\rm{sgn}})$ factors through the map $\phi$ defined by Equation (2.3.1) to induce an $S_{n}$-module structure on $j^{D_{n}}_{nA_{1}}({\rm{sgn}})$. The resulting $S_{n}$-module is isomorphic to the Specht module $S^{(k,k)}\cong V(n,k,0)$. In particular, it is irreducible.
Remark 6.1.3.
The nonnesting and noncrossing bases for the $S_{n}$-module $j^{D_{n}}_{nA_{1}}({\rm{sgn}})\cong S^{(k,k)}$ are also studied extensively in the works of Russelll–Tymoczko [42], Im–Zhu [28], Hwang–Jang–Oh [27], and Heard–Kujawa [25]. In these papers, the noncrossing basis is called the web basis, and the nonnesting basis can be naturally identified with the standard basis (or the polytabloid or Specht basis) as explained in [28, Lemma 3.1] and [27, Section 1]. Under this identification, the isomorphism of [42, Theorem 2.2] associates each nonnesting perfect matching with the unique noncrossing matching in the same $\sigma$-equivalence class, and Theorem 5.5 of [42] follows from Theorem 5.5.2 as a special case. The restriction of the quasiparabolic order to the noncrossing basis gives rise to the web graph of [42, Section 2.3], which therefore has the structure of a distributive lattice by Theorem 5.4.5 (ii). Our definition of the nesting number (Definition 4.1.2 (i)) agrees with the nesting number of [42] when restricted to noncrossing $n$-roots, and is inspired by [42]. It also follows from [27, Corollary 4.2] that if $W$ has type $D_{2k}$ and we expand the maximally crossing $n$-root $\theta_{C}$ as a linear combination of the noncrossing basis, $\theta_{C}=\sum\lambda_{\beta}\beta$, then the sum $\sum\lambda_{\beta}$ of the nonnegative integers $\lambda_{\beta}$ is given by the number $E_{k+1}$ in the family (1,1,1,2,5,16,272, …) of Euler numbers, which are characterized by the equation
$\sec(x)+\tan(x)=\sum_{i=0}^{\infty}E_{i}\frac{x^{i}}{i!}.$
Coefficients in the expansion of the maximally nesting $2k$-root $\theta_{N}$ into the noncrossing basis have a combinatorial interpretation in terms of the so-called web permutations in $S_{k}$ by [27, Theorem 1.2].
6.2. Type $E_{7}$
Suppose $W$ has type $E_{7}$. We define $\nu_{A}$ to be the positive $7$-root with the following components:
$\alpha_{2},\ \alpha_{4},\ \alpha_{6},\ \alpha_{7},\ ,\alpha_{2}+2\alpha_{3}+\alpha_{4}+\alpha_{7},$
[…]
We define $\nu_{C}$ to be the positive $7$-root with the following components:
$\alpha_{2}+\alpha_{3}+\alpha_{4}+\alpha_{5}+\alpha_{7},\quad\alpha_{2}+2\alpha_{3}+\alpha_{4}+\alpha_{7},\quad\alpha_{3}+\alpha_{4}+\alpha_{5}+\alpha_{6}+\alpha_{7},\quad\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}+\alpha_{7}$
$\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}+\alpha_{5}+\alpha_{7},\quad\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4}+\alpha_{5}+\alpha_{6}+\alpha_{7},\quad\alpha_{1}+2\alpha_{2}+3\alpha_{3}+3\alpha_{4}+2\alpha_{5}+\alpha_{6}+\alpha_{7}.$
We define $\nu_{N}$ to be the positive 7-root with the following components:
$\alpha_{7},\quad\alpha_{2}+2\alpha_{3}+\alpha_{4}+\alpha_{7},\quad\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}+\alpha_{5}+\alpha_{7},\quad\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4}+\alpha_{5}+\alpha_{6}+\alpha_{7},$
$\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4}+\alpha_{5}+\alpha_{6}+\alpha_{7},\quad\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4}+\alpha_{5}+\alpha_{7},\quad\alpha_{2}+2\alpha_{3}+2\alpha_{4}+2\alpha_{5}+\alpha_{6}+\alpha_{7}.$
Finally, we define the element $w\in W$ to be the element expressed by the word
$\mathbf{w}=(s_{1}s_{3}s_{5})(s_{2}s_{4})(s_{3})(s_{7}).$
The heap of $w$ is shown in Figure 2 (b).
Proposition 6.2.1.
If $W$ has type $E_{7}$, then the $7$-roots $\nu_{A}$, $\nu_{C}$ and $\nu_{N}$ given above are respectively the maximally aligned, maximally crossing, and maximally nesting $n$-roots of $W$. The element $w$ is the nonnesting element, and $\mathbf{w}$ a reduced word for it. The Macdonald representation $j^{E_{7}}_{7A_{1}}({\rm{sgn}})$ has dimension $15$.
Proof.
Recall from Corollary 3.3.5 that the number of coplanar quadruples in any $7$-root is $M=7$. Direct verification shows that $\text{\rm ht}(\sigma(\nu_{C}))-\text{\rm ht}(\sigma(\nu_{A}))=49-35=2M$ and that $w(\nu_{C})=\nu_{A}$, which implies the assertions about $\nu_{A}$, $\nu_{C}$, $\mathbf{w}$ and $w$ by Theorem 5.4.5 (iii). The dimension of the Macdonald representation $j^{E_{7}}_{7A_{1}}({\rm{sgn}})$ equals the cardinality of the set $\{v\in W:v\leq_{L}w\}$ by Theorem 5.2.2 (i) and Theorem 5.4.5 (ii). As explained in Section 6.1, this set is in bijection with the order filters of the heap of the fully commutative $w_{N}$, and direct computation shows that this heap has 15 filters, so the dimension of $j^{E_{7}}_{7A_{1}}({\rm{sgn}})$ is 15.
It remains to show that $\nu_{N}$ is the maximally aligned $n$-root $\theta_{A}$. By inspection, the heights of the components of $\nu_{N}$ are 1, 5, 7, 9, 9, 9, 9 when listed in increasing order. The sum of the first three terms of this sequence is bigger than the largest term, so $\nu_{N}$ cannot contain any alignments by Proposition 3.2.5 (ii). If $\nu_{N}$ contains a crossing, it follows from Proposition 3.2.5 (iv) that this crossing cannot contain the root of height 1, and that it can contain at most one root of height 9. There is no subset of four roots with this property, so $\nu_{N}$ contains no crossing either. It follows that all the $M$ coplanar quadruples in $\nu_{N}$ are nestings, so that $\nu_{N}=\theta_{N}$ by Proposition 4.3.1 (ii). ∎
The set $X_{I}$ of alignment-free positive $n$-roots in type $E_{7}$ is intimately related to the combinatorics of the Fano plane (Figure 3), the finite projective plane of order 2 over the field $\mathbb{F}_{2}$ with two elements. We recall that any two points in the Fano share a unique line that contains them both, so that the vertex labellings of the Fano points using the labels 1,2,..,71,2,..,71 , 2 , . . , 7 correspond precisely to the Steiner triple systems $S(2,3,7)$ via the bijection that associates each line in the Fano plane with the triple of labels for the vertices in that line. For example, the labellings shown in Figure 3 (a) and (b) correspond respectively to the Steiner systems $L_{C}$ and $L_{N}$ from Proposition 6.2.2. It is well known that the automorphism group of the Fano plane is the simple group $GL(3,2)$ of order $168$, so that the number of inequivalent vertex labellings is $7!/168=30$.
5555666622224444333377771111
(a) $L_{C}$
6666555522227777444433331111
(b) $L_{N}$
Figure 3. The inequivalent labellings of the Fano plane corresponding to $\theta_{C}$ and $\theta_{N}$
Proposition 6.2.2.
If $W$ has type $E_{7}$, then every component of every 7-root $\gamma\in X_{I}$ has the form
$\eta_{abc}=\left(\sum_{i=0}^{7}\varepsilon_{i}\right)-2(\varepsilon_{0}+\varepsilon_{a}+\varepsilon_{b}+\varepsilon_{c})$
for a 3-element subset $abc:=\{a,b,c\}$ of the set [7], and the map $\varphi$ sending each 7-root $\gamma\in X_{I}$ to the set
$L_{\gamma}=\{abc:\eta_{abc}\,|\,\gamma\}$
gives a canonical bijection from $X_{I}$ to the 30 inequivalent labellings of the Fano plane. Under this bijection, the minimal element $\theta_{C}$ of $X_{I}$ corresponds to the labelling
$L_{C}=L_{\theta_{C}}=\{136,145,127,235,246,347,567\},$
and the maximal element $\theta_{N}$ corresponds to the labelling
$L_{N}=L_{\theta_{N}}=\{123,145,246,257,347,356,167\}.$
Proof.
In the Fano coordinates, the components of the maximally crossing and maximally nesting $n$-roots $\theta_{C}$ and $\theta_{N}$ are given by the rows of the following matrices $M_{C}$ and $M_{N}$, respectively:
(6.2.1) $M_{C}=\left[\begin{matrix}{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}\\ {\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}\\ {\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}\\ {\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}\\ {\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}\\ {\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}\\ {\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{--1}}\\ \end{matrix}\right],\;M_{N}=\left[\begin{matrix}{\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}\\ {\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}\\ {\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}\\ {\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}\\ {\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}\\ {\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}\\ {\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}\\ \end{matrix}\right].$
By inspection, the components of $\theta_{C}$ have the properties (a) each of them is a root of the form $\eta_{abc}$ for some triple $abc\subseteq[7]$, and (b) the triples corresponding to components form a Steiner triple system.
By Proposition 6.2.1, the rightmost generator appearing in $w_{N}$ is $s_{7}$, which implies that $I=S\backslash\{s_{7}\}$. It follows from Section 2.3 that $W_{I}$ is a Weyl group of type $A_{6}$, isomorphic to $S_{7}$, and that $W_{I}$ acts on $X_{I}$ by permuting the Fano coordinates. Since all elements of $X_{I}$ are conjugate to $\theta_{C}$ under the action of $W_{I}$ by Proposition 5.1.2 (ii) and (iii), it now follows from the last paragraph that for every 7-root $\gamma\in X_{I}$, the components of $\gamma$ satsify the properties (a) and (b) satisfied by the components of $\theta_{C}$. This implies that the map $\varphi$ takes each element to a Steiner triple system (and thus one of the 30 inequivalent labellings of the Fano plane). The map $\varphi$ is clearly injective, and it is surjective because all Steiner triple systems are isomorphic via the permutation action of $S_{7}$ by Remark 3.3.2. This proves the first sentence of the proposition. The second sentence holds by inspection of the matrices $M_{C}$ and $M_{N}$. ∎
Remark 6.2.3.
The labellings canonically corresponding to $X_{I}$ have the following additional properties.
  1. (i)
    The triples $ijk$ in the labelling $L_{C}$ corresponding to the 7-root $\theta_{C}$ appear in [48, Section IV] as the triples indexing the “globally invariant linear forms” ±$\pm x_{i}\pm x_{j}\pm x_{k}$ of type $E_{7}$.
  2. (ii)
    The labelling $L_{N}$ corresponding $\theta_{N}$ is the unique labelling with the property that if the digits are written in binary, then the third digit of each triple is the bitwise exclusive or (XOR) of the other two.
  3. (iii)
    Recall from Remark 5.1.3 that $X_{I}$ naturally splits into two equal-sized components that are interchanged by the action of a reflection in $W_{I}$. As discussed in [43], any two distinct labellings in the same component have precisely one triple in common.
  4. (iv)
    The level $\lambda(\gamma)$ of each 7-root $\gamma$ in $X_{I}$ equals $(14-d)$, where $d$ is the number of 3-element subsets $E$ of the set $[7]$ with the property that the labelling $L_{\gamma}$ contains no blocks of the form $(E\setminus\{j\})\cup\{i\}$ where $i\leq j$ and $j\in E$. This fact can be verified computationally, and is similar to the interpretation of the level function in type $D_{2k}$ via Steiner systems $S(1,2,2k)$ given in Section 6.1.
Remark 6.2.4.
The noncrossing basis in type $E_{7}$ is illustrated by the diagram labelled ${\mathfrak{M}}_{6}$ in [13, Appendix], where each rectangle can be identified with a noncrossing basis element $\beta$. A label of $i$ on a rectangle indicates that $\alpha_{i}|\beta$. The edges connecting rectangles refer to star operations in the sense of [30], which can be interpreted directly in terms of $n$-roots as follows. If $i$ and $j$ are adjacent vertices of the Dynkin diagram, then two noncrossing basis elements $\beta$ and $\beta^{\prime}$ such that $\alpha_{i}|\beta$ and $\alpha_{j}|\beta^{\prime}$ are joined by an edge if we have $s_{i}s_{j}(b)=b^{\prime}$, or equivalently $s_{j}s_{i}(b^{\prime})=b$. (A similar construction appears in [24, Lemma 2.8].) Note that the Dynkin diagram of type $E_{7}$ in [13] differs from the Dynkin diagram of type $E_{7}$ shown in Figure 1 (d) in the labelling of vertices, but it can be obtained by removing the vertex “8” and its incident edge from the Dynkin diagram of type $E_{8}$ shown in Figure 1 (e).
Remark 6.2.5.
Ren–Sam–Schrader–Sturmfels [40, Theorem 4.1] give an “utterly explicit” basis for the 15-dimensional Macdonald representation in type $E_{7}$ that is natural in the context of the Göpel variety in algebraic geometry. The elements of the nonnesting basis and the noncrossing basis in type $E_{7}$ all factorize into linear factors in $\text{\rm Sym}(V^{*})$ by construction, but not all the basis elements of [40, Theorem 4.1] do, even after extending scalars to ${\mathbb{C}}$. It follows that the basis of [40] is not the same as either the noncrossing basis or the nonnesting basis, even after applying a change of basis of $V^{*}$.
6.3. Type $E_{8}$
Suppose $W$ has type $E_{8}$. We define $\nu_{A}$ to be the positive $8$-root with the following components:
$\alpha_{2},\ \alpha_{3},\ \alpha_{5},\ \alpha_{7},\ \alpha_{2}+\alpha_{3}+2\alpha_{4}+\alpha_{5},$
[…]
$2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+6\alpha_{4}+5\alpha_{5}+4\alpha_{6}+3\alpha_{7}+2\alpha_{8}.$
We define $\nu_{C}$ to be the positive $8$-root with the following components:
$\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}+\alpha_{5}+\alpha_{6}+\alpha_{7}+\alpha_{8},\quad\alpha_{1}+\alpha_{2}+\alpha_{3}+2\alpha_{4}+\alpha_{5}+\alpha_{6}+\alpha_{7},$
$\alpha_{1}+\alpha_{2}+\alpha_{3}+2\alpha_{4}+2\alpha_{5}+\alpha_{6},\quad\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4}+\alpha_{5}+\alpha_{6},$
$\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4}+2\alpha_{5}+\alpha_{6}+\alpha_{7},\quad\alpha_{1}+\alpha_{2}+2\alpha_{3}+3\alpha_{4}+2\alpha_{5}+\alpha_{6}+\alpha_{7}+\alpha_{8},$
$\alpha_{1}+\alpha_{2}+2\alpha_{3}+3\alpha_{4}+3\alpha_{5}+3\alpha_{6}+2\alpha_{7}+\alpha_{8},\quad\alpha_{1}+3\alpha_{2}+3\alpha_{3}+5\alpha_{4}+4\alpha_{5}+3\alpha_{6}+2\alpha_{7}+\alpha_{8}.$
We define $\nu_{N}$ to be the positive $8$-root with the following components:
$\alpha_{1},\quad\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4}+\alpha_{5},\quad\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4}+2\alpha_{5}+2\alpha_{6}+\alpha_{7},$
$\alpha_{1}+\alpha_{2}+2\alpha_{3}+3\alpha_{4}+2\alpha_{5}+2\alpha_{6}+\alpha_{7}+\alpha_{8},\quad\alpha_{1}+\alpha_{2}+2\alpha_{3}+3\alpha_{4}+3\alpha_{5}+2\alpha_{6}+2\alpha_{7}+\alpha_{8},$
$\alpha_{1}+2\alpha_{2}+2\alpha_{3}+3\alpha_{4}+2\alpha_{5}+2\alpha_{6}+2\alpha_{7}+\alpha_{8},\quad\alpha_{1}+2\alpha_{2}+2\alpha_{3}+3\alpha_{4}+3\alpha_{5}+2\alpha_{6}+\alpha_{7}+\alpha_{8},$
$\alpha_{1}+2\alpha_{2}+2\alpha_{3}+4\alpha_{4}+3\alpha_{5}+2\alpha_{6}+\alpha_{7}.$
Finally, we define the element $w\in W$ to be the element expressed by the word
$\mathbf{w}=(s_{1}s_{4}s_{6}s_{8})(s_{3}s_{5}s_{7})(s_{4}s_{6})(s_{2}s_{5})(s_{4})(s_{3})(s_{1}).$
The heap of $w$ is shown in Figure 2 (c).
Proposition 6.3.1.
If $W$ has type $E_{8}$, then the $8$-roots $\nu_{A}$, $\nu_{C}$ and $\nu_{N}$ given above are respectively the maximally aligned, maximally crossing, and maximally nesting $n$-roots of $W$. The element $w$ is the nonnesting element, and $\mathbf{w}$ a reduced word for it. The Macdonald representation $j^{E_{8}}_{8A_{1}}({\rm{sgn}})$ has dimension $50$.
Proof.
The statements can be proved using the same strategy used in the proof of Proposition 6.2.1 except for the following changes in numerical details. The number $M$ of coplanar quadruples in an $n$-root is now 14, and the number of order filters in the heap of the nonnesting element $w_{N}$ is 50. The components of the components of $\nu_{N}$ have heights 1, 7, 11, 13, 15, 15, 15, and 15, which implies that $\nu_{N}$ has no alignments by Proposition 3.2.5 (ii). Furthermore, if $\nu_{N}$ had a crossing, then Proposition 3.2.5 (iv) would imply that the only possibility is for the crossing to contain roots of heights 7, 11, 13, and 15, but this is also not possible either because $11+13>7+15$. ∎
Remark 6.3.2.
  1. (i)
    Schmidt [44, Lemma 3.4] gives an explicit partition of the 120 positive-negative pairs of roots in type $E_{8}$ into 15 sets of size 8. The components of $\theta_{A}$, $\theta_{N}$, and $\theta_{C}$ appear in Schmidt’s list as $V_{1}$, $V_{14}$, and $V_{15}$, respectively.
  2. (ii)
    In standard coordinates, the components of the maximally crossing and maximally nesting $n$-roots $\theta_{C}=\nu_{C}$ and $\theta_{N}=\nu_{N}$ are given by the rows of the following matrices $M^{\prime}_{C}$ and $M^{\prime}_{N}$, respectively:
    $M^{\prime}_{C}=\left[\begin{matrix}{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}\\ {\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}\\ {\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}\\ {\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}\\ {\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}\\ {\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}\\ {\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}\\ {\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}\\ \end{matrix}\right],\;M^{\prime}_{N}=\left[\begin{matrix}{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}\\ {\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}\\ {\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}\\ {\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}\\ {\text{--1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}\\ {\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}\\ {\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}\\ {\text{+1}}&{\text{--1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{+1}}&{\text{--1}}&{\text{+1}}\\ \end{matrix}\right].$
    All rows in $M^{\prime}_{C}$ other than the bottom row contain four “$+1$” and four “$-1$”, and the 14 quadruples recording the column numbers of the positive and negative entries in these rows form a Steiner quadruple system. Furthermore, these 14 quadruples are precisely the ones indexing the “globally invariant linear forms” of type $E_{8}$ in [48, Section V].
  3. (iii)
    The matrix $M^{\prime}_{C}$ above is a Hadamard matrix, meaning a matrix with entries in $\{+1,-1\}$ that has orthogonal rows (and, therefore, orthogonal columns). By rearranging the rows, the matrix can be expressed more simply as the Kronecker product $H\otimes H\otimes H$, where
    $H=\left[\begin{matrix}\text{+1}&\text{+1}\\ \text{--1}&\text{+1}\end{matrix}\right].$
Remark 6.3.3.
The noncrossing basis in type $E_{8}$ is illustrated by the diagram labelled ${\mathfrak{M}}_{50}$ in [12, Appendix].
The set $X_{I}$ of alignment-free positive $n$-roots in type $E_{8}$ can be used to give a convenient realization of the graph $\bar{\Gamma}_{1}$ studied by Schmidt in [44]. The graph $\bar{\Gamma}_{1}$, which is the complement of another graph $\Gamma_{1}$, has the property that it is quantum isomorphic (in the sense of [3]) but not isomorphic to the orthogonality graph $G_{E_{8}}$ of the roots of type $E_{8}$. The vertices of $G_{E_{8}}$ are the 120 positive roots of type $E_{8}$, and two roots are adjacent in $G_{E_{8}}$ if and only if they are orthogonal.
To realize the graph $\bar{\Gamma}_{1}$ via $X_{I}$, recall from Remark 5.1.3 that $X_{I}$ naturally splits into two equal-sized components, $X_{I}^{e}$ and $X_{I}^{o}$, which consist of all the elements in $X_{I}$ with even levels and odd levels, respectively. The components each have $240/2=120$ elements since $\lvert X_{I}\rvert=240$ by Proposition 7.1.1 (ii).
Definition 6.3.4.
We define $\Gamma$ to be the following graph: the vertex set is $X_{I}^{e}$, the set of all alignment-free positive $8$-roots of even parabolic level, and two vertices are adjacent if and only if they have no common components.
The next four paragraphs recall Schmidt’s construction of the graph $\Gamma_{1}$ and explain why $\Gamma$ is isomorphic to $\bar{\Gamma}_{1}$. (The isomorphism will also hold if we replace $X_{I}^{e}$ with $X_{I}^{o}$ in Definition 6.3.4.)
Start with the folded halved 8-cube, where the vertices are the 64 pairs of the form $\{x,\mathbf{1}+x\}$ for all length-8 binary strings $x\in\mathbb{F}_{2}^{8}$ with an even number of 1s (where $\mathbf{1}$ is the string with all entries equal to $1$). We can naturally identify these vertices with the 64 positive roots of the form $\alpha=(\sum_{i=1}^{7}\pm\varepsilon_{i})+\varepsilon_{8}$, via the bijection sending $\alpha$ to the pair $\{x_{\alpha},\mathbf{1}+x_{\alpha}\}$ where $x_{\alpha}$ is the string $(x_{i})_{i=1}^{8}$ such that $x_{i}=1$ if and only if $\varepsilon_{i}$ appears with coefficient $-1$ in $\alpha$ for all $i\in[8]$. The 64 positive roots of the form $(\sum_{i=1}^{7}\pm\varepsilon_{i})+\varepsilon_{8}$ are precisely the positive roots of $x$-height 1 in the sense of Remark 5.1.3, so they are also precisely the roots that can appear as a component of an 8-root in $X_{I}$ by Remark 5.1.3.
By definition, two vertices $\{x,\mathbf{1}+x\}$ and $\{y,\mathbf{1}+y\}$ in the folded halved cube are adjacent if and only if $x$ and $y$ differ in 2 or 6 entries. It follows that in the complement of the folded halved cube, two distinct vertices $\{x,\mathbf{1}+x\}$ and $\{y,\mathbf{1}+y\}$ are adjacent if and only if $x$ and $y$ differ in 4 positions. This complement is denoted $VO_{6}^{+}(2)$. The condition that $x$ and $y$ differ in 4 entries holds if and only if the positive roots corresponding to $\{x,\mathbf{1}+x\}$ and $\{y,\mathbf{1}+y\}$ are orthogonal, therefore each clique of size 8 in $VO_{6}^{+}(2)$ corresponds to an 8-root in $X_{I}$.
Schmidt defines the vertex set of $\Gamma_{1}$ to be any orbit of cliques of size 8 under the group ${\mathbb{Z}}_{2}^{6}\times A_{8}$ in $VO_{6}^{+}(2)$, where $A_{8}$ is the alternating subgroup of $S_{8}$. There are two such orbits, both of size 120, and the choice of the orbit does not matter, so we may assume that the orbit contains a clique corresponding to an 8-root with even level, i.e., to a vertex, $\beta$, of $\Gamma$. The vertices of $\Gamma_{1}$ then matches precisely the vertices of $\Gamma$ for the reasons sketched below. We have $I=S\setminus\{s_{1}\}$ by Proposition 6.3.1, therefore we have $W_{I}\cong W(D_{7})$. The action of $W_{I}=W(D_{7})$ on $X_{I}$ can be extended to an action of a larger subgroup $G\leq W(E_{8})$, generated by $W_{I}$ together with the reflection $s_{\theta}$ corresponding to the highest root $\theta=2(\varepsilon_{7}+\varepsilon_{8})$. We have $G\cong W(D_{8})\cong N\rtimes S_{8}$, where $N\cong{\mathbb{Z}}_{2}^{7}$ is the elementary abelian group of order $2^{7}$. By considerations involving $x$-heights (in the sense of Remark 5.1.3), each reflection in $G$ changes every 8-root in $X_{I}$, and when it does so it changes the parabolic level by an odd number because it moves a $C$ to an $N$ or vice versa. It follows that the commutator subgroup $G^{\prime}\cong{\mathbb{Z}}_{2}^{7}\times A_{8}$ of $G$ acts on $X_{I}$ with $X_{I}^{e}$ and $X_{I}^{o}$ as its orbits. This action induces a transitive action of $G^{\prime}/Z(G)\cong{\mathbb{Z}}_{2}^{6}\times A_{8}$ on $\Gamma$ that matches the action of ${\mathbb{Z}}_{2}^{6}\times A_{8}$ on $\Gamma_{1}$.
Two vertices in $\Gamma_{1}$ are defined to be adjacent if and only if they are cliques that intersect in exactly two elements from $VO_{6}^{+}(2)$. This occurs if and only if their correponding $8$-roots have two components in common. With some more work, or using computation, one can show that two distinct 8-roots whose levels have the same parity either have disjoint components or have exactly two components in common. It follows that $\Gamma$ is isomorphic to $\bar{\Gamma}_{1}$. To summarize, we have the following result.
Remark 6.3.5.
The graph $\Gamma$ from Definition 6.3.4, i.e., the graph whose vertices are the alignment-free positive $8$-roots of even parabolic level and where two vertices are adjacent if and only if they have no common component, is isomorphic to the graph $\bar{\Gamma}_{1}$ from [44]. As a consequence, the graph $\Gamma$ is quantum isomorphic but not isomorphic to the orthogonality graph $G_{E_{8}}$ of the $E_{8}$ root system.
The graphs $\Gamma_{E_{8}}$ and $\bar{\Gamma}_{1}$ are known to be strongly strongly regular graphs with parameters $(120,63,30,36)$. It follows that $\Gamma$ is also such a graph. Mathon and Street [34, Table 2.2] mention that the graphs $\Gamma_{E_{8}}$ and $\bar{\Gamma}_{1}$ each have 2025 $8$-cliques. The 8-cliques of $\Gamma_{E_{8}}$ are the positive $8$-roots of $E_{8}$, and it would be interesting to classify the 8-cliques of $\Gamma\cong\bar{\Gamma}_{1}$. Finally, we note that, as [44] points out, there are other constructions of $\Gamma_{1}$ in the literature [8, 34]. However, the construction in terms of $8$-roots has the advantages of being concise, and being clearly related to the $E_{8}$ root system.
Remark 6.3.6.
The group $\text{\rm Aut}(\Gamma_{1})$ acts as a permutation group of rank 4 on $\Gamma_{1}$ [8]. It follows that if two $n$-roots $x,y\in X_{I}$ both have even or odd levels, then $x$ and $y$ can be in one of four relative positions. These can be shown to be the following (where $N$ is the elementary abelian group of order $2^{7}$ mentioned above):
  • (i)
    $x=y$;
  • (ii)
    $x$ and $y$ have precisely two common components;
  • (iii)
    $x$ and $y$ have disjoint components, and $y=n.x$ for some $n\in N$;
  • (iv)
    $x$ and $y$ have disjoint components, and $y\neq n.x$ for any $n\in N$.
The situations in (iii) and (iv) correspond to the edges in the graph $\Gamma$, and they show that the edges of $\Gamma$ naturally split into two types. This is not the case for the graph $G_{E_{8}}$: the automorphism group of $G_{E_{8}}$ has rank 3, and two vertices can only be in three relative positions: equality, adjacency, and non-adjacency.
7. Concluding remarks
7.1. Poincaré polynomials
Rains and Vazirani [37, Section 8] define the Poincaré series of a quasiparabolic set to be $PS(q)=\sum_{x\in X}q^{\lambda(x)}$. They point out that in many cases, the Poincaré series factorizes in a very simple way, and the factors are often quantum integers
$[d]_{q}:=\frac{q^{d}-1}{q-1}=1+q+q^{2}+\cdots+q^{d-1}.$
These quantum integers often behave as if they were the degrees of polynomial invariants of a Coxeter group, and in some cases, the integers can be interpreted in terms of degrees of invariants in characteristic 2 (see [37, Section 8, Example 9.4]).
Proposition 7.1.1.
Let $W$ be a Weyl group of type $E_{7},E_{8}$, or $D_{n}$ for $n$ even.
  1. (i)
    For the quasiparabolic set $X$ for $W$ consisting of all positive $n$-roots of $W$, equipped with the level function such that $\lambda(\theta_{A})=0$, we have
    $PS_{X}(q)=\begin{cases}\prod_{i=2}^{k}[2i-1]_{q}&\text{\ if \ }W\text{\ is\ of\ type\ }D_{2k},\\ [3]_{q}[5]_{q}[9]_{q}&\text{\ if \ }W\text{\ is\ of\ type\ }E_{7},\\ [3]_{q}[5]_{q}[9]_{q}[15]_{q}&\text{\ if \ }W\text{\ is\ of\ type\ }E_{8}.\end{cases}$
  2. (ii)
    For the quasiparabolic set $X_{I}\subseteq X$ for $W_{I}$ consisting of the alignment-free positive $n$-roots of $W$ (with its level function inherited from $X$), we have
    $PS_{X_{I}}(q)=\begin{cases}q^{M}\prod_{i=2}^{k}[i]_{q}&\text{\ if \ }W\text{\ is\ of\ type\ }D_{2k},\\ q^{M}[2]_{q}[3]_{q}[5]_{q}&\text{\ if \ }W\text{\ is\ of\ type\ }E_{7},\\ q^{M}[2]_{q}[3]_{q}[5]_{q}[8]_{q}&\text{\ if \ }W\text{\ is\ of\ type\ }E_{8},\end{cases}$
    where $M$ is the level $\lambda(\theta_{C})$ of the minimal element of $X_{I}$ (which is also the number of coplanar quadruples in each $n$-root). In particular, each of the factors $[i]_{q}$ of $PS_{X_{I}}(q)$ corresponds to a factor $[2i-1]_{q}$ of $PS_{X}(q)$ in (i).
Proof.
We have verified both (i) and (ii) in types $E_{7}$ and $E_{8}$ computationally. (We do not have conceptual proofs at the moment.) For type $D_{n}$, part (i) follows from [45, Equation (5.4)], or [19, Theorem 4], or [14, Corollary 3.3], after we identify $X$ with the perfect matchings of the set $[n]$ as usual. Finally, part (ii) for type $D_{n}$ follows from Proposition 6.1.1 and the well-known form $\prod_{i=2}^{k}[i]_{q}$ for the Poincaré series $\sum_{\tau\in S_{k}}q^{l(\tau)}$ of the symmetric group $S_{k}$. ∎
Remark 7.1.2.
The exponents $3,5,9$ and $3,5,9,15$ that respectively appear in the Poincaré series $PS_{X}(q)$ of types $E_{7}$ and $E_{8}$ show up as the degrees of generators in the cohomology modulo 2 of compact exceptional Lie groups [2], and as the codimensions of generators of Chow rings associated to linear algebraic groups in characteristic 2 [36, Section 4].
Remark 7.1.3.
Recall from Proposition 5.5.1 that the set $X_{I}$ is the top $\sigma$-equivalence class with respect to the order $\leq_{\sigma}$. It turns out that for every $\sigma$-equivalence class $C$ in $X$, the polynomial $PS_{C}(q)=\sum_{x\in C}q^{\lambda(x)}$ has the form
(7.1.1) $PS_{C}(q)=\prod_{d\in D}q^{d-1}[d]_{q}$
for some set of nonnegative integers $D$. This is particularly remarkable because in general, there is no obvious way to turn a $\sigma$-equivalence class into a $W$-set for a suitable Weyl group $W$. In type $D_{n}$, the integers from $D$ have an interpretation in terms of rook placements [51, Theorem 1]. Summing over all $\sigma$-equivalence classes gives rise to the expression for $PS_{X}(q)$ that appears in the abstract of [6]. In types $E_{7}$ and $E_{8}$, we verified Equation (7.1.1) by computation.
7.2. Coxeter elements
Let $d_{1},d_{2},\ldots,d_{r}$ be the numbers that appear in the factorization $\prod_{i=1}^{r}[d_{i}]_{q}$ of the Poincare series $PS_{X}(q)$ in Proposition 7.1.1 (i). It follows easily from the definitions that $\prod_{i=1}^{r}d_{i}$ is the number of positive $n$-roots, and that $\sum_{i=1}^{r}(d_{i}-1)=2M$, where $M$ is the number of coplanar quadruples in an $n$-root. It also turns out that the largest integer $d_{r}$ in each case (which is $n-1$ in type $D_{n}$, is $9$ in type $E_{7}$, and is $15$ in type $E_{8}$) is equal to $h/2$, where $h$ is the Coxeter number.
We recall that, by definition, a Coxeter element is a product of all the simple reflections in some order, and the Coxeter number is the order of any Coxeter element $c$. All such elements are conjugate, and therefore have the same order. It turns out that $c^{h/2}$ acts as $-1$ in the reflection representation, and therefore acts trivially on the set $X$. If $C$ is the cyclic group of order $h/2$ generated by the action of $c$ on the positive $n$-roots, then it can be shown that the nonidentity elements of $C$ act without fixed points on the positive $n$-roots. The factor of $[h/2]_{q}$ in $PS_{X}(q)$ then implies that the triple $(X,PS_{X}(q),C)$ satisfies the cyclic sieving phenomenon of Reiner, Stanton, and White [39]: the number of fixed points of $c^{d}$ is equal to $PS_{X}(e^{2\pi id/m})$, where $m=h/2$.
It is possible, by choosing a suitable Coxeter element $c$ and $n$-root $\beta$, to find an orbit of $n$-roots
$O=\{c^{d}(\beta):0\leq d<h/2\}$
that contains every positive root exactly once as one of its components. This can be achieved in type $D_{n}$ by taking $\beta=\theta_{N}$ and $c=s_{1}s_{2}\cdots s_{n-1}s_{n}$. We also verified that such an orbit can also be found in types $E_{7}$ and $E_{8}$, although it is necessary to make a different choice of $\beta$. The existence of such an orbit $O$ in type $E_{8}$ is related to the Kochen–Specker theorem in quantum mechanics [50].
7.3. Feature-avoidance via quasiparabolic structure
It can be shown that each of the three types of feature-avoiding $n$-roots in the set $X$ can be characterized using only the parabolic structure of $X$, without reference to the combinatorics of $n$-roots. Specifically, the following holds for all $n$-roots $x\in X$:
  1. (i)
    $x$ is alignment-free if and only if there does not exist a reflection $r$ such that $\lambda(r(x))-\lambda(x)$ is a strictly positive even number;
  2. (ii)
    $x$ is noncrossing if and only if there is a sequence
    $x<_{Q}r_{1}(x)<_{Q}r_{2}r_{1}(x)<_{Q}\cdots<_{Q}x_{1}$
    where $x_{1}$ is the unique maximal element of $X$ and the level increases by $2$ at each step;
  3. (iii)
    $x$ is nonnesting if and only if there does not exist a reflection $r$ such that $\lambda(r(x))-\lambda(x)$ is a strictly negative even number.
In addition, Remark 5.5.3 shows that the $\sigma$-equivalence classes can be characterized as the Eulerian intervals between nonnesting and noncrossing elements. It may be interesting to use these characterizations to extend the notions of feature-avoiding elements and $\sigma$-equivalence to more general quasiparabolic sets.
Acknowledgements
We thank D.C. Ernst, H. Russell, and N. Thiem for some helpful conversations.